### v2

# Two-step estimation of latent trait models

Abstract
We consider two-step estimation of latent variable models, in which just the measurement model is estimated in the first step and the measurement parameters are then fixed at their estimated values in the second step where the structural model is estimated. We show how this approach can be implemented for latent trait models (item response theory models) where the latent variables are continuous and their measurement indicators are categorical variables. The properties of two-step estimators are examined using simulation studies and applied examples. They perform well, and have attractive practical and conceptual properties compared to the alternative one-step and three-step approaches. These results are in line with previous findings for other families of latent variable models. This provides strong evidence that two-step estimation is a flexible and useful general method of estimation for different types of latent variable models.
Key words: Item response theory models; latent variable models; structural equation models; pseudo-maximum likelihood estimation
1 Introduction
Latent variable models have normally two main parts: a measurement model which describes how the latent variables that appear in the model are measured by observed indicators of them, and a structural model which describes the associations among the latent variables and any observed explanatory and response variables which are not treated as measurement indicators. For instance, in the illustrative example considered in Section 5 of this paper the structural model specifies how individuals’ extrinsic and intrinsic work value orientations are associated with characteristics of the individuals, and the measurement model specifies how these value orientations are measured by a set of survey questions. Here, as in many applications, the structural model is the focus of substantive interest, but the measurement model also needs to be included and estimated in order for the structural model to be estimable.
Estimation of these elements can be organised in different ways. In joint or one-step estimation, both parts of the model are estimated together. When this is done by maximizing the joint likelihood, one-step estimates are the maximum likelihood (ML) estimates of the model parameters. In contrast, “stepwise” approaches divide the estimation into separate steps. The most familiar of them is three-step estimation. In its first step, the measurement model is estimated from a specification which omits all or most of the structural model. In the second step, this estimated measurement model is used to assign predicted values of the latent variables to the units of analysis, such as factor scores for continuous latent variables. In the third step, these scores are used in place of the latent variables to estimate the structural model. There are “naive” and “adjusted” versions of three-step estimation, depending on whether it attempts to account for the measurement error that results from using the predicted values.
The focus of this paper is on a different stepwise approach, two-step estimation of latent variable models. Its first step is estimation of the measurement model, as in the three-step method. In the second step, instead of being used to calculate explicit predictions of the latent variables, the parameters of the measurement model are simply fixed at their estimated values. In other words, the second step takes the same form as one-step estimation, except that all the measurement parameters are treated as known numbers rather than unknown estimands. Likelihood-based two-step estimation derives its justification and properties from the general theory of pseudo maximum likelihood estimation. Its large-sample properties are very similar to those of one-step ML estimation.
Two-step estimates avoid the measurement error bias of naive three-step estimates, and are typically comparable in performance but more generally usable and practically simpler than adjusted three-step estimates. Compared to one-step estimation, the two-step approach has in principle attractive practical and conceptual advantages. Because it is split into two steps, both of them will be computationally less demanding than the single step of one-step estimation. It is also easy to estimate the measurement and structural models using fully or partially different sets of data, which is useful in some applications. The conceptual advantage arises from an inherent difference in how the two methods use the available data. Because the one-step approach fits all parts of the model at once, the estimated measurement model is informed not only by the measurement indicators but also by observed covariates and responses in the structural model. This can cause what Burt (1973, 1976) terms “interpretational confounding”, a situation where the implied definition of a latent variable is partly determined by variables that should be conceptually separate from it. In other words, if we change the specificaton of the structural model, this also changes the estimated measurement model and hence in effect changes the definition of the latent variables that appear in the structural model. This circularity is avoided by the two-step approach because it estimates the measurement model only once and using only the measurement indicators.
Two-step estimation is a general idea, and one of the points that we want to make in this paper is that it can in principle be applied to any latent variable models. However, there may be differences in performance of the estimates, ease of implementation, or other points to consider when it is used for different broad classes of models. In the literature so far, two-step estimation has accordingly been described separately for different types of models. It was proposed for latent class analysis, that is for models where both the latent variables and their indicators are categorical, by Bandeen-Roche et al. (1997), Xue and Bandeen-Roche (2002) and Bakk and Kuha (2018), and has since been extended to further versions of them (e.g. multilevel latent class analysis in Di Mari et al. 2023). For structural equation models (SEMs), where the latent variables and indicators are both continuous, the idea was introduced already by Burt (1973, 1976), but detailed exploration of it is much more recent. In particular, Rosseel and Loh (2022) proposed a general two-step approach for SEMs, with the title of “structural-after-measurement” (SAM) estimation. They describe two (usually equivalent) forms of it: “local SAM” which summarises the first step in the form of the estimated means and covariance matrix of the variables in the structural model, and “global” SAM which is analogous to the implementation that we describe in this paper for latent trait models (the moment-based “local” approach is not really applicable to them). A method for SEMs is also proposed by Levy (2023), using Bayesian (MCMC) estimation for both steps. Another closely related paper is Skrondal and Kuha (2012), who use two-step estimation to correct for covariate measurement error in regression models. Two-step methods for these classes of models are now also being implemented in general-purpose software for latent variable modelling, in Latent Gold for latent class analysis (Vermunt and Magidson, 2021b) and in the R package lavaan for SAM (Rosseel, 2012; Rosseel and Loh, 2022).
In this paper we propose and examine two-step estimation for another family of latent variable models, one where the latent variables are continuous but their indicators are categorical. These are commonly known as Item Response Theory (IRT) models. We refer to them also as latent trait models. They resemble SEMs in having continuous latent variables, and standard latent class models in having categorical indicators. Like SEMs, they are often used with complex structural models which include multiple latent variables, whereas latent class analysis more often includes only one latent class variable. Computationally, latent trait models are typically the most demanding of these families of models, at least for ML estimation, because their log likelihood cannot be expressed in a closed form.
We describe the theory and implementation of two-step estimation of latent trait models. The presentation of this broadly parallels that of Bakk and Kuha (2018) for latent class models, but with the changes that are needed now that the latent variables are continuous. We then use simulation studies and an application example to explore the properties of the method in this context. The focus of comparisons is between two-step and one-step estimates, because generally applicable adjusted three-step estimates are not available for latent trait models. One question of interest is then whether the relative performance of the two-step method is similar to what it is for previously considered types of models. The conclusion is that it is, with some differences of emphasis (these findings are discussed in more detail in Sections 46). Two-step estimates of the structural model again perform very well. In complex models their computational advantage over one-step estimates is even larger than for latent class models and SEMs. On the other hand, interpretational confounding of one-step estimates may be less worrying here than for latent class models, because shifts in interpretation tend to be less abrupt for continuous latent variables than for categorical ones.
The latent trait model specifications that we consider are introduced in Section 2. The definition, implementation and properties of two-step estimation for them are described in Section 3. Simulation studies are reported in Section 4 and the real-data example in Section 5, and concluding discussion is given in Section 6. Extended tables of some of the results from the simulations and data analysis are given in a supplementary appendix. Computer code for the estimation in R and Mplus (both of which are needed) and replication code for the applied example are also provided as supplementary materials.
2 Models and variables
To begin with a general formulation of a latent variable model, let $\boldsymbol{\eta}_{i}$ be a vector of latent variables, and $\mathbf{Y}_{i}$ and $\mathbf{X}_{i}$ vectors of observed variables, for a unit of analysis $i$. Consider a model of the form $p(\mathbf{Y}_{i},\boldsymbol{\eta}_{i}|\mathbf{X}_{i};\boldsymbol{\theta})=p(\mathbf{Y}_{i}|\boldsymbol{\eta}_{i},\mathbf{X}_{i};\boldsymbol{\theta}_{1})\,p(\boldsymbol{\eta}_{i}|\mathbf{X}_{i};\boldsymbol{\theta}_{2})$, where $p(\cdot|\cdot)$ denotes a conditional distribution and $\boldsymbol{\theta}=(\boldsymbol{\theta}_{1}^{\prime},\boldsymbol{\theta}_{2}^{\prime})^{\prime}$ are parameters. Here $p(\mathbf{Y}_{i}|\boldsymbol{\eta}_{i},\mathbf{X}_{i};\boldsymbol{\theta}_{1})$ is the measurement model for $\boldsymbol{\eta}_{i}$ and $p(\boldsymbol{\eta}_{i}|\mathbf{X}_{i};\boldsymbol{\theta}_{2})$ the structural model. The observed variables $\mathbf{Y}_{i}$ (the measurement items) serve as measures of the latent $\boldsymbol{\eta}_{i}$, and $\mathbf{X}_{i}$ are exogenous observed explanatory variables. Endogenous observed variables can also be included, by representing them as elements of $\boldsymbol{\eta}_{i}$ which are perfectly measured by single items in $\mathbf{Y}_{i}$. The parameters in $\boldsymbol{\theta}_{1}$ determine the measurement model, and we refer to them as the measurement parameters, and $\boldsymbol{\theta}_{2}$ (structural parameters) determine the structural model. It is assumed that $\boldsymbol{\theta}_{1}$ and $\boldsymbol{\theta}_{2}$ are distinct, in the sense that the joint parameter space for $\boldsymbol{\theta}$ is the product of the spaces for $\boldsymbol{\theta}_{1}$ and $\boldsymbol{\theta}_{2}$.
In two-step estimation, $\boldsymbol{\theta}_{1}$ are estimated first and then fixed at their estimated values in the second step where $\boldsymbol{\theta}_{2}$ are estimated. This is a very general approach which can in principle be used for any instance of this general model. For specificity, however, in this paper we focus on its use for a particular class of latent variable models. Here the items in $\mathbf{Y}_{i}$ are categorical variables, $\boldsymbol{\eta}_{i}$ are continuous and normally distributed, the structural model is recursive, and the measurement models exclude direct effects from $\mathbf{X}_{i}$ to $\mathbf{Y}_{i}$. This specification is described in this section. Some further variants and extensions of it are discussed in Section 3.5.
Consider first the structural model $p(\boldsymbol{\eta}_{i}|\mathbf{X}_{i};\boldsymbol{\theta}_{2})$. A general expression of it for the class of models that we focus on is
$\boldsymbol{\eta}_{i}=\mathbf{B}_{0}+\mathbf{B}_{\eta}\boldsymbol{\eta}_{i}+\mathbf{B}_{x}\mathbf{X}_{i}+\boldsymbol{\zeta}_{i},$ (1)
where $\mathbf{B}_{0}$, $\mathbf{B}_{\eta}$ and $\mathbf{B}_{x}$ are matrices of parameters, and $\boldsymbol{\zeta}_{i}\sim N(\mathbf{0},\boldsymbol{\Psi})$ are normally distributed model residuals. This is a variant of a formulation that is commonly used to specify structural equation models in a compact and general form. For our purposes, however, it is more convenient to write the model in a different way, in terms of a chain of conditional distributions. This is possible because we consider only recursive models, obtained from (1) by imposing appropriate constraints on $\mathbf{B}_{\eta}$. Suppose that $\boldsymbol{\eta}_{i}=(\eta_{i1},\dots,\eta_{iK})^{\prime}$ includes $K\geq 1$ variables. Suppose further that the model specification groups them into $1\leq M\leq K$ blocks of variables $\boldsymbol{\eta}_{i(m)}$, so that we can also write $\boldsymbol{\eta}_{i}=(\boldsymbol{\eta}_{i(1)}^{\prime},\dots,\boldsymbol{\eta}_{i(M)}^{\prime})^{\prime}$ in this pre-specified order. The structural model can then be written as
$p(\boldsymbol{\eta}_{i}|\mathbf{X}_{i};\boldsymbol{\theta}_{2})$ $=$ $p(\boldsymbol{\eta}_{i(1)}|\mathbf{X}_{i};\boldsymbol{\theta}_{21})\,p(\boldsymbol{\eta}_{i(2)}|\mathbf{X}_{i},\boldsymbol{\eta}_{i(1)};\boldsymbol{\theta}_{22})$
××$\hskip 10.00002pt\times\cdots\times p(\boldsymbol{\eta}_{i(M)}|\mathbf{X}_{i},\boldsymbol{\eta}_{i(1)},\dots,\boldsymbol{\eta}_{i(M-1)};\boldsymbol{\theta}_{2M})$ (2)
where the structural parameters are partitioned correspondingly as $\boldsymbol{\theta}_{2}=(\boldsymbol{\theta}_{21}^{\prime},\dots,\boldsymbol{\theta}_{2M}^{\prime})^{\prime}$. It follows from (1) that all of these conditional distributions are (multivariate or univariate) normal and specified by the linear models
$p(\boldsymbol{\eta}_{i(1)}|\mathbf{X}_{i};\boldsymbol{\theta}_{21})$ $\sim$ $N(\boldsymbol{\beta}_{10}+\boldsymbol{\beta}_{1x}\mathbf{X}_{i},\,\boldsymbol{\Psi}_{1})$ (3)
$p(\boldsymbol{\eta}_{i(m)}|\mathbf{X}_{i},\boldsymbol{\eta}_{i(1)},\dots,\boldsymbol{\eta}_{i(m-1)};\boldsymbol{\theta}_{2m})$ $\sim$ $N(\boldsymbol{\beta}_{m0}+\boldsymbol{\beta}_{mx}\mathbf{X}_{i}+\sum_{l=1}^{m-1}\boldsymbol{\beta}_{ml}\boldsymbol{\eta}_{i(l)},\,\boldsymbol{\Psi}_{m})$ (4)
for $m=2,\dots,M$, so that each of $\boldsymbol{\theta}_{21},\dots,\boldsymbol{\theta}_{2M}$ consists of the corresponding $\boldsymbol{\beta}$ and $\boldsymbol{\Psi}$ parameters. Constraints on elements of $\boldsymbol{\theta}_{2}$ can be included, most often zero constraints on the $\boldsymbol{\beta}$ parameters which omit some of the explanatory variables from some parts of the model and thus yield different non-saturated choices for the overall structural model (2). If observed covariates $\mathbf{X}_{i}$ are not present, they are omitted from (2)–(4). If any $\boldsymbol{\eta}_{i(m)}$ are actually observed (rather than latent) variables, it would also be straightfoward to replace the linear model with some other model for them. For example, if $\boldsymbol{\eta}_{i(M)}$ is a single binary observed variable, we could specify a binary logistic model for it given $\mathbf{X}_{i}$ and $\boldsymbol{\eta}_{i(1)},\dots,\boldsymbol{\eta}_{i(M-1)}$.
Consider then the measurement model. We assume throughout that $p(\mathbf{Y}_{i}|\boldsymbol{\eta}_{i},\mathbf{X}_{i};\boldsymbol{\theta}_{1})=p(\mathbf{Y}_{i}|\boldsymbol{\eta}_{i};\boldsymbol{\theta}_{1})$, so that there is no non-equivalence of measurement (differential item functioning) with respect to observed covariates $\mathbf{X}_{i}$. The overall model can then be written as
$p(\mathbf{Y}_{i},\boldsymbol{\eta}_{i}|\mathbf{X}_{i};\boldsymbol{\theta})=p(\mathbf{Y}_{i}|\boldsymbol{\eta}_{i};\boldsymbol{\theta}_{1})\,p(\boldsymbol{\eta}_{i}|\mathbf{X}_{i};\boldsymbol{\theta}_{2}),$ (5)
which also implies a model for the observed variables as
[…]
We make the common assumption that the items in $\mathbf{Y}_{i}$ are conditionally independent of each other given $\boldsymbol{\eta}_{i}$. Typically each latent variable in $\boldsymbol{\eta}_{i}$ is measured by only a subset of the items in $\mathbf{Y}_{i}$, and the simplest structure is obtained when each item measures exactly one variable. In that case $\mathbf{Y}_{i}=(\mathbf{Y}_{i1}^{\prime},\dots,\mathbf{Y}_{iK}^{\prime})^{\prime}$ can be partitioned corresponding to $\boldsymbol{\eta}_{i}=(\eta_{i1},\dots,\eta_{iK})^{\prime}$, so that only items in $\mathbf{Y}_{ik}=(Y_{ik1},\dots,Y_{ikp_{k}})^{\prime}$ are measures of $\eta_{ik}$, for each $k=1,\dots,K$. The measurement model can then be written as
$p(\mathbf{Y}_{i}|\boldsymbol{\eta}_{i};\boldsymbol{\theta}_{1})=\prod_{k=1}^{K}\,p(\mathbf{Y}_{ik}|\eta_{ik};\boldsymbol{\theta}_{1k})=\prod_{k=1}^{K}\prod_{j=1}^{p_{k}}\,p(Y_{ikj}|\eta_{ik};\boldsymbol{\theta}_{1k})$ (7)
so that $\boldsymbol{\theta}_{1}=(\boldsymbol{\theta}_{11}^{\prime},\dots,\boldsymbol{\theta}_{1K}^{\prime})^{\prime}$, with the different $\boldsymbol{\theta}_{1k}$ taken to be distinct from each other. We consider measurement models of this form in most of our specific examples for simplicity, but two-step estimation is not limited to this case. In the simulation studies Section 4 we also include models where some items measure more than one latent trait.
Specification of the measurement models for individual items $Y_{ijk}$ depends on the forms of the items. We consider situations where each item is a categorical variable with a finite number of possible values. Most of our specific examples will be ones where the items are binary. Coding their values as 0 and 1, in that case we specify the model for each item in (7) as the logit model
$\text{logit}[P(Y_{ikj}=1|\eta_{ik};\boldsymbol{\theta}_{1k})]=\tau_{kj}+\lambda_{kj}\eta_{ik}$ (8)
so that $\boldsymbol{\theta}_{1k}=(\tau_{k1},\dots,\tau_{kp_{k}},\lambda_{k1},\dots,\lambda_{kp_{k}})^{\prime}$. In the language of item response theory modelling, this is a two-parameter logistic (2-PL) model. There can be parameter constraints within $\boldsymbol{\theta}_{1k}$, such as taking all $\lambda_{kj}$ for the same $k$ to be equal. Other model specifications will be used if an item has $L>2$ categories. If they are regarded as ordered, the binary model can be replaced by, for example, the ordinal logistic model $\text{logit}[P(Y_{ikj}\leq l|\eta_{ik};\boldsymbol{\theta}_{1k})]=\tau_{kjl}-\lambda_{kj}\eta_{ik}$ for categories $l=1,\dots,L-1$. For items with unordered categories, we can use the multinomial logistic model $\log[P(Y_{ikj}=l|\eta_{ik};\boldsymbol{\theta}_{1k})/P(Y_{ikj}=1|\eta_{ik};\boldsymbol{\theta}_{1k})]=\tau_{kjl}+\lambda_{kjl}\eta_{ik}$ for $l=2,\dots,L$.
3 Two-step estimation of the structural model
3.1 Two-step point estimates
Suppose that we observe $(\mathbf{Y}_{i},\mathbf{X}_{i})$ for units $i=1,\dots,n$, assumed to be independent of each other. We use likelihood-based estimation in a parametric framework. The log-likelihood function for the class of models models that we consider is
[…]
where the structural and measurement models are specified as described in Section 2. The measurement model on the second line of (9) is for the simple measurement structure where each item measures exactly one latent variable $\eta_{ik}$, and it would be modified appropriately if this was not the case. Maximizing (9) with respect to $\boldsymbol{\theta}=(\boldsymbol{\theta}_{1}^{\prime},\boldsymbol{\theta}_{2}^{\prime})^{\prime}$ would give the overall maximum likelihood estimates (one-step estimates) of $\boldsymbol{\theta}$; we will discuss them further in Section 3.4.1 The integrals in (9) and in (10) below are not available in a closed form, so numerical integration is needed to evaluate them in estimation.
In step 1 of two-step estimation, the measurement parameters $\boldsymbol{\theta}_{1}$ are estimated from a simpler specification. When the measurement structure is simple, this can be done separately for each $\eta_{ik}$, using the log-likelihoods
[…]
for $k=1,\dots,K$, where $p(\eta_{ik};\mu_{k},\sigma^{2}_{k})\sim N(\mu_{k},\sigma^{2}_{k})$ and $\boldsymbol{\psi}_{k}=(\boldsymbol{\theta}_{1k}^{\prime},\mu_{k},\sigma^{2}_{k})^{\prime}$. Maximizing (10) gives estimate $\tilde{\boldsymbol{\psi}}_{k}$, from which $\tilde{\boldsymbol{\theta}}_{1k}$ are the step-1 estimates of $\boldsymbol{\theta}_{1k}$ and $\tilde{\mu}_{k}$ and $\tilde{\sigma}_{k}^{2}$ are discarded.
In step 2 of two-step estimation, the parameters $\boldsymbol{\theta}_{2}$ of the structural model are then estimated, holding $\boldsymbol{\theta}_{1}$ fixed at their estimates $\tilde{\boldsymbol{\theta}}_{1}$ from step 1. Here the log-likelihood is $\ell(\tilde{\boldsymbol{\theta}}_{1},\boldsymbol{\theta}_{2})$, which is of the same form as the one-step log-likelihood $\ell(\boldsymbol{\theta})=\ell(\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2})$ in (9) but with the fixed values $\tilde{\boldsymbol{\theta}}_{1}$ substituted for $\boldsymbol{\theta}_{1}$. Maximizing this with respect to $\boldsymbol{\theta}_{2}$ gives the two-step estimate of the parameters of the structural model, which we denote by $\tilde{\boldsymbol{\theta}}_{2}$.
The same step-1 estimate $\tilde{\boldsymbol{\theta}}_{1}$ can be used for any structural models for the same $\boldsymbol{\eta}_{i}$ in step 2, for example if we want to compare models with different choices of the covariates $\mathbf{X}_{i}$ or with some associations omitted in the structural model. Note also that for simplicity the notation in (9) and (10) assumes that $\ell(\boldsymbol{\theta})$ and all $\ell(\boldsymbol{\psi}_{k})$ use the same set of $n$ observations. This is not necessary in general. Two-step estimation also allows partially or completely different sets of data to be used for estimating $\tilde{\boldsymbol{\theta}}_{1}$ in step 1 and $\tilde{\boldsymbol{\theta}}_{2}$ in step 2, or for estimating $\tilde{\boldsymbol{\theta}}_{1k}$ for different $k$ in step 1, as long as we can assume that the same measurement model holds in all of them.
Compared with latent class models (as discussed in Bakk and Kuha 2018), there are some additional considerations here where the latent variables $\boldsymbol{\eta}_{i}$ are continuous. First, as usual for any models with continuous latent variables, some parameter constraints are needed to fix the scales of $\boldsymbol{\eta}_{i}$ and thus to identify the parameters $\boldsymbol{\theta}$. In two-step estimation they are imposed in the first step, and the scale implied by them then carries over to the second step where no further constraints are required. We identify the scale of each $\eta_{ik}$ by fixing parameters in the measurement model, for example $\tau_{kj}=0$ and $\lambda_{kj}=1$ for one $j$ in the logistic model in (8). This anchors the latent scale in the items $\mathbf{Y}_{i}$, leaving the distribution of $\boldsymbol{\eta}_{i}$ free. An alternative first-step constraint would be to set $\mu_{k}=0$ and $\sigma^{2}_{k}=1$ for all $k$, implying that each latent variable is on a scale where its marginal distribution is standard normal.
Another difference to typical latent class models is that in applications of latent trait models it is more common for the latent $\boldsymbol{\eta}_{i}$ to be multivariate. The measurement model can then be correspondingly more complex. Our recommendation is that step 1 of the estimation should then be divided into the simplest models that allow correct estimation of the parameters of the specified measurement model. When this model has simple structure, this means estimating single-trait models (10) for each latent trait separately. In other cases this is generalised only as far as is needed. For example, suppose that the only deviation from a simple measurement structure is that some items measure both $\eta_{i1}$ and $\eta_{i2}$. Then the measurement parameters for them would be estimated from a two-trait model with a bivariate normal distribution for ($\eta_{i1},\eta_{i2}$), and the rest of $\tilde{\boldsymbol{\theta}}_{1}$ from one-trait models for any other traits. This keeps step 1 as simple as possible, which is a key benefit of two-step estimation. An alternative would be to estimate all of the measurement models at once, from a $K$-trait model for a $K$-variate normal $\boldsymbol{\eta}_{i}$. In other words, this would be obtained by omitting just $\mathbf{X}_{i}$ from the structural model, and ignoring any parameter constraints in it. This, however, would make the first step more complex, often with no clear benefit. Rosseel and Loh (2022) give a good discussion of the considerations of how to organise step 1 in this respect. They too recommend estimating the measurement model of each distinct block of latent variables separately in most cases.
In practice some of the observed variables $\mathbf{X}_{i}$ and $\mathbf{Y}_{i}$ may be missing. Any missing items in $\mathbf{Y}_{i}$ are taken to be missing at random (MAR), and their contributions are simply omitted from (9) and (10). In (9) we take $\mathbf{X}_{i}$ to be complete, implying that units with any missing values in $\mathbf{X}_{i}$ are omitted from estimation of the structural model. This gives valid estimates for the model as long as the missingness in $\mathbf{X}_{i}$ does not depend on the response variables $\boldsymbol{\eta}_{i}$ (or on $\mathbf{Y}_{i}$). Multiple imputation could be used to include also incomplete observations of $\mathbf{X}_{i}$. These assumptions are conventional, and their implications are mostly the same for both two-step and one-step estimation. There are, however, some small differences which favour the two-step approach here. First, if the MAR assumption fails only for some items, say those in $\mathbf{Y}_{il}$, this affects only the estimates of the measurement model they contribute to, i.e. only $\tilde{\boldsymbol{\theta}}_{1l}$. This is because $\tilde{\boldsymbol{\theta}}_{1k}$ for different $k$ are estimated separately from each other in step 1. Second, even if we omit some units with incomplete data on $\mathbf{X}_{i}$ in step 2, we can still use them in step 1 where $\mathbf{X}_{i}$ are not involved, hence reducing the loss of information from missing data.
3.2 Asymptotic properties and variance estimation
The properties of the step-2 estimates $\tilde{\boldsymbol{\theta}}_{2}$ follow from the general theory of pseudo-maximum likelihood (PML) estimation (Gong and Samaniego, 1981). Much of the presentation here is similar to that of Bakk and Kuha (2018) for latent class models, because the general form of these results is the same irrespective of the details of the model specification.
PML refers to the general approach where one set of parameters of a model (here $\boldsymbol{\theta}_{1}$) are estimated first, and their estimates are then treated as known in a log likelihood (here $\ell(\tilde{\boldsymbol{\theta}}_{1},\boldsymbol{\theta}_{2})$) that is maximized to estimate a second set of parameters ($\boldsymbol{\theta}_{2}$). PML estimators are consistent and asymptotically normally distributed under very general regularity conditions. In our situation these require, broadly, that the joint model is such that the one-step ML estimator $\hat{\boldsymbol{\theta}}$ is consistent for $\boldsymbol{\theta}$, that the values of $\boldsymbol{\theta}_{1}$ and $\boldsymbol{\theta}_{2}$ can vary independently of each other, and that the step-1 estimator $\tilde{\boldsymbol{\theta}}_{1}$ is consistent for $\boldsymbol{\theta}_{1}$. Here these conditions are satisfied, with the partial exception of one approximation which is discussed further at the end of this section.
To get the variance matrix of $\tilde{\boldsymbol{\theta}}_{2}$, let the Fisher information matrix for $\boldsymbol{\theta}$ in the full model be
[…]
where $\boldsymbol{\theta}^{*}$ denotes the true value of $\boldsymbol{\theta}$ and the partitioning corresponds to $\boldsymbol{\theta}_{1}$ and $\boldsymbol{\theta}_{2}$. Let $\mathbf{\Sigma}_{11}/n$ be the asymptotic variance matrix of the step-1 estimator $\tilde{\boldsymbol{\theta}}_{1}$. The asymptotic variance matrix of the two-step estimator $\tilde{\boldsymbol{\theta}}_{2}$ is then $\mbox{var}(\tilde{\boldsymbol{\theta}}_{2})=\mathbf{V}/n$, where
[…]
(if $\tilde{\boldsymbol{\theta}}_{1}$ was obtained using $n_{1}\neq n$ of observations, $\mathbf{\Sigma}_{11}$ is multiplied by ($n/n_{1}$) in this; see Xue and Bandeen-Roche (2002) and Bakk and Kuha (2018)). Here $\mathbf{V}_{2}$ describes the variability in $\tilde{\boldsymbol{\theta}}_{2}$ if $\boldsymbol{\theta}_{1}$ were known, and $\mathbf{V}_{1}$ the additional variability arising from estimating $\boldsymbol{\theta}_{1}$ by $\tilde{\boldsymbol{\theta}}_{1}$.
[…]
In practical data analysis, we could also choose to omit $\mathbf{V}_{1}$ from the variance. This would mean, in effect, that once step 1 was done we would be treating the measurement model obtained from it, and thus the definition of the latent variables $\boldsymbol{\eta}_{i}$, as known and fixed rather than as an estimable characteristic. Estimated structural models obtained from step 2 would then be models for this definition of latent variables. This is close in spirit to the approach that is also almost always adopted in naive three-step estimation, as discussed further in Section 3.4.2 below. This may be desirable or unproblematic in many applications. When it is not, however, omitting the contribution from $\mathbf{V}_{1}$ may result in serious underestimation of the variance of $\tilde{\boldsymbol{\theta}}_{2}$.
[…]
3.3 Implementation of the estimation
Software that can do one-step estimation of a model can also be used for two-step estimation of it. We have used Mplus 6.12 software (Muthén and Muthén, 2010) for the simulation studies and application example of this paper. We have supplemented it with functions in R (R Core Team, 2022) to manage the estimation process. This automatically sequences the two steps of estimation, passing the estimated parameter values from the first step to the code for the second and the estimates back to R. We have used the MplusAutomation package in R (Hallquist and Wiley, 2018) to control Mplus from R, and the brew package (Horner, 2011) to automatically edit the input files. Examples of the estimation code are included in supplementary materials for the paper.
Because the log likelihood functions can have multiple local maxima, it is desirable to use multiple starting values for the estimation algorithm. This is a standard feature in software such as Mplus. High-quality software is also needed to implement and control the numerical integration that is needed in all parts of the estimation. These numerical challenges apply to both two-step and one-step estimation of latent trait models, although the starting point issue may be somewhat smaller for the simpler two-step approach.
[…]
3.4 Alternatives: One- and three-step estimation
3.4.1 One-step estimation
The main alternative to two-estimation is one-step estimation. It means obtaining maximum likelihood (ML) estimates $\hat{\boldsymbol{\theta}}=(\hat{\boldsymbol{\theta}}_{1},\hat{\boldsymbol{\theta}}_{2})$ of the measurement parameters and structural parameters together, by maximizing the log-likelihood (9) with respect to $\boldsymbol{\theta}$. This can be done using the same software (in our applications Mplus) that can also be used for two-step estimation.
[…]
3.4.2 Naive three-step estimation
Three-step estimation is a stepwise approach. Its first step is the same as in the two-step method, i.e. obtaining estimates $\tilde{\boldsymbol{\theta}}_{1}$ of the measurement parameters from simplest possible models. In the second step, these are used to calculate predicted values $\tilde{\eta}_{ik}$ for the latent variables, for each $i=1,\dots,n$; $k=1,\dots,K$. In the third step, the structural model is estimated with $\tilde{\boldsymbol{\eta}}_{i}=(\tilde{\eta}_{i1},\dots,\tilde{\eta}_{iK})^{\prime}$ substituted for $\boldsymbol{\eta}_{i}$ and treated as observed variables. For the recursive model (2)–(4) this means simply fitting the $M$ regression models for $\tilde{\boldsymbol{\eta}}_{i(1)},\dots,\tilde{\boldsymbol{\eta}}_{i(M)}$ separately. Estimated variances of the parameter estimates are also obtained as in standard regression, in effect treating $\tilde{\eta}_{ik}$ like any observed variables. We refer to this version of the approach as “naive” three-step estimation.
The predictions (or “latent trait scores”) $\tilde{\eta}_{ik}$ could be obtained in different ways. The most common, and the ones we use here, are the conditional expected values (empirical Bayes predictions) $\tilde{\eta}_{ik}=\mbox{E}(\eta_{k}|\mathbf{Y}_{i(k)};\tilde{\boldsymbol{\psi}}_{(k)})$, where $\mathbf{Y}_{i(k)}$ denotes those items which were included in step-1 estimation of the measurement model for $\eta_{ik}$ and $\tilde{\boldsymbol{\psi}}_{(k)}$ the estimated parameters of this model. These $\tilde{\eta}_{ik}$ require numerical integration, but they are normally provided by estimation software.
This approach is “naive” because in most cases it produces biased estimates of the parameters of the structural model. This is because $\tilde{\eta}_{ik}$ are erroneously measured versions of $\eta_{ik}$. Trying to correct for this bias leads to “adjusted” versions of three-step estimation, which we discuss in the next section. Before that, however, it is worth noting that whether we would want to do that depends on what we think the goal of three-step estimation is. Measurement error bias is a problem if we want to treat the step-3 models as estimates of the regression models (2)–(4) for the latent variables $\boldsymbol{\eta}_{i}$. But it is not a problem if we view them instead as estimates for the same models for the scores $\tilde{\boldsymbol{\eta}}_{i}$ themselves. In other words, we can choose to treat $\tilde{\boldsymbol{\eta}}_{i}$ as variables of interest rather than as poor measures of $\boldsymbol{\eta}_{i}$. In this view, steps 1 and 2 of the analysis become data reduction steps which are used only to develop a calculation that reduces $\mathbf{Y}_{i}$ to these $\tilde{\boldsymbol{\eta}}_{i}$. This also an entirely coherent approach to data analysis, if it is substantively interesting and justified. Here, however, we focus on the situation where models for the latent $\boldsymbol{\eta}_{i}$ are the goal, because it is also what motivates two-step and one-step estimation.
3.4.3 Adjusted three-step estimation
The goal of adjusted three-step estimation is to obtain valid estimates of the structural model $p(\boldsymbol{\eta}_{i}|\mathbf{X}_{i};\boldsymbol{\theta}_{2})$ based on a model for $p(\tilde{\boldsymbol{\eta}}_{i}|\mathbf{X}_{i};\boldsymbol{\theta}_{2})$, treating $\tilde{\boldsymbol{\eta}}_{i}$ as a known function of $\mathbf{Y}_{i}$. Naive three-step estimation is based on the approximation $p(\tilde{\boldsymbol{\eta}}_{i}|\mathbf{X}_{i};\boldsymbol{\theta}_{2})\approx p(\boldsymbol{\eta}_{i}|\mathbf{X}_{i};\boldsymbol{\theta}_{2})$, which is generally incorrect. The correct distribution is of the form
[…]
Note that here we deviate slightly from previous notation, to make one point clearer below. This is that the latent variables are denoted not as $\boldsymbol{\eta}_{i}$, but as $\boldsymbol{\eta}=(\eta_{.1},\dots,\eta_{.K})^{\prime}$ to emphasise the fact that in (14) they are integration variables and not fixed values for units $i$. Adjusted three-step estimation could be based on the right-hand side or left-hand side of (14), and this has been done for other families of latent variable models. For latent trait models, however, this approach is less promising, because it is either redundant or unavailable as a general method.
Consider first the right-hand side of (14). This is still a latent variable model, but with $\tilde{\boldsymbol{\eta}}_{i}$ rather than $\mathbf{Y}_{i}$ as the observed measurement indicators. It is thus like step 2 of two-step estimation, but applied to $\tilde{\boldsymbol{\eta}}_{i}$. To implement it, we would first need to derive the form of the distribution $p(\tilde{\boldsymbol{\eta}}_{i}|\boldsymbol{\eta})$. This is fairly easy for latent class models, where $\mathbf{Y}_{i}$, $\boldsymbol{\eta}$ and $\tilde{\boldsymbol{\eta}}_{i}$ are all categorical, and estimation based on this approach has been developed there by Vermunt (2010) and Bakk et al. (2013). It could also be straightforward for normal structural equation models (SEMs), but it has not been examined in full in that literature (perhaps because other three-step methods discussed below are then more natural and straightforward). Savalei (2018) considered it for SEMs in the special case where $\tilde{\eta}_{ik}$ are equally-weighted sum scores of the measurement items.
To examine the situation for latent trait models, suppose for simplicity that all the items $Y_{ikj}$ are binary and that the measurement model is (7), i.e. that the different traits $\eta_{\cdot k}$ are measured by the distinct sets of items $\mathbf{Y}_{ik}=(Y_{ik1},\dots,Y_{ikp_{k}})^{\prime}$. We can then write (14) as
[…]
Here $\tilde{\eta}_{ik}=\mbox{E}(\eta_{.k}|\mathbf{Y}_{ik})$ are the empirical Bayes predictions, which are not available in a closed form. For purposes of this discussion, however, it is reasonable to approximate them by the weighted sums $\tilde{\eta}_{ik}\approx\sum_{j}a_{kj}Y_{ikj}$, where $a_{kj}$ are functions of the step-1 parameter estimates $\tilde{\boldsymbol{\psi}}_{k}$. As a function of the random vector $\mathbf{Y}_{ik}$, this $\tilde{\eta}_{ik}=\tilde{\eta}_{ik}(\mathbf{Y}_{ik})$ is a discrete random variable with $2^{p_{k}}$ possible values, many of which will have very low probability for a given $\eta_{.k}$, and with mean $\sum_{j}a_{kj}\pi_{ikj}(\eta_{.k})$ and variance $\mbox{var}(\tilde{\eta}_{ik}|\eta_{.k})=\sum_{j}a_{kj}^{2}\pi_{ikj}(\eta_{.k})[1-\pi_{ikj}(\eta_{.k})]$ where $\pi_{ikj}(\eta_{.k})=P(Y_{ikj}=1|\eta_{.k})$ is given by the measurement model (8). The $p(\tilde{\eta}_{ik}|\eta_{.k})$ in (15) is the probability function of this distribution, evaluated at the observed value of $\tilde{\eta}_{ik}$.
Adjusted three-step estimation of this kind has been proposed for latent trait models by Lai and Hsiao (2022). They approximate $p(\tilde{\eta}_{ik}|\eta_{.k})$ by a normal distribution with mean $\eta_{.k}$ (when $\eta_{.k}$ is taken to have variance 1) and variance $\hat{\mbox{var}}(\eta_{.k}|\tilde{\eta}_{ik})$ (which can be obtained from estimation software). This, however, is not correct in general, because $\tilde{\eta}_{ik}$ is not normally distributed and because the variance of $\tilde{\eta}_{ik}$ given $\eta_{.k}$ is not equal to the variance of $\eta_{.k}$ given $\tilde{\eta}_{ik}$ and not constant as a function of $\eta_{.k}$. It is not clear when this approximation would be adequate in practice. It seems that it would require, at least, that $\tilde{\eta}_{ik}$ and $\eta_{.k}$ are scaled to have the same marginal variance and that $\tilde{\eta}_{ik}$ is approximately normally distributed (which requires at least large $p_{k}$) and has roughly the same variance given all the values of $\eta_{.k}$ which make a non-trivial contribution to the integral in (15).
There is, however, a different way of obtaining these three-step estimates. Consider again binary items, and suppose that all estimated measurement models (8) for different items $Y_{ikj}$ for $\eta_{.k}$ are different from each other. The weights $a_{kj}$ in the approximation $\tilde{\eta}_{ik}\approx\sum_{j}a_{kj}Y_{ikj}$ are then also all different, so that every possible value of $\mathbf{Y}_{ik}$ gives a distinct value of $\tilde{\eta}_{ik}=\tilde{\eta}_{ik}(\mathbf{Y}_{ik})$ almost surely. When this is the case, $p(\tilde{\eta}_{ik}|\eta_{.k})=p(\mathbf{Y}_{ik}|\eta_{.k})$ and $p(\tilde{\boldsymbol{\eta}}_{i}|\mathbf{X}_{i};\boldsymbol{\theta}_{2})$ in (15) is equal to the $p(\mathbf{Y}_{i}|\mathbf{X}_{i};\boldsymbol{\theta}_{2})$ that would be used for step 2 of two-step estimation. It would seem that the same argument holds also for the exact values of $\tilde{\boldsymbol{\eta}}_{i}$, and when any items $Y_{ikj}$ are polytomous. In other words, if all the items have different estimated measurement parameters, there is almost surely a one-to-one relationship between distinct values of $\mathbf{Y}_{i}$ and $\tilde{\boldsymbol{\eta}}_{i}$. Adjusted three-step estimation based on the right-hand side of (14), if correctly implemented, would then give the same estimates of $\boldsymbol{\theta}_{2}$ as two-step estimation which uses the same step-1 estimates of $\boldsymbol{\theta}_{1}$ — but would deliver them with more effort for no gain. If different items had identical measurement models (most likely because they were constrained to be so), this kind of three-step estimation would require separate implementation, but this would again be effectively pointless because it would in the end reflect essentially the same information that two-step estimation uses more easily. We do not consider such three-step estimates further in this paper.
Adjusted three-step estimation based on the left-hand side of (14) means estimating $\boldsymbol{\theta}_{2}$ from a closed-form model $p(\tilde{\boldsymbol{\eta}}_{i}|\mathbf{X}_{i};\boldsymbol{\theta}_{2})$ which is obtained by first evaluating the integral in (14). This is again relatively straightforward for latent class models (see Bolck et al. 2004, Vermunt 2010, and Bakk et al. 2013). It is also convenient for linear SEMs, where it means obtaining closed-form expressions for the mean vector and covariance matrix of $\tilde{\boldsymbol{\eta}}_{i}$ and using them to estimate $\boldsymbol{\theta}_{2}$ (see Croon 2002, Devlieger and Rosseel 2017, and references therein).
For latent trait models this would require a sufficiently good approximation of the integral which would also yield a conveniently estimable model. That is not generally available. In some cases, however, a version of it can be derived. Suppose that the interesting part of the structural model is $\eta_{2i}=\beta_{20}+\beta_{21}\eta_{1i}+\psi_{2i}$, with no covariates $\mathbf{X}_{i}$. Suppose further that we use for $\eta_{1i}$ the Empirical Bayes score $\tilde{\eta}_{1i}=\mbox{E}(\eta_{.1}|\mathbf{Y}_{i1})$, and for $\eta_{2i}$ a unbiased score $\breve{\eta}_{2i}$ for which $\mbox{E}(\breve{\eta}_{2i}|\eta_{2i})=\eta_{2i}$. Then we have $\breve{\eta}_{2i}=\beta_{20}+\beta_{21}\tilde{\eta}_{1i}+\psi_{2i}^{*}$, so that $(\beta_{20},\beta_{21})$ can be estimated from a linear regression model for $\tilde{\eta}_{2i}$ given $\tilde{\eta}_{1i}$ (which has a heteroscedastic residual $\psi_{2i}^{*}$, but this can be allowed for in the estimation). For this model and these trait scores, naive and adjusted three-step estimation are thus the same. This situation was described for SEMs by Skrondal and Laake (2001), using regression factor scores as $\tilde{\eta}_{1i}$ and Bartlett factor scores as $\breve{\eta}_{2i}$. Lu and Thomas (2008) showed that the same construction would work also for latent trait models, but they also noted that for these models there are no unbiased scores to be used as $\breve{\eta}_{2i}$ (the best that can be obtained is consistency as the number of items increases).
This approach would not generalise well to more complex models. For example, suppose that the structural model includes a further latent variable $\eta_{3i}$ and the model $\eta_{3i}=\beta_{30}+\beta_{31}\eta_{1i}+\beta_{32}\eta_{2i}+\psi_{3i}$. The scores that should be assigned for $(\eta_{1i},\eta_{2i})$ to estimate this model would be $\mbox{E}[(\eta_{1i},\eta_{2i})|\mathbf{Y}_{1i},\mathbf{Y}_{2i}]$, i.e. predictions given both of their indicators (Skrondal and Laake 2001, Lu and Thomas 2008); these should also be conditional on any covariates $\mathbf{X}_{i}$ in the structural model. To get these, we would need to fit the step-1 model for $\eta_{1i}$ and $\eta_{2i}$ jointly rather than separately. Furthermore, the appropriate trait scores to assign for $\eta_{1i}$ and $\eta_{2i}$ would be different for the model for $\eta_{3i}$ than for the model for $\eta_{2i}$, and different again for any further parts of the structural model. This could quickly get very impracticable for larger models. We do not consider this version of adjusted three-step estimation further in this paper, except for some examples with only a single explanatory latent trait in the simulation studies below.
3.5 Two-step estimation for other types of models
Two-step estimation could in principle be applied to any instance of the general latent variable model that was defined at the start of Section 2. To provide a focus for this paper, we concentrate on the family of latent trait models that was introduced later in Section 2. Before proceeding to empirical investigations of them, we comment briefly on other settings where the two-step method could be used. Some of them are obvious extensions or variants of the models considered here, while others could involve more consequential differences. Implementation and performance of two-step estimation in many of these other cases remain to be examined.
In the measurement model we made the assumption that $p(\mathbf{Y}_{i}|\boldsymbol{\eta}_{i},\mathbf{X}_{i};\boldsymbol{\theta}_{1})=p(\mathbf{Y}_{i}|\boldsymbol{\eta}_{i};\boldsymbol{\theta}_{1})$. Relaxing this would take us to models with non-equivalence of measurement with respect to the observed covariates $\mathbf{X}_{i}$. Two-step estimation in that situation can introduce new issues, and deserves further investigation. For latent class models, it has been studied by Di Mari and Bakk (2018), Janssen et al. (2019) and Vermunt and Magidson (2021a).
We made the common assumption that all items in $\mathbf{Y}_{i}$ are conditionally independent given the latent traits $\boldsymbol{\eta}_{i}$. This could be relaxed by allowing conditional associations between some items, requiring only corresponding modifications of the log likelihoods (9) and (10). Another obvious class of extensions would be obtained by considering models with mixtures of different types of measurement items. In this paper we focus on categorical items, while linear structural equation models (as in Rosseel and Loh 2022) take the items to be continuous. But some applications may involve both of these types, or others such as counts or survival times as measurement items. Similarly, the structural model could include both categorical and continuous latent variables, and categorical variables could be measured not only by categorical items (as in standard latent class analysis discussed by Bakk and Kuha 2018) but also by continuous ones (the case known as latent profile analysis). It seems plausible that the behaviour of the two-step method would not be fundamentally different in these different combinations of settings.
Generalisations of the structural model could also include such possibilities as interactions or nonlinear terms of the latent variables, or non-recursive models. Two-step estimation in these cases has not been examined. It could behave differently there, if only because such models are inherently more complex and difficult howsoever they are estimated.
One extension that is easily accommodated is multigroup modelling. This means models for data on units from different groups, such as respondents from different countries in cross-national research. The focus of interest is then on how the parameters of the structural model do or do not vary between the groups. The simplest version of this is to include dummy variables for the groups in $\mathbf{X}_{i}$, thus giving separate intercepts for different groups in the structural models (3)–(4). Beyond this, latent variable modelling software such as Mplus also have dedicated multigroup syntax which allows any of the parameters to vary by the group. This can also be combined with two-step estimation. In the applied example in Section 5 we use it to fit a model to cross-sectional survey data from 36 countries, where all the regression coefficients and all residual variances and covariances in the structural model vary between the countries.
4 Simulations
4.1 First set of simulations
We have used simulation studies to examine the properties of two-step estimates for the types of models considered in this paper. In this section we describe them for one model setting in detail. In Section 4.2 we then describe further simulations in several other settings.
The properties that we examine include the performance of the point estimates and accuracy of the estimated standard errors. The main comparison of interest is with one-step estimates of the same models. Since they are full maximum likelihood estimates, they serve as a benchmark whose theoretical properties are well understood. In large samples one-step estimates should be unbiased and fully efficient, and it is then interesting to see if two-step estimates do roughly as well. In smaller samples, either approach could have an advantage. In this section all the models are correctly specified, but in the next we also include simulations with a misspecified measurement model in order to examine the robustness of the estimates in such cases. Another point of comparison is computational demands, where two-step estimation should have the advantage. The simulation results also include (naive) three-step estimates, but they are not discussed here. We will comment on the results for them separately in Section 4.3.
In these first simulations the structural model for unit $i=1,\dots,n$ is of the form
$\eta_{i2}=\beta_{0}+\beta_{1}\eta_{i1}+\epsilon_{i}$ (16)
where $\eta_{1i}$ and $\eta_{2i}$ are latent variables and $\epsilon_{i}$ is the model residual. Here $\eta_{1i}\sim N(0,1)$ and $\epsilon_{i}\sim N(0,\sigma^{2}_{\epsilon})$, $\beta_{0}=0$, and $(\beta_{1},\sigma^{2}_{\epsilon})$ are such that marginally $\eta_{2i}\sim N(0,1)$ in all simulation settings. The value of $\beta_{1}$ is set so that the $R^{2}$ statistic for the response $\eta_{i2}$ is $R^{2}=0,0.2$, or $0.4$. We refer to this as simulation case A. Path diagram for this model is shown in Figure A.1 in the supplementary appendix.
The latent variable $\eta_{i1}$ is measured by $p$ binary items $Y_{i1j}$, coded with values 0 and 1, and $\eta_{i2}$ by a different set of $p$ binary items similarly, with $p=4$ or $p=8$. All items are conditionally independent of each other given the latent variables. The measurement model is of the form $\text{logit}[P(Y_{ikj}=1|\eta_{ik})]=\tau_{kj}+\lambda_{kj}\eta_{ik}$ for $k=1,2$ and $j=1,\dots,p$. The measurement parameters $(\tau_{kj},\lambda_{kj})$ are estimated as distinct parameters for different $k,j$, but their true values $\tau_{kj}=\tau$ and $\lambda_{kj}=\lambda$ in the data-generating model are set to be equal for all items in a given simulation. The loading parameter $\lambda$ is set with reference to the linear model for a notional continuous latent variable $Y_{ikj}^{*}$ which implies the logistic model for $Y_{ikj}$. Here $R^{2}$ for $Y_{ikj}^{*}$ given $\eta_{ik}$ is $R^{2}_{Y}=\lambda^{2}\mbox{var}(\eta_{ik})/(\lambda^{2}\mbox{var}(\eta_{ik})+\pi^{2}/3)$ where $\mbox{var}(\eta_{ik})=1$, and $\lambda$ is set so that $R^{2}_{Y}=0.4$ or 0.6. The values of $\tau$ are then set so that the marginal probability $\pi_{Y}=P(Y_{ikj}=1)$ is 0.5 or (approximately) 0.8. When the model is estimated, we constrain $\tau_{k1}=0$ and $\lambda_{k1}=1$ for the first item $Y_{ik1}$ of each latent variable, leaving $\tau_{kj}$ and $\lambda_{kj}$ for the other $p-1$ items estimable. This affects the implied scales of the latent variables, so that the value of $\beta_{1}$ that the estimators should actually be estimating will depend also on the value of $\lambda$ in the data-generating model. These true values of $\beta_{1}$ are shown in the results tables.
In each setting we generate 1000 simulated datasets with $n=200$ or $n=1000$ independent observations $i$. All combinations of $n$, $R^{2}$, $p$, $\pi_{Y}$ and $R^{2}_{Y}$ are considered, resulting in 48 settings. The two-step, one-step and naive three-step estimates are calculated. Two-step estimation was implemented as described in Section 3.3. The same R functions we wrote for it (combined with Mplus) also produced the one-step and three-step estimates.
We focus on the results for the estimates of the regression coefficient $\beta_{1}$ in the structural model (simulation results for the measurement model are discussed briefly in Section 4.2). These results are shown in Table 1 for the point estimates and Table 2 for the standard errors. For the point estimates, the table includes their mean bias, root mean squared error (RMSE) and median absolute error (MAE). To examine the estimated standard errors, we report the standard deviation of the point estimates of $\beta_{1}$ and the mean of their estimated standard errors across the simulations, and simulation coverage of the 95% confidence intervals. For the two-step estimates we also report two quantities for the case where we ignore the uncertainty from step 1 (i.e.  omit the contribution from $\mathbf{V}_{1}$ to (12)) and include only the uncertainty from step 2 in the standard errors: the average proportion of the full estimated variance of the two-step estimate of $\beta_{1}$ that this accounts for, and the coverage of the resulting 95% confidence interval.
Consider first the point estimates of $\beta_{1}$ in Table 1. There is no meaningful difference between two-step and one-step estimates for the larger sample size of $n=1000$. They also behave very similarly in almost all settings when $n=200$, except for some differences when the number of items is small ($p=4$), the measurement model is weak ($R^{2}_{Y}=0.4$) and one of the values of the items is rare ($\pi_{Y}=0.8$). This is the setting where the items provide the least information about the latent variables. When $\beta_{1}$ is 0, the two-step estimates have in these cases a slightly better RMSE and MAE than the one-step estimates. With larger values of $\beta_{1}$, on the other hand, in a few simulations these cases yield an extreme value of the two-step estimate, inflating its RMSE. These extremes are reduced for the one-step estimates, most likely because the measurement models of $\eta_{1}$ and $\eta_{2}$ stabilise each other when they are estimated together.
Results for the estimated standard errors of the estimates of $\beta_{1}$ are reported in Table 2. They are very satisfactory when $n=1000$: for both the two-step and one-step methods the standard errors are good estimates of the simulation standard deviations of estimates of $\beta_{1}$, and coverages of confidence intervals are close to the nominal 95%. With the smaller sample size of $n=200$ this coverage is less correct, ranging from 89% to 100% in different settings. The estimated standard errors are even then reasonably accurate when the estimated measurement model is sufficiently informative. In the most difficult settings, with a small sample, weak measurement model and small number of items, the standard errors tend to underestimate the true variability.
The last two columns of Table 2 examine what happens if we ignore the uncertainty from step 1 of two-step estimation, i.e. if we calculate its standard errors using only the term $\mathbf{V}_{2}$ in the variance formula (12). Here this step-2 variance accounts for almost all of the variance of the estimate of $\beta_{1}$ when the true $\beta_{1}$ is 0, but much less otherwise, down to less than a third of the overall variance in the settings with the largest $\beta_{1}$. The coverage of confidence intervals is correspondingly reduced, down to less than 70%. When the association in the structural model is strong, the uncertainty in its estimates is thus dominated by the uncertainty in the estimated measurement parameters from step 1.
Table 1: Simulation results for point estimates of structural regression coefficient $\beta_{1}$ of a latent covariate $\eta_{1}$ for a conditionally normally distributed latent response $\eta_{2}$ (simulation case A).
Bias RMSE MAE
$p$ $\pi_{Y}$ $R^{2}_{Y}$ $R^{2}_{\eta}$ $\beta_{1}$ 2-step 1-step 3-step 2-step 1-step 3-step 2-step 1-step 3-step
$n=200$
4 0.5 0.4 0.0 0.000 -0.004 -0.003 -0.002 0.133 0.141 0.081 0.077 0.080 0.046
4 0.8 0.4 0.0 0.000 -0.007 -0.004 -0.000 0.240 0.276 0.110 0.101 0.113 0.044
4 0.5 0.6 0.0 0.000 0.007 0.007 0.005 0.115 0.117 0.080 0.069 0.070 0.048
4 0.8 0.6 0.0 0.000 0.002 0.002 0.003 0.153 0.154 0.091 0.081 0.084 0.048
8 0.5 0.4 0.0 0.000 -0.001 -0.001 -0.001 0.104 0.104 0.078 0.063 0.062 0.047
8 0.8 0.4 0.0 0.000 0.008 0.009 0.006 0.147 0.151 0.087 0.081 0.083 0.047
8 0.5 0.6 0.0 0.000 -0.003 -0.003 -0.002 0.097 0.095 0.077 0.062 0.059 0.050
8 0.8 0.6 0.0 0.000 -0.002 -0.002 -0.000 0.114 0.116 0.080 0.072 0.073 0.050
4 0.5 0.4 0.2 0.447 0.016 0.035 -0.162 0.236 0.237 0.221 0.132 0.132 0.202
4 0.8 0.4 0.2 0.447 0.031 0.070 -0.208 0.404 0.385 0.300 0.183 0.177 0.267
4 0.5 0.6 0.2 0.447 0.019 0.025 -0.116 0.182 0.189 0.176 0.117 0.116 0.148
4 0.8 0.6 0.2 0.447 0.030 0.039 -0.144 0.257 0.244 0.225 0.135 0.134 0.183
8 0.5 0.4 0.2 0.447 0.013 0.010 -0.104 0.177 0.188 0.170 0.109 0.110 0.137
8 0.8 0.4 0.2 0.447 0.023 0.040 -0.152 0.249 0.250 0.222 0.144 0.136 0.197
8 0.5 0.6 0.2 0.447 0.005 -0.008 -0.081 0.143 0.159 0.143 0.092 0.093 0.116
8 0.8 0.6 0.2 0.447 0.024 0.028 -0.100 0.183 0.181 0.170 0.114 0.111 0.137
4 0.5 0.4 0.4 0.632 0.028 0.047 -0.223 0.309 0.285 0.301 0.176 0.162 0.268
4 0.8 0.4 0.4 0.632 0.042 0.081 -0.284 0.608 0.457 0.436 0.228 0.199 0.363
4 0.5 0.6 0.4 0.632 0.048 0.052 -0.143 0.275 0.254 0.247 0.140 0.147 0.199
4 0.8 0.6 0.4 0.632 0.038 0.048 -0.195 0.326 0.283 0.295 0.174 0.161 0.255
8 0.5 0.4 0.4 0.632 0.020 0.022 -0.142 0.213 0.222 0.216 0.131 0.134 0.176
8 0.8 0.4 0.4 0.632 0.033 0.054 -0.204 0.344 0.346 0.309 0.177 0.166 0.258
8 0.5 0.6 0.4 0.632 0.033 0.006 -0.087 0.196 0.227 0.182 0.127 0.124 0.138
8 0.8 0.6 0.4 0.632 0.024 0.030 -0.138 0.223 0.218 0.220 0.137 0.135 0.180
$n=1000$
4 0.5 0.4 0.0 0.000 0.000 0.000 0.000 0.053 0.054 0.032 0.036 0.036 0.021
4 0.8 0.4 0.0 0.000 -0.001 -0.001 -0.000 0.072 0.073 0.032 0.049 0.049 0.021
4 0.5 0.6 0.0 0.000 -0.001 -0.001 -0.001 0.045 0.045 0.031 0.030 0.030 0.021
4 0.8 0.6 0.0 0.000 0.001 0.001 0.001 0.054 0.055 0.032 0.036 0.036 0.021
8 0.5 0.4 0.0 0.000 0.000 0.000 0.000 0.045 0.045 0.033 0.031 0.031 0.023
8 0.8 0.4 0.0 0.000 0.001 0.001 0.001 0.059 0.059 0.034 0.038 0.038 0.022
8 0.5 0.6 0.0 0.000 -0.001 -0.001 -0.000 0.040 0.040 0.032 0.026 0.026 0.021
8 0.8 0.6 0.0 0.000 0.000 0.000 0.000 0.045 0.045 0.032 0.030 0.030 0.021
4 0.5 0.4 0.2 0.447 0.006 0.012 -0.174 0.092 0.090 0.183 0.061 0.059 0.178
4 0.8 0.4 0.2 0.447 0.011 0.021 -0.227 0.124 0.124 0.236 0.078 0.077 0.234
4 0.5 0.6 0.2 0.447 0.001 0.003 -0.131 0.075 0.074 0.142 0.053 0.052 0.135
4 0.8 0.6 0.2 0.447 0.011 0.013 -0.160 0.091 0.090 0.171 0.057 0.059 0.168
8 0.5 0.4 0.2 0.447 0.004 0.005 -0.112 0.073 0.073 0.125 0.047 0.047 0.116
8 0.8 0.4 0.2 0.447 0.010 0.015 -0.163 0.094 0.093 0.174 0.062 0.061 0.166
8 0.5 0.6 0.2 0.447 0.002 0.003 -0.084 0.062 0.062 0.098 0.039 0.039 0.087
8 0.8 0.6 0.2 0.447 0.008 0.009 -0.113 0.074 0.074 0.126 0.048 0.048 0.118
4 0.5 0.4 0.4 0.632 0.009 0.014 -0.244 0.118 0.112 0.255 0.076 0.076 0.250
4 0.8 0.4 0.4 0.632 0.003 0.014 -0.319 0.156 0.150 0.329 0.103 0.099 0.329
4 0.5 0.6 0.4 0.632 0.006 0.007 -0.177 0.093 0.090 0.190 0.061 0.058 0.182
4 0.8 0.6 0.4 0.632 0.009 0.010 -0.218 0.117 0.109 0.232 0.074 0.071 0.223
8 0.5 0.4 0.4 0.632 0.003 0.005 -0.158 0.088 0.086 0.172 0.059 0.057 0.164
8 0.8 0.4 0.4 0.632 -0.001 0.007 -0.231 0.114 0.111 0.244 0.078 0.072 0.237
8 0.5 0.6 0.4 0.632 0.007 0.008 -0.110 0.080 0.079 0.128 0.054 0.053 0.112
8 0.8 0.6 0.4 0.632 0.006 0.008 -0.152 0.088 0.086 0.167 0.060 0.058 0.156
Note: $p$ denotes number of measurement items $Y_{j}$, $\pi_{Y}$ marginal proportion of $Y_{j}=1$,
         $R_{Y}^{2}$ and $R^{2}_{\eta}$ the $R^{2}$ statistics in models for $Y_{j}$ and $\eta_{2}$, and $\beta_{1}$ the true value of $\beta_{1}$.
Table 2: Simulation results for standard error estimates of estimated structural regression coefficient $\beta_{1}$ of a latent covariate $\eta_{1}$ for a conditionally normally distributed latent response $\eta_{2}$ (simulation case A).
Simulation standard deviation 2-step
vs. mean of est. std. error Coverage of without
2-step 1-step 95% conf. interval step-1 var.
$p$ $\pi_{Y}$ $R^{2}_{Y}$ $R^{2}_{\eta}$ $\beta_{1}$ s.d. m(se) s.d. m(se) 2-st. 1-st. 3-st. %var. cover.
$n=200$
4 0.5 0.4 0.0 0.000 0.133 0.137 0.141 0.138 99.5 99.3 95.6 87.3 96.0
4 0.8 0.4 0.0 0.000 0.240 0.224 0.276 0.235 99.9 100.0 94.7 78.0 96.0
4 0.5 0.6 0.0 0.000 0.115 0.114 0.116 0.114 98.9 98.5 94.3 91.4 94.9
4 0.8 0.6 0.0 0.000 0.153 0.146 0.154 0.145 99.2 98.6 94.2 88.2 95.1
8 0.5 0.4 0.0 0.000 0.105 0.105 0.105 0.102 98.5 98.5 95.7 92.5 95.8
8 0.8 0.4 0.0 0.000 0.147 0.143 0.151 0.143 99.3 99.0 95.0 87.1 95.9
8 0.5 0.6 0.0 0.000 0.097 0.095 0.095 0.090 98.4 98.1 95.5 94.3 95.7
8 0.8 0.6 0.0 0.000 0.115 0.112 0.116 0.112 99.1 99.0 94.5 92.1 95.0
4 0.5 0.4 0.2 0.447 0.236 0.223 0.235 0.217 90.5 90.5 36.8 41.3 77.9
4 0.8 0.4 0.2 0.447 0.403 0.333 0.378 0.331 88.5 90.3 28.2 41.6 75.4
4 0.5 0.6 0.2 0.447 0.181 0.183 0.187 0.178 92.8 92.7 48.6 43.3 82.8
4 0.8 0.6 0.2 0.447 0.256 0.230 0.241 0.219 91.5 91.9 38.4 42.8 79.3
8 0.5 0.4 0.2 0.447 0.176 0.162 0.188 0.160 90.7 89.3 48.7 39.7 77.0
8 0.8 0.4 0.2 0.447 0.248 0.226 0.247 0.226 91.1 91.8 36.5 39.7 75.1
8 0.5 0.6 0.2 0.447 0.143 0.143 0.159 0.137 92.1 90.0 57.4 42.3 78.9
8 0.8 0.6 0.2 0.447 0.181 0.175 0.179 0.171 92.1 92.6 51.4 40.9 78.5
4 0.5 0.4 0.4 0.632 0.308 0.295 0.281 0.263 91.9 92.7 28.6 27.3 72.2
4 0.8 0.4 0.4 0.632 0.607 0.439 0.450 0.385 89.7 92.8 19.9 27.9 69.6
4 0.5 0.6 0.4 0.632 0.271 0.254 0.248 0.229 94.0 94.6 36.7 28.8 73.8
4 0.8 0.6 0.4 0.632 0.324 0.316 0.279 0.267 93.2 94.5 29.5 28.3 71.2
8 0.5 0.4 0.4 0.632 0.212 0.208 0.221 0.203 92.5 92.5 40.7 24.9 66.1
8 0.8 0.4 0.4 0.632 0.342 0.284 0.342 0.280 91.0 92.7 27.6 24.6 65.1
8 0.5 0.6 0.4 0.632 0.193 0.189 0.227 0.175 94.6 90.2 47.7 25.9 66.8
8 0.8 0.6 0.4 0.632 0.221 0.218 0.216 0.211 93.2 94.1 38.4 26.2 68.5
$n=1000$
4 0.5 0.4 0.0 0.000 0.053 0.054 0.054 0.054 95.9 95.7 95.1 97.1 95.1
4 0.8 0.4 0.0 0.000 0.072 0.076 0.073 0.076 97.9 97.2 96.3 95.1 96.3
4 0.5 0.6 0.0 0.000 0.045 0.046 0.045 0.046 96.5 96.4 95.2 98.3 95.4
4 0.8 0.6 0.0 0.000 0.054 0.056 0.055 0.056 96.8 96.8 95.9 97.6 96.2
8 0.5 0.4 0.0 0.000 0.045 0.044 0.045 0.044 95.3 95.2 94.5 98.2 94.5
8 0.8 0.4 0.0 0.000 0.059 0.056 0.059 0.056 95.0 94.8 93.7 96.9 94.0
8 0.5 0.6 0.0 0.000 0.040 0.040 0.040 0.040 96.1 95.9 95.2 98.8 95.2
8 0.8 0.6 0.0 0.000 0.045 0.046 0.046 0.046 96.3 96.2 95.1 98.4 95.3
4 0.5 0.4 0.2 0.447 0.092 0.091 0.089 0.089 94.4 95.4 5.9 41.6 79.5
4 0.8 0.4 0.2 0.447 0.124 0.123 0.122 0.120 93.6 94.1 3.2 42.6 80.2
4 0.5 0.6 0.2 0.447 0.075 0.076 0.074 0.074 94.8 94.6 12.6 43.8 79.3
4 0.8 0.6 0.2 0.447 0.090 0.090 0.089 0.089 96.0 95.8 8.4 43.8 81.3
8 0.5 0.4 0.2 0.447 0.073 0.070 0.073 0.069 93.9 94.6 18.8 39.3 77.7
8 0.8 0.4 0.2 0.447 0.093 0.091 0.092 0.091 93.9 94.6 8.5 38.7 76.6
8 0.5 0.6 0.2 0.447 0.062 0.062 0.062 0.062 94.6 94.7 28.6 41.9 79.2
8 0.8 0.6 0.2 0.447 0.074 0.073 0.073 0.072 95.1 94.8 17.9 41.4 80.1
4 0.5 0.4 0.4 0.632 0.118 0.118 0.112 0.109 94.3 94.9 2.3 28.9 70.9
4 0.8 0.4 0.4 0.632 0.156 0.155 0.149 0.143 93.4 93.8 1.6 29.1 70.4
4 0.5 0.6 0.4 0.632 0.093 0.097 0.089 0.092 95.7 95.9 6.8 30.5 73.9
4 0.8 0.6 0.4 0.632 0.116 0.115 0.109 0.108 94.7 94.7 4.0 30.4 73.9
8 0.5 0.4 0.4 0.632 0.088 0.089 0.086 0.087 95.5 96.2 7.9 24.6 68.0
8 0.8 0.4 0.4 0.632 0.114 0.114 0.110 0.111 93.3 94.5 2.6 24.5 65.3
8 0.5 0.6 0.4 0.632 0.080 0.080 0.078 0.077 95.1 94.6 19.8 26.3 69.2
8 0.8 0.6 0.4 0.632 0.088 0.092 0.086 0.089 95.5 95.8 10.5 26.3 70.8
Note: $p$ denotes number of measurement items $Y_{j}$, $\pi_{Y}$ marginal proportion of $Y_{j}=1$,
         $R_{Y}^{2}$ and $R^{2}_{\eta}$ the $R^{2}$ statistics in models for $Y_{j}$ and $\eta_{2}$, and $\beta_{1}$ the true value of $\beta_{1}$.
$\dagger$ Average % of the variance of 2-step estimator accounted for by step-2 variance alone,
       and coverage of 95% confidence interval if only this variance is included.
4.2 Further simulations
Here we describe some further simulations for different settings. Path diagrams for these models are shown in Figures A.1–A.2, and results in Tables A.1–A.25, in the supplementary appendix. The discussion is short, because we focus on results that extend or deviate from the conclusions from the first set of simulations in Section 4.1. The following situations are included here:
  • Observed covariate, latent response (simulation case B). Structural model is as in case A (equation 16), except that $\eta_{i1}$ is replaced by an observed covariate $X_{i}$. Three distributions for $X_{i}$ are used: $X_{i}\sim N(0,1)$ [case B1], $X_{i}\sim 2\,\text{Bernoulli}(0.5)-1$ [B2], and $X_{i}\sim\text{Gamma}(\alpha,\beta)$ with shape $\alpha=0.06$ and rate $\beta=1$, scaled to have sample mean 0 and variance 1 [B3].
  • Latent covariate, observed response (case C). Structural model is as in case A, except that $\eta_{i2}$ is replaced by an observed response which is here labelled $Z_{i}$. Two conditional distributions for $Z_{i}$ given $\eta_{1i}$ (i.e. distributions for $\epsilon_{i}$) are considered, a normal distribution [case C1] and a skew-normal distribution (Azzalini and Capitanio, 2014) set to have a skewness coefficient of just under 1 [C2], with $Z_{i}$ having marginal mean of 0 and marginal variance 1 in all settings.
  • Ordinal measurement items (case D). Structural model is the same as in case A, but each measurement item $Y_{ikj}$ has 4 rather than 2 categories. Measurement models are again derived from a linear model for a latent continuous $Y_{ijk}^{*}$, here resulting in an ordinal logistic measurement model for each $Y_{ijk}$. Its loading parameter $\lambda$ is set as in case A. The measurement intercepts are such that the marginal probabilities of the four categories of each $Y_{ikj}$ are (0.25,0.25,0.25,0.25) when $R^{2}_{Y}=0.6$ and about (0.20,0.30,0.30,0.20) when $R^{2}_{Y}=0.4$ (these correspond to the setting $\pi_{Y}=0.5$ in case A).
  • Three latent variables (case E). The structural model has two parts, $\eta_{2i}=\alpha_{0}+\alpha_{1}\eta_{1i}+\epsilon_{2i}$ and $\eta_{3i}=\beta_{0}+\beta_{1}\eta_{1i}+\beta_{2}\eta_{2i}+\epsilon_{3i}$ for a further latent variable $\eta_{3i}$. The latent variables are each marginally distributed as $N(0,1)$, and each measured by $p$ separate binary items $Y_{ikj}$. We set $R_{Y}^{2}=0.4$ for both $\eta_{2i}$ and $\eta_{3i}$. What we vary here is the relative sizes of the coefficients $\beta_{1}$ and $\beta_{2}$, so that the proportion of the “indirect effect” of $\eta_{1i}$ on $\eta_{3i}$ via $\eta_{2i}$ ($\alpha_{1}\beta_{2}$) out of the “total effect” ($\beta_{1}+\alpha_{1}\beta_{2}$) is 0, 1/3 or 1/2. Here and for cases F and G we report simulation results for estimates of $\beta_{2}$.
  • Measurement model with cross-loadings (case F). Structural model is the same as in case E. The difference is that the items $Y_{i23}$ and $Y_{i24}$, which were previously measures of $\eta_{2i}$ only, are here also measures of $\eta_{1i}$. The measurement loadings for $Y_{i23}$ and $Y_{i24}$ given $\eta_{1i}$ have the same value $\lambda$ as all other loadings in the measurement model. In step 1 of two-step estimation the measurement model for $\eta_{1}$ and $\eta_{2}$ was accordingly estimated for these variables together, including the cross-loadings.
  • Misspecified measurement model(case G). The data-generating model is as in case F, but the measurement model is estimated as for case E, i.e. incorrectly omitting the cross-loadings from $\eta_{1i}$ to $Y_{i23}$ and $Y_{i24}$.
Cases B and C include the same 48 parameter settings as in case A, but cases D–F omit some of them for simplicity. There we consider only settings with the smaller number of $p=4$ measurement items for each latent variable, and balanced items probabilities with $\pi_{Y}=0.5$. What is then left to vary are the strength of the measurement models (with $R^{2}_{Y}=0.4,0.6$), strengths of the associations in the structural model (with $R^{2}_{\eta}=0,0.2,0.4$ in case D, and as described above for cases E–G), and sample sizes ($n=200,1000$). In case E we also included $n=100$ to further examine the performance of the estimators in such small samples. The results are again based on 1000 simulated datasets in each setting.
Consider first some of the results for parameters of the measurement models. Supplementary Table A.1 shows simulation means of the estimated measurement loadings ($\lambda_{kj}$) and table A.2 means of the estimated measurement intercepts ($\tau_{kj}$) from one-step estimation and step 1 of two-step estimation in simulation cases A, B2 and B3. These means are over both the 1000 simulations and all $p$ or (in case A) $2p$ loadings or intercepts in the model. The true values of all loadings and intercepts are 1 and 0 respectively in each case. In case A these are results for maximum likelihood (ML) estimates of these parameters for correctly specified models, two one-trait models (for two-step estimation) or one two-trait model (for one-step estimation). They are, as expected, approximately unbiased when $n=1000$. With the smaller sample size of $n=200$, however, the estimates show some bias. The simulation medians (which are not shown) are even then very close to the true values, so the observed bias is entirely due to skewness of the sampling distributions of these estimates in small samples.
Measurement models in cases B2 and B3 are of interest because they are examples of the situation (as discussed at the end of Section 3.2) where step 1 of two-step estimation is distributionally misspecified because a non-normal covariate $X$ induces a non-normal marginal distribution of a latent variable (here $\eta_{2}$). In contrast, the model for one-step estimation (and two-step when $R^{2}_{\eta}=0$) is correctly specified. Previous literature on misspecified latent trait models suggests that this should be a minor deviation in case B2, because the Bernoulli distribution of $X$ implies a symmetric two-component mixture model for $\eta_{2}$. Case B3 could be more difficult, because there the marginal distribution of $\eta_{2}$ is skewed. It has a skewness coefficient of about 0.7 when $R^{2}_{\eta}=0.2$ and about 2.0 (i.e. comparable to an exponential distribution) when $R^{2}_{\eta}=0.4$. However, the results do not show any new problems in either of these cases: two-step estimates still behave similarly with the one-step estimates. This is the case even though case B3 was designed to be so extreme that it should not really occur in practice: the observed $X$ had a skewness coefficient of 8.2, which any data analyst would be likely to notice and address by taking a (log) transformation of $X$. It thus seems that it is difficult for a non-normal covariate to induce a serious bias in the estimated measurement parameters from two-step estimation. It should then be even more difficult for this to lead to bias in the estimated structural parameters from step 2 — which are the focus of interest — and this is indeed what we find below.
Consider then results for estimates of structural regression coefficients. First, in cases B1, C1, D, E and F (and B2 and B3 when $R^{2}_{\eta}=0$) the estimation model is correctly specified. These results (and those for the cases discussed below) are reported in supplementary Tables A.3–A.25. The broad conclusion is that most of these results are similar to the results in case A. Both estimates perform mostly well, and two-step estimates at least as well as one-step estimates. These conclusions are thus not substantially affected when the models are larger than in case A, in having ordinal items, more latent variables, or non-simple measurement models. This is unsurprising, given that the large-sample properties of the estimators (as discussed in Sections 3.2 and 3.4.1) should not be affected by the complexity of the model.
There are, however, some new findings in these cases. First, in case C1 the estimates perform very similarly even in the most difficult settings (where the two-step ones had some extreme values in case A). A possible explanation is that when the latent $\eta_{1}$ is only an explanatory variable, a weak estimated measurement model for it tends to attenuate rather than exaggerate the estimate of $\beta_{1}$, preventing occasional extreme estimates. Second, in cases with three latent variables (E and F) there are several settings where the one-step estimates have a much larger bias and RMSE than the two-step estimates. This is due to occasional extreme values of the one-step estimates, as can be seen by observing that the MAEs and simulation medians (the latter are shown separately in Tables A.18 and A.21) are similar even in these settings. These extremes occur most prominently with the largest sample size ($n=1000$), which shows that they are not due to poor small-sample behaviour. Instead, they indicate numerical difficulties with one-step estimation, and of the estimation algorithm failing to converge to (or near) the maximum likelihood estimate. The code for the simulation was set to increase the number of starting values for datasets where the initial settings were not enough, but here the procedure we used for this was not always sufficient. Thus we could probably have removed some of these extreme estimates by manually increasing the starting values further. However, we left them in the results because they illustrate another difference between the estimators, which is that the two-step approach is computationally substantially more manageable than the one-step one. For the two-step estimates there were also some numerical extreme values in cases E (with $n=100$) and F, but for the estimated standard errors rather than point estimates of $\beta_{2}$.
Even when one-step estimation converges without problems, it is more time-consuming; most of the computing time in these simulations was spent on it. It is likely that this is the most important practical difference between these two methods in large models, where one-step estimation quickly becomes more and more demanding.
In case E we also included a small sample size of $n=100$, where asymptotic properties of the estimators are least appropriate. This would arguably be uncomfortably small for substantive applications of models of this type. Even here the estimators perform mostly similarly and mostly well, although inevitably less well than with larger samples, and again with some extreme estimates. The two-step estimates have consistently the smaller RMSEs and MAEs, suggesting that they may have somewhat better properties in very small samples.
Consider then cases where the estimation model is misspecified in some way (B2, B3, C2, and G). There is essentially no difference between C2 and C1, suggesting that misspecification of the distribution of an observed response $Z$ is inconsequential. In cases B2 and B3 there is misspecification in step 1 of two-step estimation, because the distribution of the covariate $X$ is not normal. As discusssed above, estimated measurement models were nevertheless not seriously affected. It follows that two-step estimates of the structural regression coefficient are also unaffected. They still behave similarly to one-step estimates, for which this model is correctly specified. In case B2 the results are essentially similar to the correctly specified B1. In case B3 the estimates are noticeably biased, at least with $n=200$. This, however, affects both estimates — and is indeed worse for one-step estimates — so it is not due to the distributional misspecification in two-step estimation. Instead, it must be caused by the extreme skeweness of $X$, which hurts the small-sample performance of both estimates.
Finally, in case F the estimation model is misspecified for both estimators. Both of them give biased estimates of all the structural coefficients in this case, with $\beta_{2}$ being overestimated. The bias is noticeably smaller for two-step estimators, but still substantial. This is not surprising. Two-step estimation may reduce bias in estimated parameters of the measurement models because of the way it estimates them separately, but it still uses the same misspecifified form of the model (i.e. of which items measure which latent variables). So we would not in general expect it to offer much additional protection against bias from model misspecification.
The conclusions from these simulations are broadly similar to previous ones for other types of models by Bakk and Kuha (2018) and Rosseel and Loh (2022). In all of these contexts two-step and one-step estimates mostly perform very similarly. Differences between them emerge in difficult situations where the measurement models are weak and sample sizes are small. Here there is some variation between the findings of the different studies. In the latent class examples of Bakk and Kuha (2018), two-step estimation does relatively poorly in difficult settings, possibly because an observed covariate or response provides particularly useful stabilisation for one-step estimates there. In contrast, in the simulations of structural equation models by Rosseel and Loh (2022) it is one-step estimation that performs much less well in the most difficult settings. In our simulations the differences in the hardest situations were somewhat smaller than in these previous studies. It should be noted that the exact simulation settings are not easily comparable. For example, Rosseel and Loh (2022) considered models with even more latent factors (five rather than three), which may be particularly challenging for the small-sample performance of one-step estimation (it would also be computationally very demanding for it for latent trait models). They also considered various cases where the analysis models were misspecified. The conclusions were qualitatively the same as in our simulations, i.e. that both estimates were then biased but the two-step estimates often somewhat less so.
4.3 Comments on three-step estimates in the simulations
The simulation tables include also results for naive three-step estimates which replace the latent variables $\eta_{ik}$ with their empirical Bayes predictions $\tilde{\eta}_{ik}$. They are typically badly biased for the true structural parameters. This to be expected, based on the general results discussed in Sections 3.4.2 and 3.4.3. These naive estimates are clearly an inadequate way of estimating the structural model for $\eta_{ik}$. As discussed in Section 3.4.2, it is more constructive to regard them as estimates for a model for $\tilde{\eta}_{ik}$, treated as a model of interest in its own right.
An exception here is case C, where the only latent variable $\eta_{1}$ appears as the only explanatory variable in the model. As observed in Section 3.4.3, naive three-step estimation which uses the empirical Bayes predictions then becomes a form of adjusted three-step estimation which avoids measurement error bias. Accordingly, estimates of $\beta_{1}$ in cases C1 and C2 are essentially unbiased and very similar to the two-step and one-step estimates (see supplementary Tables A.11 and A.13). This is not the case for the other simulations here, and generalising this type of adjustment to them would not be as easy.
[…]
5 Application: Extrinsic and intrinsic work values
Here we apply the methods to a real-data example. We use it, in particular, to further explore two topics which were not the focus of the simulations in Section 4: the possibility of interpretative confounding and computational demands of the different estimators.
The example concerns individuals’ work value orientations, conceptualised in two dimensions: intrinsic and extrinsic work values (for more information about substantive research in this area, see for example Gesthuizen et al. (2019) and Bacher et al. (2022), and references cited therein). We use data from the fifth wave of the European Values Study (EVS), conducted in 2017–20 (EVS, 2022). The survey includes six items on work values, introduced (in the English-language version) by Here are some aspects of a job that people say are important. Please look at them and tell me which ones you personally think are important in a job? The responses are binary, coded as 1 if the respondent mentioned that aspect and 0 if they did not (and with “Don’t know” responses coded as missing). Three of the items are treated as measures of extrinsic work values — “Good pay” (labelled Pay in the tables below), “Good hours” (Hours), and “Generous holidays” (Holidays) — and three of intrinsic work values — “An opportunity to use initiative” (Initiative), “A job in which you feel you can achieve something” (Achievement), and “A responsible job” (Responsibility). For more discussion of the measurement of work values in the EVS, see Gesthuizen et al. (2019).
To provide a focus for this illustrative application, we consider specifically gender differences in work value orientations between men and women (in EVS, gender is coded as a binary variable). This question was also examined, for example, by Bacher et al. (2022), using data for Austria from a different survey (The Social Survey of Austria).
We carry out two analyses. The second of them will be a cross-sectional analysis of all 36 countries in the EVS data. But first we consider just one country, the Netherlands. In addition to the work value items and gender, a set of covariates are included: the respondent’s age (in categories: 15–29, 30–49 or 50– years), whether they are in a registered partnership and/or live with a partner (yes/no), the age of the youngest person in the household (0–5, 6-17, or older), respondent’s highest level of education (less than upper secondary, upper secondary, or higher), occupation-based social class (European Socioeconomic Classification [ESeC08]; grouped by type of employment regulation of the occupation as Service relationship [ESeC08 classes 1,2], Mixed [3,6], Labour contract [7,8,9] or Self-employed [4,5]), whether they are currently in paid employment (yes/no), and the highest education level of the respondent’s parents (coded as for own education). These parallel, as far as possible with the EVS data, the covariates used by Bacher et al. (2022). This analysis dataset for the Netherlands includes $n=1374$ respondents with observed values of all the covariates and at least one of the six value items.
In the notation of Section 2, we consider here the structural model $p(\boldsymbol{\eta}_{i(1)}|\mathbf{X}_{i};\boldsymbol{\theta}_{21})$ where $\boldsymbol{\eta}_{i(1)}=(\eta_{i1},\eta_{i2})^{\prime}$ and $\eta_{i1}$ stands for extrinsic and $\eta_{i2}$ for intrinsic work values. They are responses to covariates $\mathbf{X}_{i}$, and there are no observed response variables. As defined in (3), the model includes separate linear models for $\eta_{i1}$ and $\eta_{i2}$, plus their conditional association which we report as the correlation $\rho=\text{corr}(\eta_{i1},\eta_{i2}|\boldsymbol{X}_{i})$. The two latent variables are measured by the corresponding 3 items each, with logistic measurement models (8).
We consider three models, with different sets of covariates: Model (1) which includes only gender, Model (2) which adds respondent’s age, partnership status and age of the youngest person in the household, and Model (3) which includes all the covariates. Estimates are shown in Table 3. This includes only the loading parameters ($\lambda_{kj}$) of the measurement models, the regression coefficient of gender (as dummy variable for men) and the conditional correlation $\rho$. The full sets of structural coefficients are shown in supplementary Tables B.1 and B.2.
Table 3: Models for extrinsic and intrinsic work values given gender, controlling for different sets of covariates, estimated from EVS2017 data for Netherlands. The table shows two-step (“2-st.”), one-step (“1-st.”) and naive three-step (“3-st.”) estimates. “Step 1” denotes the estimated measurement loadings from step 1 of two-step estimation, which are used for two-step and (for latent trait scores of) three-step estimation of all three models.
Model (1) Model (2) Model (3)
Step 1 2-st. 1-st. 3-st. 2-st. 1-st. 3-st. 2-st. 1-st. 3-st.
Model for Extrinsic work values
Measurement loadings:
Pay 1.000 1.000 1.000 1.000
Hours 1.161 1.101 1.148 1.190
Holidays 3.080 3.123 2.586 2.573
Structural model: coefficient of Gender (man)
Male -0.042 -0.033 -0.033 0.025 0.022 0.024 0.018 0.009 0.018
(s.e.) (0.103) (0.104) (0.082) (0.101) (0.104) (0.080) (0.102) (0.104) (0.081)
Model for Intrinsic work values
Measurement loadings:
Initiative 1.000 1.000 1.000 1.000
Achievement 0.431 0.453 0.500 0.556
Responsibility 0.563 0.589 0.631 0.804
Structural model: coefficient of Gender (man)
Male 0.607 0.603 0.460 0.657 0.638 0.500 0.516 0.506 0.400
(s.e.) (0.193) (0.186) (0.142) (0.194) (0.176) (0.142) (0.188) (0.151) (0.139)
Correlation of Extrinsic and Intrinsic work values given covariates
0.550 0.551 0.345 0.543 0.549 0.334 0.559 0.567 0.334
* Model (1) includes only gender as covariate, Model (2) also respondent’s age, partnership status and age of
    the youngest person in the household, and Model (3) also respondent’s education, social class and work status,
    and their parents’ education.
The full sets of estimated coefficients of the structural model are shown in the supplementary appendix.
Consider first the estimated measurement models in Table 3. The column “Step 1” shows the measurement loadings that are obtained from step 1 of the two-step method. The measurement parameters from it are used for step 2 for all choices of the set of covariates $\mathbf{X}_{i}$. They are also used to calculate the trait scores that are used for three-step estimation of all the models.
In contrast, the one-step method estimates the measurement model anew whenever the structural model changes; these estimated loadings are shown under the “1-st.” columns of the table. This raises the possibility of interpretational confounding, where the definition of a latent variable is affected not only by its indicators but also by its predictors. Previously, Bakk and Kuha (2018) examined this for latent class models. In one of their examples, different choices of covariates flipped the measurement model between two qualitatively quite different configurations of the classes, so that comparisons of one-step estimates between different structural models became effectively meaningless. The differences are not similarly dramatic in our example here. All of the one-step estimates of the loadings are positive, indicating that for both the extrinsic and intrinsic factors higher values correspond to a respondent placing higher importance on that dimension of work values, and the loadings of different items are in the same order of size in each case. This is perhaps not surprising, in that we might expect that for continuous latent variables such changes in the measurement model will be quantitative shifts rather than qualitative jumps. Nevertheless, in this more subtle way the one-step estimates for the three models in Table 3 are still for three slightly different response variables which differ both from each other and from the response in all of the two-step estimates.
Consider then the estimated structural models for the Netherlands. In terms of the settings of the simulations in Section 4, these data have $n=1374$ and $R^{2}_{\eta}$ around 0.1 for both latent factors, each of them measured by $p=3$ items with $R^{2}_{Y}$ of around 0.3–0.8 and $\pi_{Y}$ of 0.4–0.7 (different for different items). These place this example around the middle of the simulation settings in terms of the strength of the measurement model, but with a somewhat larger sample size than even the larger $n=1000$ considered there. The simulations suggest that the two-step and one-step estimates should both behave well here. They are also similar to each other, with somewhat higher standard errors for the two-step estimates. For the naive three-step estimates, for better comparability we first re-scaled the factor scores so that their variances equal the estimated marginal variances of the corresponding latent variables from step 1 of two-step estimation. Even after this, the three-step estimates of regression coefficients are attenuated toward zero, as would be expected given the theoretical and simulation results above.
Substantively, the results in Table 3 show that individuals’ intrinsic and extrinsic work values, as measured by these survey items, are strongly positively correlated even given covariates. Between genders, there is essentially no difference in extrinsic values, but for intrinsic ones men indicate on average significantly (at the 5% level) higher level of importance on them. Some other significant associations with covariates are also found, as shown by the estimated coefficients in supplementary Tables B.1 and B.2. The importance of extrinsic rewards of a job (pay, hours and holidays) tends to be higher for respondents who are younger, have young children at home, or are currently in paid employment. For intrinsic values (related to the scope of initiative, achievement and responsibility), the expected levels are higher for individuals who have (or whose parents have) higher levels education and who are in professional, administrative or managerial occupations.
Summarising previous literature on gender differences in work values, Bacher et al. (2022) observed that “[…]intrinsic work values, especially altruistic or social work values, seem still to be more important for women than men, at least in some countries and/or in combination with other factors[…]” and that “Gender differences in extrinsic work value orientations are absent or smaller than those for intrinsic work value orientations”. In their own analysis in Austria, they also found higher levels of intrinsic values for women. Our conclusion from the Netherlands is the opposite. This may be because the three intrinsic items in EVS do not include ones of “altruistic or social work” kind. But it could also be an instance of variation between different country contexts. To examine this further, we can use the EVS to carry out a comparative cross-national analysis. Here we include as covariates the respondents’ gender, age, education, social class and work status; the other covariates were omitted from this illustrative example because they had large proportions of missing data in some countries. The list of the 36 countries and their sample sizes in this analysis are shown in Tables B.3 and B.4 of the supplementary appendix.
Figure 1: Estimates of the coefficient (with 95% confidence intervals) of dummy variable for men in models for intrinsic work values (circles and solid lines) and extrinsic work values (triangles and dashed lines). Two-step estimates obtained separately for each country in the 2017 European Values Survey, but with the same measurement model for each country estimated from pooled data in step 1.
Note: The coefficients are expressed on a scale where the residual variance of both factors for the Netherlands is 1. The models also include as covariates the respondent’s age, education, social class and employment status (working vs. not).
We allow all parameters of the structural model (regression coefficients and residual variances and covariance) to have different values in each country. Every parameter of the measurement model, in contrast, is constrained to have the same value in all countries, to specify full measurement equivalence across the countries. We note that because the step-1 estimates are here based on much more data than the step-2 estimates (with nearly 59,000 observations in the pooled data but 1000–2000 in each country), step 1 contributes essentially nothing to the standard errors of the estimated structural coefficients. So in this example we could safely treat the measurement parameters from step 1 as known and omit the term for them from the estimated standard errors.
Estimates of gender differences in the work values in each country are shown in supplementary Tables B.3 and B.4. We focus here on the two-step estimates, which are also shown in Figure 1. For extrinsic values, most of the point estimates indicate higher average importance of them for women, but these differences are small and mostly not statistically significant (at the 5% level). For intrinsic values there are a few more countries with significant differences, but they include ones in both directions and have no very clear geographic regularities in the variation across countries. In other words, these results do not show strong evidence of consistent gender differences in work value orientations, suggesting at best small differences with much cross-national variation in them.
As mentioned at the start of this this section, this application example also allowed us to examine two topics in the relative performance of two-step estimation. From the analysis for the Netherlands above, we discussed the observed level of interpretative confounding in one-step estimates. The second, cross-national analysis provides an extreme illustration of the difference in computational convenience between two- and one-step estimation. The model has 836 estimable parameters. One-step estimates have to be calculated for all of them at once, from the pooled dataset of all the countries. On our computer this took 4.5 hours, using just one starting value. Two-step estimation, in contrast, splits the task in a way which considerably simplifies it. In its step 1, the measurement models are estimated from the pooled dataset, as two one-trait models for three binary items each and without any covariates. Step 2 can then be done separately for each country, because their structural parameters are distinct. Even with multiple starting values (Starts=30 10 in Mplus syntax) this took just 1-2 minutes for both step 1 and each country in step 2, for a total of a little over 0.5 hours for the whole model.
6 Conclusions and discussion
Two-step estimation is a general approach which is in principle applicable to any types of latent variable models where the focus of interest is on the structural model. In this paper we have described it for the broad family of latent trait (or item response theory) models, where continuous latent variables are measured by observed categorical indicators. In our simulations and examples, two-step estimation performed very well for them, as was also expected on theoretical grounds. We compared it with one-step estimation which estimates all model parameters at once (another possible alternative, adjusted three-step estimation, is not generally available for latent trait models). The two approaches behaved generally similarly, with two-step estimates losing next to no efficiency in large samples and often having somewhat better properties in small samples. This is in line with previous findings on two-step estimation for other families of models, in particular latent class analysis of categorical variables and structural equation modelling of continuous variables. Together these studies provide increasingly strong evidence and reassurance that the two-step method is a useful and convenient approach to estimation of latent variable models, and an attractive alternative to the existing one-step and three-step methods.
We also argue that two-step estimation has two major advantages in general. One of them is practical and computational convenience, compared to one-step estimation of all parts of the model at once. This may be particularly important for latent trait models, because they are more often used with complex structural models than are latent class models, and are also inherently more computationally demanding than structural equation models. Another important characteristic of the two-step approach is that it avoids interpretational confounding, where the effective definition of the latent variables is affected by the specification of the structural model. This may be more consequential for latent class models, where the interpretation of a categorical latent class variable can change in a discontinuous way, than for latent trait models and structural equation models which have continuous latent variables.
To build on these promising results so far, different directions of further research on different elements of two-step estimation could be pursued. The accessibility of the approach would be much improved by integrated implementation of it (including calculation of correct standard errors of the estimates) in general-purpose software for latent variable models, including latent trait models. This would extend and complement the procedures that are already available for latent class and structural equation modelling. For better understanding of the scope of two-step estimation, its implementation and performance could be studied for other categories of models, for example ones which combine different types of latent variables and measurement items, or for extensions such as multilevel models and ones with non-equivalent measurement models. These and other areas of further research remain to be explored.
Declarations
Funding: No funding was received for conducting this study.
Competing interests: The authors have no relevant financial or non-financial interests to disclose.
Supplementary appendices
These supplementary materials provide the following information:
Appendix A Results of the simulations (Section 4 of the paper)
  • Path diagrams that represent the cases that are considered in the simulations of Section 4 of the paper are shown in Figures A.1 and A.2:
    • Cases with 1 or 2 latent variables (simulation cases A–D) in Figure A.1.
    • Cases with 3 latent variables (simulation cases E–G) in Figure A.2.
  • Results of the simulations for point estimates of measurement parameters in simulation cases A, B2 and B3 are given in Tables A.1 (for the measurement loadings) and A.2 (for the measurement intercepts):
  • Results of the simulations for estimates of structural regression coefficients are given in Tables A.3–A.25:
    • Case A (latent covariate, latent response) in Tables A.2 (point estimates) and A.3 (standard errors and confidence intervals). These tables are also shown in Section 4 of the main paper, as Tables 1 and 2 respectively.
    • Case B1, B2 and B3 (observed covariate, latent response) in Tables A.5, A.7, and A.9 (point estimates)and A.6, A.8, and A.10 (standard errors and confidence intervals).
    • Cases C1 and C2 (latent covariate, observed response) in Tables A.11 and A.13 (point estimates) and A.12 and A.14 (standard errors and confidence intervals).
    • Case D (like case A, but with ordinal measurement items) in Tables A.15 (point estimates) and A.16 (standard errors and confidence intervals).
    • Case E (three latent variables) in in Tables A.17 (and with medians rather than means of the estimates in A.18) and A.19 (standard errors and confidence intervals).
    • Case F (like case E, but with cross-loadings in the measurement model) in in Tables A.20 (and with medians rather than means of the estimates in A.21) and A.22 (standard errors and confidence intervals).
    • Case G (misspecified model: true model as in case F, but estimation model as in case E) in in Tables A.23 (and with medians rather than means of the estimates in A.24) and A.25 (standard errors and confidence intervals).
Appendix B Results of the example (Section 5 of the paper)
Further results of the applied example discussed in Section 5 are given Tables B.1–B.4:
  • Full sets of coefficients for the structural models for Netherlands in Tables B.1–B.2.
    • Of these, the results for the coefficient of the dummy variable for Male are also shown in Section 5, as part of Table 3.
  • The list of countries in the cross-national analysis, their sample sizes, and estimated coefficients of the dummy variable for Male in each of them in Tables B.3–B.4.
    • Of these, the two-step estimates of the coefficients and their 95% confidence intervals are also displayed in Figure 1 in Section 5 (with the further standardisation that there the coefficients are expressed on a scale where the residual variance of both intrinsic and extrinsic values for the Netherlands is 1).
Figure A.1: Path diagrams representing cases A–D that are considered in the simulation studies of Section 4 of the paper. Here the dashed arrows represent the measurement model that is estimated in step 1 of two-step estimation, and the solid arrow the structural model (regression model for $\eta_{2}$ given $\eta_{1}$, $\eta_{2}$ given $X$, or $Z$ given $\eta_{1}$) that is estimated in its step 2.
Figure A.2: Path diagrams representing cases E–G that are considered in the simulation studies of Section 4 of the paper. Here the dashed and dotted arrows represent the measurement model that is estimated in step 1 of two-step estimation, and the solid arrows the structural model (regression models for $\eta_{2}$ given $\eta_{1}$ and for $\eta_{3}$ given $(\eta_{1},\eta_{2}))$ that is estimated in its step 2.
Table A.1: Simulation means of two-step and one-step point estimates of all loadings parameters ($\lambda_{kj}$) of the measurement models in simulation cases A, B2 and B3. In each case the true value of all the loadings is 1.
Case A Case B2 Case B3
$p$ $\pi_{Y}$ $R^{2}_{Y}$ $R^{2}_{\eta}$ 2-step 1-step 2-step 1-step 2-step 1-step
$n=200$
4 0.5 0.4 0.0 1.095 1.046 1.116 1.118 1.118 1.120
4 0.8 0.4 0.0 1.345 1.632 1.349 1.358 1.331 1.397
4 0.5 0.6 0.0 1.068 0.916 1.081 1.082 1.080 1.081
4 0.8 0.6 0.0 1.107 1.172 1.117 1.115 1.139 1.141
8 0.5 0.4 0.0 1.045 -0.597 1.035 1.036 1.049 1.049
8 0.8 0.4 0.0 1.089 1.096 1.071 1.071 1.079 1.080
8 0.5 0.6 0.0 1.031 -4.574 1.036 1.036 1.042 1.042
8 0.8 0.6 0.0 1.050 1.059 1.056 1.056 1.054 1.054
4 0.5 0.4 0.2 1.095 1.087 1.080 1.071 1.132 1.123
4 0.8 0.4 0.2 1.279 1.314 1.288 1.178 1.588 1.436
4 0.5 0.6 0.2 1.084 0.659 1.072 1.066 1.058 1.058
4 0.8 0.6 0.2 1.126 1.127 1.086 1.058 1.110 1.115
8 0.5 0.4 0.2 1.048 -0.662 1.055 1.050 1.038 1.039
8 0.8 0.4 0.2 1.088 1.090 1.101 1.091 1.105 1.106
8 0.5 0.6 0.2 1.028 -1.730 1.036 1.033 1.032 1.033
8 0.8 0.6 0.2 1.047 1.049 1.058 1.055 1.053 1.054
4 0.5 0.4 0.4 1.096 1.084 1.103 1.054 1.132 1.116
4 0.8 0.4 0.4 1.281 1.182 1.298 1.112 2.336 1.836
4 0.5 0.6 0.4 1.071 0.930 1.060 1.038 1.095 1.092
4 0.8 0.6 0.4 1.107 1.092 1.118 1.084 1.150 1.162
8 0.5 0.4 0.4 1.054 0.040 1.056 1.049 1.071 1.073
8 0.8 0.4 0.4 1.105 1.098 1.084 1.074 1.200 1.203
8 0.5 0.6 0.4 1.034 -2.222 1.037 1.032 1.029 1.031
8 0.8 0.6 0.4 1.055 1.055 1.046 1.044 1.066 1.067
$n=1000$
4 0.5 0.4 0.0 1.019 1.022 1.020 1.020 1.021 1.021
4 0.8 0.4 0.0 1.031 1.034 1.025 1.025 1.037 1.036
4 0.5 0.6 0.0 1.015 1.015 1.011 1.011 1.012 1.012
4 0.8 0.6 0.0 1.019 1.023 1.018 1.018 1.017 1.017
8 0.5 0.4 0.0 1.005 1.006 1.003 1.003 1.006 1.006
8 0.8 0.4 0.0 1.017 1.020 1.015 1.015 1.007 1.007
8 0.5 0.6 0.0 1.005 1.008 1.009 1.009 1.006 1.006
8 0.8 0.6 0.0 1.007 1.010 1.005 1.005 1.014 1.014
4 0.5 0.4 0.2 1.009 1.011 1.018 1.012 1.022 1.023
4 0.8 0.4 0.2 1.028 1.025 1.037 1.024 1.046 1.046
4 0.5 0.6 0.2 1.013 1.012 1.014 1.012 1.013 1.012
4 0.8 0.6 0.2 1.011 1.011 1.009 1.010 1.008 1.008
8 0.5 0.4 0.2 1.008 1.009 1.010 1.009 1.011 1.011
8 0.8 0.4 0.2 1.017 1.020 1.011 1.011 1.014 1.014
8 0.5 0.6 0.2 1.008 1.009 1.002 1.003 1.006 1.006
8 0.8 0.6 0.2 1.008 1.009 1.007 1.006 1.014 1.014
4 0.5 0.4 0.4 1.018 1.014 1.015 1.006 1.024 1.022
4 0.8 0.4 0.4 1.026 1.023 1.028 1.016 1.038 1.042
4 0.5 0.6 0.4 1.016 1.012 1.013 1.009 1.016 1.016
4 0.8 0.6 0.4 1.012 1.012 1.017 1.013 1.019 1.020
8 0.5 0.4 0.4 1.008 1.008 1.010 1.010 1.018 1.018
8 0.8 0.4 0.4 1.013 1.018 1.015 1.013 1.011 1.011
8 0.5 0.6 0.4 1.006 1.007 1.007 1.005 1.008 1.008
8 0.8 0.6 0.4 1.008 1.009 1.009 1.009 1.007 1.008
Note: $p$ denotes number of measurement items $Y_{j}$,
          $\pi_{Y}$ marginal proportion of $Y_{j}=1$, and
         $R_{Y}^{2}$ and $R^{2}_{\eta}$ the $R^{2}$ statistics in models for $Y_{j}$ and $\eta_{2}$.
Table A.2: Simulation means of two-step and one-step point estimates of all intercept parameters ($\tau_{kj}$) of the measurement models in simulation cases A, B2 and B3. In each case the true value of all the intercepts is 0.
Case A Case B2 Case B3
$p$ $\pi_{Y}$ $R^{2}_{Y}$ $R^{2}_{\eta}$ 2-step 1-step 2-step 1-step 2-step 1-step
$n=200$
4 0.5 0.4 0.0 -0.013 -0.008 -0.002 -0.002 0.001 0.002
4 0.8 0.4 0.0 0.462 0.743 0.442 0.452 0.403 0.490
4 0.5 0.6 0.0 0.005 0.020 -0.003 -0.003 -0.001 -0.001
4 0.8 0.6 0.0 0.085 0.148 0.092 0.090 0.136 0.138
8 0.5 0.4 0.0 0.009 0.023 0.001 0.001 -0.006 -0.006
8 0.8 0.4 0.0 0.103 0.110 0.093 0.093 0.110 0.111
8 0.5 0.6 0.0 -0.006 0.002 0.005 0.005 0.007 0.007
8 0.8 0.6 0.0 0.036 0.045 0.046 0.046 0.061 0.061
4 0.5 0.4 0.2 -0.003 -0.004 0.002 $0$0.001 -0.004
4 0.8 0.4 0.2 0.339 0.348 0.349 0.215 0.847 0.548
4 0.5 0.6 0.2 -0.006 -0.014 -0.013 -0.013 0.003 0.001
4 0.8 0.6 0.2 0.097 0.099 0.058 0.035 0.104 0.105
8 0.5 0.4 0.2 -$0$0.005 0.005 0.005 -0.002 -0.002
8 0.8 0.4 0.2 0.109 0.111 0.132 0.118 0.146 0.147
8 0.5 0.6 0.2 0.005 -0.024 -0.010 -0.010 0.006 0.005
8 0.8 0.6 0.2 0.050 0.053 0.054 0.051 0.045 0.045
4 0.5 0.4 0.4 -0.002 -0.000 0.008 0.007 0.007 0.003
4 0.8 0.4 0.4 0.327 0.207 0.398 0.143 2.015 1.166
4 0.5 0.6 0.4 0.002 -0.002 0.014 0.014 -0.004 -0.009
4 0.8 0.6 0.4 0.094 0.074 0.099 0.062 0.129 0.139
8 0.5 0.4 0.4 -0.007 0.003 0.003 0.002 -0.004 -0.007
8 0.8 0.4 0.4 0.137 0.128 0.096 0.082 0.286 0.286
8 0.5 0.6 0.4 0.003 0.035 -0.010 -0.010 0.005 0.003
8 0.8 0.6 0.4 0.047 0.047 0.043 0.038 0.083 0.083
$n=1000$
4 0.5 0.4 0.0 -0.001 -0.001 -0.001 -0.001 -0.002 -0.002
4 0.8 0.4 0.0 0.045 0.048 0.035 0.034 0.047 0.046
4 0.5 0.6 0.0 -0.002 -0.002 0.000 0.000 -0.001 -0.001
4 0.8 0.6 0.0 0.016 0.019 0.006 0.006 0.012 0.013
8 0.5 0.4 0.0 0.000 0.000 0.003 0.003 -0.004 -0.004
8 0.8 0.4 0.0 0.020 0.023 0.018 0.018 0.012 0.012
8 0.5 0.6 0.0 0.003 0.003 0.002 0.002 -0.003 -0.003
8 0.8 0.6 0.0 0.004 0.007 0.010 0.010 0.020 0.020
4 0.5 0.4 0.2 -0.004 -0.004 0.002 0.002 $0$0.000
4 0.8 0.4 0.2 0.033 0.029 0.043 0.027 0.058 0.058
4 0.5 0.6 0.2 -0.002 -0.002 0.001 0.001 -0.003 -0.003
4 0.8 0.6 0.2 0.008 0.009 0.009 0.010 0.004 0.004
8 0.5 0.4 0.2 -0.001 -0.001 0.001 0.001 -$0$0.000
8 0.8 0.4 0.2 0.021 0.025 0.013 0.013 0.014 0.014
8 0.5 0.6 0.2 0.004 0.004 -0.004 -0.004 -0.003 -0.003
8 0.8 0.6 0.2 0.005 0.006 0.006 0.006 0.006 0.006
4 0.5 0.4 0.4 0.001 0.001 -0.007 -0.007 -0.003 -0.004
4 0.8 0.4 0.4 0.030 0.026 0.033 0.017 0.059 0.063
4 0.5 0.6 0.4 -0.003 -0.003 -0.005 -0.004 0.001 -0.000
4 0.8 0.6 0.4 0.007 0.008 0.016 0.012 0.019 0.019
8 0.5 0.4 0.4 -0.001 -0.001 0.001 0.001 0.005 0.004
8 0.8 0.4 0.4 0.018 0.023 0.014 0.012 0.017 0.017
8 0.5 0.6 0.4 0.001 0.001 0.006 0.006 -$0$0.000
8 0.8 0.6 0.4 0.006 0.007 0.002 0.002 0.009 0.009
Note: $p$ denotes number of measurement items $Y_{j}$,
          $\pi_{Y}$ marginal proportion of $Y_{j}=1$, and
         $R_{Y}^{2}$ and $R^{2}_{\eta}$ the $R^{2}$ statistics in models for $Y_{j}$ and $\eta_{2}$.
Table A.3: Simulation results for point estimates of structural regression coefficient $\beta_{1}$ of a latent covariate $\eta_{1}$ for a conditionally normally distributed latent response $\eta_{2}$ (Case A).
Bias RMSE MAE
$p$ $\pi_{Y}$ $R^{2}_{Y}$ $R^{2}_{\eta}$ $\beta_{1}$ 2-step 1-step 3-step 2-step 1-step 3-step 2-step 1-step 3-step
$n=200$
4 0.5 0.4 0.0 0.000 -0.004 -0.003 -0.002 0.133 0.141 0.081 0.077 0.080 0.046
4 0.8 0.4 0.0 0.000 -0.007 -0.004 -0.000 0.240 0.276 0.110 0.101 0.113 0.044
4 0.5 0.6 0.0 0.000 0.007 0.007 0.005 0.115 0.117 0.080 0.069 0.070 0.048
4 0.8 0.6 0.0 0.000 0.002 0.002 0.003 0.153 0.154 0.091 0.081 0.084 0.048
8 0.5 0.4 0.0 0.000 -0.001 -0.001 -0.001 0.104 0.104 0.078 0.063 0.062 0.047
8 0.8 0.4 0.0 0.000 0.008 0.009 0.006 0.147 0.151 0.087 0.081 0.083 0.047
8 0.5 0.6 0.0 0.000 -0.003 -0.003 -0.002 0.097 0.095 0.077 0.062 0.059 0.050
8 0.8 0.6 0.0 0.000 -0.002 -0.002 -0.000 0.114 0.116 0.080 0.072 0.073 0.050
4 0.5 0.4 0.2 0.447 0.016 0.035 -0.162 0.236 0.237 0.221 0.132 0.132 0.202
4 0.8 0.4 0.2 0.447 0.031 0.070 -0.208 0.404 0.385 0.300 0.183 0.177 0.267
4 0.5 0.6 0.2 0.447 0.019 0.025 -0.116 0.182 0.189 0.176 0.117 0.116 0.148
4 0.8 0.6 0.2 0.447 0.030 0.039 -0.144 0.257 0.244 0.225 0.135 0.134 0.183
8 0.5 0.4 0.2 0.447 0.013 0.010 -0.104 0.177 0.188 0.170 0.109 0.110 0.137
8 0.8 0.4 0.2 0.447 0.023 0.040 -0.152 0.249 0.250 0.222 0.144 0.136 0.197
8 0.5 0.6 0.2 0.447 0.005 -0.008 -0.081 0.143 0.159 0.143 0.092 0.093 0.116
8 0.8 0.6 0.2 0.447 0.024 0.028 -0.100 0.183 0.181 0.170 0.114 0.111 0.137
4 0.5 0.4 0.4 0.632 0.028 0.047 -0.223 0.309 0.285 0.301 0.176 0.162 0.268
4 0.8 0.4 0.4 0.632 0.042 0.081 -0.284 0.608 0.457 0.436 0.228 0.199 0.363
4 0.5 0.6 0.4 0.632 0.048 0.052 -0.143 0.275 0.254 0.247 0.140 0.147 0.199
4 0.8 0.6 0.4 0.632 0.038 0.048 -0.195 0.326 0.283 0.295 0.174 0.161 0.255
8 0.5 0.4 0.4 0.632 0.020 0.022 -0.142 0.213 0.222 0.216 0.131 0.134 0.176
8 0.8 0.4 0.4 0.632 0.033 0.054 -0.204 0.344 0.346 0.309 0.177 0.166 0.258
8 0.5 0.6 0.4 0.632 0.033 0.006 -0.087 0.196 0.227 0.182 0.127 0.124 0.138
8 0.8 0.6 0.4 0.632 0.024 0.030 -0.138 0.223 0.218 0.220 0.137 0.135 0.180
$n=1000$
4 0.5 0.4 0.0 0.000 0.000 0.000 0.000 0.053 0.054 0.032 0.036 0.036 0.021
4 0.8 0.4 0.0 0.000 -0.001 -0.001 -0.000 0.072 0.073 0.032 0.049 0.049 0.021
4 0.5 0.6 0.0 0.000 -0.001 -0.001 -0.001 0.045 0.045 0.031 0.030 0.030 0.021
4 0.8 0.6 0.0 0.000 0.001 0.001 0.001 0.054 0.055 0.032 0.036 0.036 0.021
8 0.5 0.4 0.0 0.000 0.000 0.000 0.000 0.045 0.045 0.033 0.031 0.031 0.023
8 0.8 0.4 0.0 0.000 0.001 0.001 0.001 0.059 0.059 0.034 0.038 0.038 0.022
8 0.5 0.6 0.0 0.000 -0.001 -0.001 -0.000 0.040 0.040 0.032 0.026 0.026 0.021
8 0.8 0.6 0.0 0.000 0.000 0.000 0.000 0.045 0.045 0.032 0.030 0.030 0.021
4 0.5 0.4 0.2 0.447 0.006 0.012 -0.174 0.092 0.090 0.183 0.061 0.059 0.178
4 0.8 0.4 0.2 0.447 0.011 0.021 -0.227 0.124 0.124 0.236 0.078 0.077 0.234
4 0.5 0.6 0.2 0.447 0.001 0.003 -0.131 0.075 0.074 0.142 0.053 0.052 0.135
4 0.8 0.6 0.2 0.447 0.011 0.013 -0.160 0.091 0.090 0.171 0.057 0.059 0.168
8 0.5 0.4 0.2 0.447 0.004 0.005 -0.112 0.073 0.073 0.125 0.047 0.047 0.116
8 0.8 0.4 0.2 0.447 0.010 0.015 -0.163 0.094 0.093 0.174 0.062 0.061 0.166
8 0.5 0.6 0.2 0.447 0.002 0.003 -0.084 0.062 0.062 0.098 0.039 0.039 0.087
8 0.8 0.6 0.2 0.447 0.008 0.009 -0.113 0.074 0.074 0.126 0.048 0.048 0.118
4 0.5 0.4 0.4 0.632 0.009 0.014 -0.244 0.118 0.112 0.255 0.076 0.076 0.250
4 0.8 0.4 0.4 0.632 0.003 0.014 -0.319 0.156 0.150 0.329 0.103 0.099 0.329
4 0.5 0.6 0.4 0.632 0.006 0.007 -0.177 0.093 0.090 0.190 0.061 0.058 0.182
4 0.8 0.6 0.4 0.632 0.009 0.010 -0.218 0.117 0.109 0.232 0.074 0.071 0.223
8 0.5 0.4 0.4 0.632 0.003 0.005 -0.158 0.088 0.086 0.172 0.059 0.057 0.164
8 0.8 0.4 0.4 0.632 -0.001 0.007 -0.231 0.114 0.111 0.244 0.078 0.072 0.237
8 0.5 0.6 0.4 0.632 0.007 0.008 -0.110 0.080 0.079 0.128 0.054 0.053 0.112
8 0.8 0.6 0.4 0.632 0.006 0.008 -0.152 0.088 0.086 0.167 0.060 0.058 0.156
Note: $p$ denotes number of measurement items $Y_{j}$, $\pi_{Y}$ marginal proportion of $Y_{j}=1$,
         $R_{Y}^{2}$ and $R^{2}_{\eta}$ the $R^{2}$ statistics in models for $Y_{j}$ and $\eta_{2}$, and $\beta_{1}$ true value of $\beta_{1}$.
Table A.4: Simulation results for standard error estimates of structural regression coefficient $\beta_{1}$ of a latent covariate $\eta_{1}$ for a conditionally normally distributed latent response $\eta_{2}$ (Case A).
Simulation standard deviation 2-step
vs. mean of est. std. error Coverage of without
2-step 1-step 95% conf. interval step-1 var.
$p$ $\pi_{Y}$ $R^{2}_{Y}$ $R^{2}_{\eta}$ $\beta_{1}$ s.d. m(se) s.d. m(se) 2-st. 1-st. 3-st. %var. cover.
$n=200$
4 0.5 0.4 0.0 0.000 0.133 0.137 0.141 0.138 99.5 99.3 95.6 87.3 96.0
4 0.8 0.4 0.0 0.000 0.240 0.224 0.276 0.235 99.9 100.0 94.7 78.0 96.0
4 0.5 0.6 0.0 0.000 0.115 0.114 0.116 0.114 98.9 98.5 94.3 91.4 94.9
4 0.8 0.6 0.0 0.000 0.153 0.146 0.154 0.145 99.2 98.6 94.2 88.2 95.1
8 0.5 0.4 0.0 0.000 0.105 0.105 0.105 0.102 98.5 98.5 95.7 92.5 95.8
8 0.8 0.4 0.0 0.000 0.147 0.143 0.151 0.143 99.3 99.0 95.0 87.1 95.9
8 0.5 0.6 0.0 0.000 0.097 0.095 0.095 0.090 98.4 98.1 95.5 94.3 95.7
8 0.8 0.6 0.0 0.000 0.115 0.112 0.116 0.112 99.1 99.0 94.5 92.1 95.0
4 0.5 0.4 0.2 0.447 0.236 0.223 0.235 0.217 90.5 90.5 36.8 41.3 77.9
4 0.8 0.4 0.2 0.447 0.403 0.333 0.378 0.331 88.5 90.3 28.2 41.6 75.4
4 0.5 0.6 0.2 0.447 0.181 0.183 0.187 0.178 92.8 92.7 48.6 43.3 82.8
4 0.8 0.6 0.2 0.447 0.256 0.230 0.241 0.219 91.5 91.9 38.4 42.8 79.3
8 0.5 0.4 0.2 0.447 0.176 0.162 0.188 0.160 90.7 89.3 48.7 39.7 77.0
8 0.8 0.4 0.2 0.447 0.248 0.226 0.247 0.226 91.1 91.8 36.5 39.7 75.1
8 0.5 0.6 0.2 0.447 0.143 0.143 0.159 0.137 92.1 90.0 57.4 42.3 78.9
8 0.8 0.6 0.2 0.447 0.181 0.175 0.179 0.171 92.1 92.6 51.4 40.9 78.5
4 0.5 0.4 0.4 0.632 0.308 0.295 0.281 0.263 91.9 92.7 28.6 27.3 72.2
4 0.8 0.4 0.4 0.632 0.607 0.439 0.450 0.385 89.7 92.8 19.9 27.9 69.6
4 0.5 0.6 0.4 0.632 0.271 0.254 0.248 0.229 94.0 94.6 36.7 28.8 73.8
4 0.8 0.6 0.4 0.632 0.324 0.316 0.279 0.267 93.2 94.5 29.5 28.3 71.2
8 0.5 0.4 0.4 0.632 0.212 0.208 0.221 0.203 92.5 92.5 40.7 24.9 66.1
8 0.8 0.4 0.4 0.632 0.342 0.284 0.342 0.280 91.0 92.7 27.6 24.6 65.1
8 0.5 0.6 0.4 0.632 0.193 0.189 0.227 0.175 94.6 90.2 47.7 25.9 66.8
8 0.8 0.6 0.4 0.632 0.221 0.218 0.216 0.211 93.2 94.1 38.4 26.2 68.5
$n=1000$
4 0.5 0.4 0.0 0.000 0.053 0.054 0.054 0.054 95.9 95.7 95.1 97.1 95.1
4 0.8 0.4 0.0 0.000 0.072 0.076 0.073 0.076 97.9 97.2 96.3 95.1 96.3
4 0.5 0.6 0.0 0.000 0.045 0.046 0.045 0.046 96.5 96.4 95.2 98.3 95.4
4 0.8 0.6 0.0 0.000 0.054 0.056 0.055 0.056 96.8 96.8 95.9 97.6 96.2
8 0.5 0.4 0.0 0.000 0.045 0.044 0.045 0.044 95.3 95.2 94.5 98.2 94.5
8 0.8 0.4 0.0 0.000 0.059 0.056 0.059 0.056 95.0 94.8 93.7 96.9 94.0
8 0.5 0.6 0.0 0.000 0.040 0.040 0.040 0.040 96.1 95.9 95.2 98.8 95.2
8 0.8 0.6 0.0 0.000 0.045 0.046 0.046 0.046 96.3 96.2 95.1 98.4 95.3
4 0.5 0.4 0.2 0.447 0.092 0.091 0.089 0.089 94.4 95.4 5.9 41.6 79.5
4 0.8 0.4 0.2 0.447 0.124 0.123 0.122 0.120 93.6 94.1 3.2 42.6 80.2
4 0.5 0.6 0.2 0.447 0.075 0.076 0.074 0.074 94.8 94.6 12.6 43.8 79.3
4 0.8 0.6 0.2 0.447 0.090 0.090 0.089 0.089 96.0 95.8 8.4 43.8 81.3
8 0.5 0.4 0.2 0.447 0.073 0.070 0.073 0.069 93.9 94.6 18.8 39.3 77.7
8 0.8 0.4 0.2 0.447 0.093 0.091 0.092 0.091 93.9 94.6 8.5 38.7 76.6
8 0.5 0.6 0.2 0.447 0.062 0.062 0.062 0.062 94.6 94.7 28.6 41.9 79.2
8 0.8 0.6 0.2 0.447 0.074 0.073 0.073 0.072 95.1 94.8 17.9 41.4 80.1
4 0.5 0.4 0.4 0.632 0.118 0.118 0.112 0.109 94.3 94.9 2.3 28.9 70.9
4 0.8 0.4 0.4 0.632 0.156 0.155 0.149 0.143 93.4 93.8 1.6 29.1 70.4
4 0.5 0.6 0.4 0.632 0.093 0.097 0.089 0.092 95.7 95.9 6.8 30.5 73.9
4 0.8 0.6 0.4 0.632 0.116 0.115 0.109 0.108 94.7 94.7 4.0 30.4 73.9
8 0.5 0.4 0.4 0.632 0.088 0.089 0.086 0.087 95.5 96.2 7.9 24.6 68.0
8 0.8 0.4 0.4 0.632 0.114 0.114 0.110 0.111 93.3 94.5 2.6 24.5 65.3
8 0.5 0.6 0.4 0.632 0.080 0.080 0.078 0.077 95.1 94.6 19.8 26.3 69.2
8 0.8 0.6 0.4 0.632 0.088 0.092 0.086 0.089 95.5 95.8 10.5 26.3 70.8
Note: $p$ denotes number of measurement items $Y_{j}$, $\pi_{Y}$ marginal proportion of $Y_{j}=1$,
         $R_{Y}^{2}$ and $R^{2}_{\eta}$ the $R^{2}$ statistics in models for $Y_{j}$ and $\eta_{2}$, and $\beta_{1}$ true value of $\beta_{1}$.
$\dagger$ Average % of the variance of 2-step estimator accounted for by step-2 variance alone,
       and coverage of 95% confidence interval if only this variance is included.
Table A.5: Simulation results for point estimates of structural regression coefficient $\beta_{1}$ of a normally distributed covariate $X$ for a latent response $\eta_{2}$ (Case B1) .
Bias RMSE MAE
$p$ $\pi_{Y}$ $R^{2}_{Y}$ $R^{2}_{\eta}$ $\beta_{1}$ 2-step 1-step 3-step 2-step 1-step 3-step 2-step 1-step 3-step
$n=200$
4 0.5 0.4 0.0 0.000 -0.003 -0.004 -0.002 0.146 0.151 0.090 0.084 0.086 0.052
4 0.8 0.4 0.0 0.000 0.007 0.006 0.003 0.196 0.208 0.094 0.100 0.102 0.045
4 0.5 0.6 0.0 0.000 0.008 0.008 0.006 0.207 0.208 0.145 0.125 0.126 0.086
4 0.8 0.6 0.0 0.000 0.002 0.002 0.002 0.251 0.253 0.151 0.144 0.145 0.085
8 0.5 0.4 0.0 0.000 -0.005 -0.005 -0.004 0.132 0.132 0.098 0.085 0.085 0.063
8 0.8 0.4 0.0 0.000 0.003 0.003 0.002 0.150 0.152 0.088 0.096 0.097 0.056
8 0.5 0.6 0.0 0.000 0.004 0.004 0.003 0.187 0.185 0.150 0.118 0.118 0.095
8 0.8 0.6 0.0 0.000 -0.013 -0.013 -0.009 0.203 0.204 0.141 0.133 0.135 0.093
4 0.5 0.4 0.2 0.662 0.030 0.039 -0.239 0.251 0.229 0.292 0.137 0.132 0.267
4 0.8 0.4 0.2 0.662 0.065 0.054 -0.326 0.527 0.367 0.423 0.179 0.166 0.382
4 0.5 0.6 0.2 0.993 0.045 0.048 -0.270 0.320 0.300 0.355 0.182 0.176 0.319
4 0.8 0.6 0.2 0.993 0.085 0.071 -0.359 0.520 0.394 0.475 0.221 0.210 0.418
8 0.5 0.4 0.2 0.662 0.023 0.026 -0.153 0.183 0.180 0.208 0.119 0.115 0.175
8 0.8 0.4 0.2 0.662 0.031 0.040 -0.254 0.255 0.266 0.298 0.151 0.142 0.277
8 0.5 0.6 0.2 0.993 0.042 0.042 -0.163 0.256 0.253 0.262 0.161 0.162 0.205
8 0.8 0.6 0.2 0.993 0.052 0.056 -0.264 0.315 0.312 0.345 0.188 0.190 0.302
4 0.5 0.4 0.4 0.937 0.031 0.034 -0.348 0.305 0.251 0.401 0.185 0.148 0.382
4 0.8 0.4 0.4 0.937 0.143 0.074 -0.443 0.924 0.365 0.632 0.254 0.213 0.533
4 0.5 0.6 0.4 1.405 0.083 0.071 -0.367 0.404 0.364 0.466 0.233 0.207 0.412
4 0.8 0.6 0.4 1.405 0.093 0.075 -0.527 0.642 0.449 0.663 0.269 0.252 0.611
8 0.5 0.4 0.4 0.937 0.035 0.042 -0.216 0.234 0.233 0.280 0.136 0.131 0.245
8 0.8 0.4 0.4 0.937 0.040 0.051 -0.362 0.309 0.297 0.408 0.183 0.167 0.389
8 0.5 0.6 0.4 1.405 0.051 0.051 -0.239 0.317 0.316 0.351 0.198 0.197 0.286
8 0.8 0.6 0.4 1.405 0.051 0.054 -0.391 0.365 0.356 0.470 0.214 0.207 0.434
$n=1000$
4 0.5 0.4 0.0 0.000 0.000 0.000 0.000 0.062 0.062 0.037 0.041 0.041 0.025
4 0.8 0.4 0.0 0.000 0.005 0.005 0.002 0.073 0.074 0.033 0.046 0.046 0.020
4 0.5 0.6 0.0 0.000 0.002 0.002 0.001 0.086 0.086 0.060 0.058 0.058 0.040
4 0.8 0.6 0.0 0.000 0.007 0.007 0.004 0.095 0.095 0.055 0.065 0.065 0.038
8 0.5 0.4 0.0 0.000 0.000 0.000 0.000 0.055 0.055 0.041 0.036 0.036 0.026
8 0.8 0.4 0.0 0.000 -0.003 -0.003 -0.002 0.065 0.065 0.038 0.044 0.044 0.026
8 0.5 0.6 0.0 0.000 0.002 0.002 0.002 0.080 0.079 0.064 0.052 0.051 0.041
8 0.8 0.6 0.0 0.000 -0.001 -0.001 -0.000 0.085 0.085 0.059 0.057 0.057 0.040
4 0.5 0.4 0.2 0.662 -0.004 -0.001 -0.268 0.098 0.092 0.276 0.063 0.059 0.273
4 0.8 0.4 0.2 0.662 0.009 0.014 -0.365 0.117 0.108 0.369 0.078 0.072 0.369
4 0.5 0.6 0.2 0.993 0.015 0.016 -0.292 0.127 0.123 0.307 0.084 0.079 0.300
4 0.8 0.6 0.2 0.993 0.006 0.007 -0.409 0.143 0.137 0.419 0.095 0.093 0.413
8 0.5 0.4 0.2 0.662 0.003 0.004 -0.171 0.077 0.076 0.181 0.054 0.052 0.174
8 0.8 0.4 0.2 0.662 0.003 0.003 -0.273 0.095 0.093 0.280 0.066 0.063 0.277
8 0.5 0.6 0.2 0.993 -0.001 -0.002 -0.199 0.108 0.108 0.218 0.072 0.073 0.201
8 0.8 0.6 0.2 0.993 0.001 0.001 -0.299 0.116 0.115 0.310 0.078 0.077 0.304
4 0.5 0.4 0.4 0.937 0.008 0.011 -0.371 0.123 0.107 0.380 0.080 0.070 0.378
4 0.8 0.4 0.4 0.937 0.012 0.014 -0.517 0.158 0.133 0.523 0.096 0.086 0.524
4 0.5 0.6 0.4 1.405 0.013 0.014 -0.420 0.157 0.151 0.436 0.099 0.097 0.426
4 0.8 0.6 0.4 1.405 0.011 0.011 -0.579 0.184 0.170 0.591 0.120 0.116 0.586
8 0.5 0.4 0.4 0.937 0.006 0.007 -0.241 0.097 0.093 0.252 0.062 0.060 0.243
8 0.8 0.4 0.4 0.937 0.009 0.010 -0.385 0.121 0.114 0.392 0.082 0.075 0.388
8 0.5 0.6 0.4 1.405 0.007 0.007 -0.275 0.125 0.123 0.293 0.079 0.079 0.278
8 0.8 0.6 0.4 1.405 0.020 0.018 -0.410 0.153 0.150 0.425 0.100 0.096 0.412
Note: $p$ denotes number of measurement items $Y_{j}$, $\pi_{Y}$ marginal proportion of $Y_{j}=1$,
         $R_{Y}^{2}$ and $R^{2}_{\eta}$ the $R^{2}$ statistics in models for $Y_{j}$ and $\eta_{2}$, and $\beta_{1}$ true value of $\beta_{1}$.
Table A.6: Simulation results for standard error estimates of structural regression coefficient $\beta_{1}$ of a normally distributed covariate $X$ for a latent response $\eta_{2}$ (Case B1).
Simulation standard deviation 2-step
vs. mean of est. std. error Coverage of without
2-step 1-step 95% conf. interval step-1 var.
$p$ $\pi_{Y}$ $R^{2}_{Y}$ $R^{2}_{\eta}$ $\beta_{1}$ s.d. m(se) s.d. m(se) 2-st. 1-st. 3-st. %var. cover.
$n=200$
4 0.5 0.4 0.0 0.000 0.146 0.149 0.151 0.149 98.4 98.0 95.3 92.6 95.3
4 0.8 0.4 0.0 0.000 0.196 0.204 0.208 0.201 99.2 98.1 94.8 86.5 95.0
4 0.5 0.6 0.0 0.000 0.207 0.205 0.208 0.204 96.1 95.8 93.2 95.2 93.4
4 0.8 0.6 0.0 0.000 0.251 0.240 0.253 0.243 97.4 96.8 94.0 92.8 94.2
8 0.5 0.4 0.0 0.000 0.132 0.129 0.132 0.129 96.8 96.8 95.3 95.6 95.3
8 0.8 0.4 0.0 0.000 0.150 0.150 0.152 0.150 98.2 98.0 95.1 92.6 95.0
8 0.5 0.6 0.0 0.000 0.187 0.186 0.185 0.185 95.4 95.6 94.2 96.8 94.3
8 0.8 0.6 0.0 0.000 0.203 0.202 0.203 0.202 96.5 96.2 94.7 95.9 94.7
4 0.5 0.4 0.2 0.662 0.250 0.240 0.226 0.218 94.1 95.4 27.5 40.4 80.9
4 0.8 0.4 0.2 0.662 0.523 0.386 0.363 0.301 92.3 93.7 13.5 34.6 73.7
4 0.5 0.6 0.2 0.993 0.317 0.310 0.296 0.294 95.9 96.1 38.1 46.6 83.3
4 0.8 0.6 0.2 0.993 0.513 0.400 0.387 0.351 93.4 93.8 25.4 44.6 79.2
8 0.5 0.4 0.2 0.662 0.182 0.181 0.178 0.178 95.7 95.8 47.6 45.3 82.6
8 0.8 0.4 0.2 0.662 0.253 0.232 0.263 0.227 94.9 94.7 26.0 39.3 76.2
8 0.5 0.6 0.2 0.993 0.252 0.252 0.250 0.249 95.3 95.3 61.4 51.0 85.2
8 0.8 0.6 0.2 0.993 0.310 0.289 0.307 0.285 95.5 95.6 42.1 47.5 80.6
4 0.5 0.4 0.4 0.937 0.303 0.297 0.249 0.243 94.7 96.3 17.5 24.3 66.0
4 0.8 0.4 0.4 0.937 0.913 0.566 0.357 0.323 94.0 95.4 8.3 20.3 61.8
4 0.5 0.6 0.4 1.405 0.396 0.393 0.357 0.347 96.6 96.8 26.9 30.6 75.4
4 0.8 0.6 0.4 1.405 0.635 0.636 0.442 0.407 93.7 94.6 14.5 28.6 71.2
8 0.5 0.4 0.4 0.937 0.231 0.223 0.229 0.214 96.1 94.7 29.9 27.3 71.5
8 0.8 0.4 0.4 0.937 0.307 0.291 0.293 0.270 94.4 95.0 13.9 23.2 67.2
8 0.5 0.6 0.4 1.405 0.313 0.300 0.312 0.294 95.2 95.0 43.0 32.3 74.0
8 0.8 0.6 0.4 1.405 0.361 0.349 0.352 0.337 95.9 95.8 22.8 30.3 72.7
$n=1000$
4 0.5 0.4 0.0 0.000 0.062 0.061 0.062 0.061 95.4 95.3 94.3 98.4 94.3
4 0.8 0.4 0.0 0.000 0.073 0.072 0.074 0.072 95.5 95.2 94.8 97.3 94.8
4 0.5 0.6 0.0 0.000 0.086 0.086 0.086 0.086 95.5 95.4 95.0 99.1 95.2
4 0.8 0.6 0.0 0.000 0.095 0.093 0.095 0.093 95.1 95.0 94.7 98.7 94.9
8 0.5 0.4 0.0 0.000 0.055 0.055 0.055 0.055 94.7 94.6 94.4 99.1 94.4
8 0.8 0.4 0.0 0.000 0.065 0.062 0.065 0.062 95.4 95.2 94.2 98.4 94.3
8 0.5 0.6 0.0 0.000 0.080 0.079 0.079 0.079 95.0 95.2 94.8 99.4 94.9
8 0.8 0.6 0.0 0.000 0.085 0.085 0.085 0.085 95.4 95.2 94.9 99.2 95.2
4 0.5 0.4 0.2 0.662 0.098 0.094 0.092 0.089 93.9 94.0 0.8 41.8 76.7
4 0.8 0.4 0.2 0.662 0.117 0.121 0.107 0.112 95.9 96.1 0.0 36.5 77.8
4 0.5 0.6 0.2 0.993 0.127 0.125 0.122 0.122 95.0 95.7 3.8 47.7 81.9
4 0.8 0.6 0.2 0.993 0.143 0.142 0.137 0.138 94.7 95.0 0.4 45.6 81.4
8 0.5 0.4 0.2 0.662 0.077 0.078 0.076 0.077 94.6 95.6 7.0 45.9 81.7
8 0.8 0.4 0.2 0.662 0.095 0.095 0.093 0.093 95.2 95.0 0.1 40.4 78.7
8 0.5 0.6 0.2 0.993 0.108 0.106 0.108 0.105 93.4 93.5 17.8 51.4 82.5
8 0.8 0.6 0.2 0.993 0.116 0.119 0.115 0.119 96.6 96.4 2.5 48.7 83.4
4 0.5 0.4 0.4 0.937 0.122 0.119 0.107 0.104 94.0 94.3 0.2 25.7 67.3
4 0.8 0.4 0.4 0.937 0.158 0.155 0.133 0.128 95.3 94.8 0.0 22.4 65.3
4 0.5 0.6 0.4 1.405 0.157 0.155 0.150 0.144 94.4 94.1 1.4 31.5 75.2
4 0.8 0.6 0.4 1.405 0.184 0.178 0.170 0.164 95.3 94.1 0.1 30.2 69.4
8 0.5 0.4 0.4 0.937 0.096 0.095 0.093 0.092 94.8 95.0 2.1 27.5 69.4
8 0.8 0.4 0.4 0.937 0.121 0.120 0.114 0.113 95.2 96.0 0.1 24.0 66.8
8 0.5 0.6 0.4 1.405 0.124 0.128 0.123 0.126 94.3 94.2 6.3 32.7 75.6
8 0.8 0.6 0.4 1.405 0.152 0.148 0.149 0.144 95.5 95.2 1.3 30.8 72.1
Note: $p$ denotes number of measurement items $Y_{j}$, $\pi_{Y}$ marginal proportion of $Y_{j}=1$,
         $R_{Y}^{2}$ and $R^{2}_{\eta}$ the $R^{2}$ statistics in models for $Y_{j}$ and $\eta_{2}$, and $\beta_{1}$ true value of $\beta_{1}$.
$\dagger$ Average % of the variance of 2-step estimator accounted for by step-2 variance alone,
       and coverage of 95% confidence interval if only this variance is included.
Table A.7: Simulation results for point estimates of structural regression coefficient $\beta_{1}$ of a binary covariate $X$ for a latent response $\eta_{2}$ (Case B2).
Bias RMSE MAE
$p$ $\pi_{Y}$ $R^{2}_{Y}$ $R^{2}_{\eta}$ $\beta_{1}$ 2-step 1-step 3-step 2-step 1-step 3-step 2-step 1-step 3-step
$n=200$
4 0.5 0.4 0.0 0.000 0.001 0.001 0.000 0.144 0.147 0.090 0.090 0.092 0.055
4 0.8 0.4 0.0 0.000 0.007 0.008 0.003 0.191 0.213 0.091 0.099 0.106 0.046
4 0.5 0.6 0.0 0.000 -0.001 -0.002 -0.001 0.209 0.210 0.147 0.129 0.129 0.091
4 0.8 0.6 0.0 0.000 0.002 0.001 0.001 0.259 0.254 0.160 0.138 0.141 0.081
8 0.5 0.4 0.0 0.000 -0.002 -0.002 -0.002 0.125 0.126 0.093 0.086 0.086 0.063
8 0.8 0.4 0.0 0.000 0.004 0.005 0.003 0.160 0.162 0.095 0.091 0.092 0.053
8 0.5 0.6 0.0 0.000 -0.005 -0.005 -0.004 0.185 0.184 0.149 0.126 0.126 0.100
8 0.8 0.6 0.0 0.000 0.005 0.005 0.003 0.204 0.206 0.143 0.135 0.135 0.094
4 0.5 0.4 0.2 0.662 0.016 0.026 -0.231 0.236 0.221 0.285 0.138 0.135 0.261
4 0.8 0.4 0.2 0.662 0.077 0.050 -0.318 0.567 0.316 0.439 0.194 0.171 0.383
4 0.5 0.6 0.2 0.993 0.035 0.037 -0.236 0.299 0.297 0.332 0.191 0.182 0.290
4 0.8 0.6 0.2 0.993 0.082 0.083 -0.345 0.465 0.428 0.444 0.218 0.211 0.393
8 0.5 0.4 0.2 0.662 0.021 0.023 -0.137 0.192 0.189 0.208 0.126 0.124 0.169
8 0.8 0.4 0.2 0.662 0.018 0.026 -0.257 0.239 0.232 0.298 0.147 0.140 0.277
8 0.5 0.6 0.2 0.993 0.044 0.044 -0.123 0.256 0.255 0.254 0.157 0.158 0.191
8 0.8 0.6 0.2 0.993 0.046 0.048 -0.248 0.309 0.303 0.336 0.180 0.178 0.290
4 0.5 0.4 0.4 0.937 0.042 0.039 -0.280 0.308 0.236 0.361 0.171 0.156 0.323
4 0.8 0.4 0.4 0.937 0.062 0.065 -0.479 0.587 0.388 0.563 0.234 0.182 0.525
4 0.5 0.6 0.4 1.405 0.050 0.045 -0.254 0.398 0.351 0.425 0.223 0.198 0.337
4 0.8 0.6 0.4 1.405 0.079 0.072 -0.495 0.508 0.505 0.594 0.253 0.241 0.558
8 0.5 0.4 0.4 0.937 0.015 0.018 -0.175 0.217 0.208 0.255 0.137 0.131 0.204
8 0.8 0.4 0.4 0.937 0.025 0.037 -0.361 0.290 0.279 0.406 0.181 0.164 0.381
8 0.5 0.6 0.4 1.405 0.047 0.044 -0.114 0.321 0.308 0.320 0.185 0.175 0.229
8 0.8 0.6 0.4 1.405 0.070 0.071 -0.318 0.359 0.352 0.419 0.216 0.213 0.365
$n=1000$
4 0.5 0.4 0.0 0.000 -0.000 -0.000 0.000 0.061 0.062 0.037 0.040 0.040 0.024
4 0.8 0.4 0.0 0.000 0.000 0.000 0.000 0.073 0.074 0.032 0.046 0.046 0.020
4 0.5 0.6 0.0 0.000 -0.002 -0.002 -0.001 0.082 0.083 0.058 0.055 0.056 0.039
4 0.8 0.6 0.0 0.000 0.003 0.003 0.002 0.094 0.094 0.055 0.060 0.060 0.035
8 0.5 0.4 0.0 0.000 0.001 0.001 0.001 0.056 0.056 0.042 0.037 0.037 0.027
8 0.8 0.4 0.0 0.000 0.002 0.002 0.001 0.064 0.064 0.037 0.044 0.043 0.025
8 0.5 0.6 0.0 0.000 -0.002 -0.002 -0.001 0.081 0.081 0.065 0.056 0.055 0.045
8 0.8 0.6 0.0 0.000 0.005 0.005 0.004 0.087 0.087 0.061 0.055 0.055 0.038
4 0.5 0.4 0.2 0.662 0.004 0.005 -0.249 0.093 0.086 0.257 0.058 0.055 0.253
4 0.8 0.4 0.2 0.662 0.001 0.009 -0.366 0.128 0.114 0.371 0.081 0.072 0.370
4 0.5 0.6 0.2 0.993 0.008 0.011 -0.261 0.118 0.118 0.278 0.077 0.078 0.263
4 0.8 0.6 0.2 0.993 -0.002 -0.001 -0.401 0.145 0.140 0.412 0.093 0.093 0.408
8 0.5 0.4 0.2 0.662 0.007 0.007 -0.152 0.079 0.078 0.165 0.053 0.052 0.156
8 0.8 0.4 0.2 0.662 0.008 0.009 -0.266 0.096 0.094 0.273 0.061 0.059 0.268
8 0.5 0.6 0.2 0.993 0.003 0.002 -0.161 0.109 0.109 0.187 0.069 0.070 0.164
8 0.8 0.6 0.2 0.993 0.008 0.009 -0.276 0.118 0.118 0.290 0.082 0.079 0.279
4 0.5 0.4 0.4 0.937 0.001 0.005 -0.322 0.119 0.103 0.334 0.077 0.069 0.330
4 0.8 0.4 0.4 0.937 0.005 0.011 -0.520 0.165 0.131 0.526 0.105 0.085 0.525
4 0.5 0.6 0.4 1.405 0.007 0.012 -0.296 0.154 0.141 0.325 0.105 0.097 0.308
4 0.8 0.6 0.4 1.405 0.023 0.020 -0.531 0.182 0.166 0.545 0.122 0.102 0.541
8 0.5 0.4 0.4 0.937 0.001 0.003 -0.191 0.095 0.089 0.208 0.062 0.057 0.194
8 0.8 0.4 0.4 0.937 0.006 0.010 -0.376 0.114 0.111 0.383 0.072 0.072 0.377
8 0.5 0.6 0.4 1.405 0.014 0.014 -0.148 0.129 0.124 0.191 0.085 0.083 0.161
8 0.8 0.6 0.4 1.405 0.007 0.005 -0.364 0.148 0.143 0.382 0.097 0.094 0.372
Note: $p$ denotes number of measurement items $Y_{j}$, $\pi_{Y}$ marginal proportion of $Y_{j}=1$,
         $R_{Y}^{2}$ and $R^{2}_{\eta}$ the $R^{2}$ statistics in models for $Y_{j}$ and $\eta_{2}$, and $\beta_{1}$ true value of $\beta_{1}$.
Table A.8: Simulation results for standard error estimates of structural regression coefficient $\beta_{1}$ of a binary covariate $X$ for a latent response $\eta_{2}$ (Case B2).
Simulation standard deviation 2-step
vs. mean of est. std. error Coverage of without
2-step 1-step 95% conf. interval step-1 var.
$p$ $\pi_{Y}$ $R^{2}_{Y}$ $R^{2}_{\eta}$ $\beta_{1}$ s.d. m(se) s.d. m(se) 2-st. 1-st. 3-st. %var. cover.
$n=200$
4 0.5 0.4 0.0 0.000 0.144 0.147 0.147 0.147 97.9 97.7 95.5 92.7 95.6
4 0.8 0.4 0.0 0.000 0.191 0.239 0.213 0.214 99.2 98.8 94.4 85.2 94.7
4 0.5 0.6 0.0 0.000 0.209 0.204 0.210 0.204 95.8 95.6 93.7 95.2 93.7
4 0.8 0.6 0.0 0.000 0.260 0.243 0.254 0.243 97.5 97.3 94.1 92.9 94.4
8 0.5 0.4 0.0 0.000 0.126 0.129 0.126 0.129 97.1 97.1 96.0 95.8 96.0
8 0.8 0.4 0.0 0.000 0.160 0.151 0.162 0.151 98.4 98.1 96.4 93.2 96.3
8 0.5 0.6 0.0 0.000 0.185 0.185 0.184 0.184 96.5 96.5 95.3 96.9 95.4
8 0.8 0.6 0.0 0.000 0.204 0.201 0.206 0.201 96.9 96.8 94.7 95.7 94.8
4 0.5 0.4 0.2 0.662 0.235 0.228 0.219 0.209 94.2 94.0 27.6 38.6 76.3
4 0.8 0.4 0.2 0.662 0.562 0.486 0.312 0.288 93.3 94.1 13.7 33.8 73.1
4 0.5 0.6 0.2 0.993 0.297 0.296 0.295 0.285 94.4 94.5 44.1 45.9 82.1
4 0.8 0.6 0.2 0.993 0.459 0.378 0.420 0.354 94.3 95.5 27.6 42.9 80.2
8 0.5 0.4 0.2 0.662 0.191 0.178 0.188 0.174 93.2 93.5 51.0 44.3 78.5
8 0.8 0.4 0.2 0.662 0.239 0.226 0.231 0.218 92.8 93.1 25.0 38.7 74.9
8 0.5 0.6 0.2 0.993 0.252 0.247 0.251 0.244 95.2 95.2 64.9 49.3 83.5
8 0.8 0.6 0.2 0.993 0.305 0.285 0.299 0.281 95.1 95.0 42.1 46.5 79.6
4 0.5 0.4 0.4 0.937 0.305 0.297 0.233 0.230 95.0 95.6 21.5 22.6 66.7
4 0.8 0.4 0.4 0.937 0.584 0.444 0.382 0.349 93.1 95.8 6.5 20.5 63.2
4 0.5 0.6 0.4 1.405 0.395 0.372 0.348 0.322 94.8 94.4 34.0 28.1 69.8
4 0.8 0.6 0.4 1.405 0.502 0.470 0.501 0.402 94.6 94.1 15.6 27.8 72.1
8 0.5 0.4 0.4 0.937 0.217 0.213 0.207 0.201 94.6 93.9 38.1 26.0 68.0
8 0.8 0.4 0.4 0.937 0.289 0.285 0.277 0.265 94.8 94.9 14.8 22.3 64.9
8 0.5 0.6 0.4 1.405 0.318 0.294 0.305 0.281 94.6 94.7 51.7 29.7 70.7
8 0.8 0.6 0.4 1.405 0.352 0.348 0.345 0.335 96.6 96.3 32.8 28.5 72.4
$n=1000$
4 0.5 0.4 0.0 0.000 0.061 0.061 0.062 0.061 94.9 94.6 93.8 98.5 93.9
4 0.8 0.4 0.0 0.000 0.073 0.073 0.074 0.073 95.2 95.0 94.4 97.2 94.4
4 0.5 0.6 0.0 0.000 0.082 0.086 0.083 0.085 95.9 95.9 95.8 99.1 95.9
4 0.8 0.6 0.0 0.000 0.094 0.093 0.094 0.093 94.6 94.6 94.4 98.7 94.4
8 0.5 0.4 0.0 0.000 0.056 0.055 0.056 0.055 94.3 94.4 94.1 99.1 94.2
8 0.8 0.4 0.0 0.000 0.064 0.062 0.064 0.062 95.7 95.7 95.1 98.5 95.1
8 0.5 0.6 0.0 0.000 0.081 0.079 0.081 0.079 95.1 95.6 94.8 99.3 94.7
8 0.8 0.6 0.0 0.000 0.087 0.085 0.087 0.085 94.1 94.0 93.6 99.2 93.6
4 0.5 0.4 0.2 0.662 0.093 0.094 0.085 0.088 95.4 95.8 1.7 40.3 79.2
4 0.8 0.4 0.2 0.662 0.128 0.120 0.114 0.111 92.4 93.9 0.0 36.9 73.4
4 0.5 0.6 0.2 0.993 0.118 0.122 0.118 0.119 95.8 95.1 7.1 46.3 83.8
4 0.8 0.6 0.2 0.993 0.145 0.140 0.140 0.135 93.8 94.5 0.7 44.8 79.3
8 0.5 0.4 0.2 0.662 0.079 0.077 0.078 0.076 93.9 93.8 13.8 44.8 79.3
8 0.8 0.4 0.2 0.662 0.096 0.095 0.094 0.093 94.8 94.6 0.7 39.8 79.0
8 0.5 0.6 0.2 0.993 0.109 0.105 0.109 0.104 93.7 93.2 31.0 50.3 80.7
8 0.8 0.6 0.2 0.993 0.118 0.119 0.117 0.118 95.7 95.5 4.3 47.2 82.4
4 0.5 0.4 0.4 0.937 0.119 0.115 0.103 0.098 94.0 94.1 1.1 24.2 66.4
4 0.8 0.4 0.4 0.937 0.165 0.154 0.131 0.128 94.0 94.3 0.0 22.5 63.0
4 0.5 0.6 0.4 1.405 0.154 0.150 0.141 0.136 95.2 94.4 8.1 29.0 69.3
4 0.8 0.6 0.4 1.405 0.181 0.177 0.165 0.161 95.5 94.8 0.6 28.9 71.0
8 0.5 0.4 0.4 0.937 0.095 0.092 0.089 0.088 94.1 94.6 7.8 26.2 67.2
8 0.8 0.4 0.4 0.937 0.114 0.119 0.111 0.112 96.4 95.3 0.0 23.2 69.5
8 0.5 0.6 0.4 1.405 0.128 0.125 0.123 0.121 95.0 94.9 32.9 30.1 69.2
8 0.8 0.6 0.4 1.405 0.148 0.144 0.143 0.140 94.2 94.9 2.7 29.4 72.2
Note: $p$ denotes number of measurement items $Y_{j}$, $\pi_{Y}$ marginal proportion of $Y_{j}=1$,
         $R_{Y}^{2}$ and $R^{2}_{\eta}$ the $R^{2}$ statistics in models for $Y_{j}$ and $\eta_{2}$, and $\beta_{1}$ true value of $\beta_{1}$.
$\dagger$ Average % of the variance of 2-step estimator accounted for by step-2 variance alone,
       and coverage of 95% confidence interval if only this variance is included.
Table A.9: Simulation results for point estimates of structural regression coefficient $\beta_{1}$ of a covariate $X$ with a skewed distribution for a latent response $\eta_{2}$ (Case B3).
Bias RMSE MAE
$p$ $\pi_{Y}$ $R^{2}_{Y}$ $R^{2}_{\eta}$ $\beta_{1}$ 2-step 1-step 3-step 2-step 1-step 3-step 2-step 1-step 3-step
$n=200$
4 0.5 0.4 0.0 0.000 -0.008 -0.008 -0.004 0.226 0.232 0.119 0.105 0.108 0.062
4 0.8 0.4 0.0 0.000 0.077 0.085 0.002 0.441 0.545 0.115 0.119 0.124 0.049
4 0.5 0.6 0.0 0.000 0.003 0.004 0.002 0.278 0.281 0.173 0.144 0.148 0.100
4 0.8 0.6 0.0 0.000 0.047 0.049 -0.005 0.347 0.350 0.172 0.165 0.166 0.092
8 0.5 0.4 0.0 0.000 0.003 0.003 0.002 0.173 0.174 0.119 0.090 0.090 0.066
8 0.8 0.4 0.0 0.000 0.040 0.041 0.001 0.238 0.242 0.113 0.101 0.103 0.058
8 0.5 0.6 0.0 0.000 0.004 0.004 0.003 0.244 0.243 0.181 0.130 0.130 0.104
8 0.8 0.6 0.0 0.000 0.035 0.034 -0.003 0.297 0.298 0.176 0.148 0.147 0.101
4 0.5 0.4 0.2 0.662 0.054 0.078 -0.393 0.382 0.420 0.419 0.192 0.192 0.418
4 0.8 0.4 0.2 0.662 0.311 0.372 -0.515 1.031 1.143 0.541 0.315 0.325 0.548
4 0.5 0.6 0.2 0.993 0.169 0.182 -0.546 0.587 0.607 0.578 0.278 0.277 0.572
4 0.8 0.6 0.2 0.993 0.364 0.383 -0.718 1.065 1.118 0.738 0.375 0.384 0.748
8 0.5 0.4 0.2 0.662 0.060 0.070 -0.315 0.281 0.291 0.340 0.162 0.163 0.332
8 0.8 0.4 0.2 0.662 0.147 0.172 -0.479 0.511 0.549 0.488 0.265 0.264 0.494
8 0.5 0.6 0.2 0.993 0.117 0.120 -0.466 0.424 0.429 0.504 0.223 0.223 0.482
8 0.8 0.6 0.2 0.993 0.218 0.228 -0.657 0.693 0.708 0.674 0.333 0.330 0.687
4 0.5 0.4 0.4 0.937 0.125 0.175 -0.623 0.530 0.595 0.644 0.252 0.252 0.654
4 0.8 0.4 0.4 0.937 0.526 0.809 -0.787 1.975 3.496 0.809 0.442 0.434 0.827
4 0.5 0.6 0.4 1.405 0.153 0.178 -0.921 0.641 0.672 0.944 0.336 0.344 0.959
4 0.8 0.6 0.4 1.405 0.430 0.490 -1.137 1.312 1.452 1.151 0.523 0.513 1.173
8 0.5 0.4 0.4 0.937 0.073 0.092 -0.548 0.346 0.364 0.570 0.208 0.206 0.568
8 0.8 0.4 0.4 0.937 0.211 0.283 -0.752 0.712 0.821 0.758 0.333 0.351 0.771
8 0.5 0.6 0.4 1.405 0.167 0.175 -0.814 0.538 0.547 0.847 0.289 0.291 0.844
8 0.8 0.6 0.4 1.405 0.265 0.283 -1.067 0.811 0.831 1.079 0.415 0.421 1.098
$n=1000$
4 0.5 0.4 0.0 0.000 0.001 0.001 0.000 0.068 0.068 0.039 0.041 0.041 0.024
4 0.8 0.4 0.0 0.000 0.006 0.006 -0.002 0.082 0.082 0.033 0.049 0.049 0.021
4 0.5 0.6 0.0 0.000 0.000 0.000 0.000 0.094 0.095 0.063 0.061 0.062 0.043
4 0.8 0.6 0.0 0.000 0.010 0.010 0.000 0.103 0.104 0.058 0.064 0.064 0.038
8 0.5 0.4 0.0 0.000 0.003 0.003 0.002 0.057 0.057 0.042 0.037 0.037 0.027
8 0.8 0.4 0.0 0.000 0.003 0.003 -0.002 0.069 0.069 0.039 0.045 0.045 0.026
8 0.5 0.6 0.0 0.000 -0.002 -0.002 -0.001 0.086 0.087 0.068 0.055 0.055 0.044
8 0.8 0.6 0.0 0.000 0.009 0.009 0.002 0.094 0.094 0.062 0.061 0.061 0.042
4 0.5 0.4 0.2 0.662 0.015 0.018 -0.424 0.122 0.122 0.427 0.080 0.079 0.428
4 0.8 0.4 0.2 0.662 0.029 0.036 -0.550 0.212 0.214 0.551 0.123 0.120 0.552
4 0.5 0.6 0.2 0.993 0.021 0.023 -0.610 0.188 0.189 0.615 0.121 0.120 0.615
4 0.8 0.6 0.2 0.993 0.065 0.067 -0.772 0.289 0.290 0.774 0.168 0.165 0.777
8 0.5 0.4 0.2 0.662 0.012 0.014 -0.359 0.105 0.106 0.363 0.070 0.070 0.359
8 0.8 0.4 0.2 0.662 0.033 0.036 -0.502 0.167 0.169 0.503 0.105 0.107 0.507
8 0.5 0.6 0.2 0.993 0.017 0.017 -0.535 0.153 0.153 0.542 0.096 0.097 0.538
8 0.8 0.6 0.2 0.993 0.041 0.042 -0.710 0.223 0.224 0.712 0.146 0.142 0.713
4 0.5 0.4 0.4 0.937 0.012 0.019 -0.684 0.170 0.171 0.686 0.109 0.107 0.689
4 0.8 0.4 0.4 0.937 0.051 0.066 -0.831 0.290 0.298 0.832 0.170 0.172 0.834
4 0.5 0.6 0.4 1.405 0.029 0.034 -0.995 0.227 0.231 0.998 0.144 0.147 0.998
4 0.8 0.6 0.4 1.405 0.062 0.067 -1.193 0.362 0.366 1.194 0.216 0.212 1.198
8 0.5 0.4 0.4 0.937 0.007 0.010 -0.602 0.133 0.135 0.605 0.082 0.083 0.605
8 0.8 0.4 0.4 0.937 0.040 0.048 -0.781 0.222 0.226 0.782 0.143 0.146 0.782
8 0.5 0.6 0.4 1.405 0.025 0.026 -0.905 0.198 0.199 0.910 0.129 0.126 0.909
8 0.8 0.6 0.4 1.405 0.050 0.051 -1.126 0.265 0.266 1.128 0.175 0.176 1.128
Note: $p$ denotes number of measurement items $Y_{j}$, $\pi_{Y}$ marginal proportion of $Y_{j}=1$,
         $R_{Y}^{2}$ and $R^{2}_{\eta}$ the $R^{2}$ statistics in models for $Y_{j}$ and $\eta_{2}$, and $\beta_{1}$ true value of $\beta_{1}$.
Results for 1-step and 2-step estimates exclude simulations where the estimation did not converge.
This occurred in 1–11 of the 1000 simulations in four settings.
Table A.10: Simulation results for standard error estimates of structural regression coefficient $\beta_{1}$ of a covariate $X$ with a skewed distribution for a latent response $\eta_{2}$ (Case B3).
Simulation standard deviation 2-step
vs. mean of est. std. error Coverage of without
2-step 1-step 95% conf. interval step-1 var.
$p$ $\pi_{Y}$ $R^{2}_{Y}$ $R^{2}_{\eta}$ $\beta_{1}$ s.d. m(se) s.d. m(se) 2-st. 1-st. 3-st. %var. cover.
$n=200$
4 0.5 0.4 0.0 0.000 0.226 0.179 0.232 0.179 97.9 97.5 94.4 92.2 94.5
4 0.8 0.4 0.0 0.000 0.434 0.254 0.539 0.281 99.4 98.9 96.0 86.7 96.6
4 0.5 0.6 0.0 0.000 0.278 0.247 0.281 0.247 97.8 97.8 94.6 95.3 95.3
4 0.8 0.6 0.0 0.000 0.344 0.300 0.347 0.299 98.3 98.1 95.2 92.9 96.2
8 0.5 0.4 0.0 0.000 0.173 0.154 0.174 0.154 97.1 97.1 95.0 95.5 94.9
8 0.8 0.4 0.0 0.000 0.235 0.193 0.239 0.193 98.2 98.0 94.6 92.3 94.3
8 0.5 0.6 0.0 0.000 0.244 0.220 0.243 0.219 96.9 96.9 95.7 97.0 95.5
8 0.8 0.6 0.0 0.000 0.295 0.250 0.296 0.250 98.0 97.8 95.6 95.8 95.8
4 0.5 0.4 0.2 0.662 0.378 0.332 0.413 0.340 94.1 94.6 10.4 62.4 87.8
4 0.8 0.4 0.2 0.662 0.984 0.958 1.081 0.852 93.8 94.5 5.1 64.7 90.6
4 0.5 0.6 0.2 0.993 0.563 0.482 0.579 0.486 95.7 96.0 15.2 71.0 91.4
4 0.8 0.6 0.2 0.993 1.001 0.814 1.051 0.860 95.4 96.1 5.5 77.9 93.6
8 0.5 0.4 0.2 0.662 0.275 0.257 0.282 0.259 95.1 95.2 21.9 66.2 88.6
8 0.8 0.4 0.2 0.662 0.490 0.427 0.521 0.437 93.7 94.1 4.6 70.4 89.6
8 0.5 0.6 0.2 0.993 0.408 0.387 0.412 0.388 96.3 96.1 25.9 74.6 92.6
8 0.8 0.6 0.2 0.993 0.658 0.563 0.670 0.566 95.2 95.1 7.8 80.1 93.5
4 0.5 0.4 0.4 0.937 0.516 0.456 0.569 0.470 94.7 95.6 4.8 48.7 83.3
4 0.8 0.4 0.4 0.937 1.904 1.195 3.403 2.629 91.3 93.1 3.6 54.1 83.9
4 0.5 0.6 0.4 1.405 0.622 0.584 0.649 0.592 95.0 96.0 3.6 62.8 89.4
4 0.8 0.6 0.4 1.405 1.240 1.093 1.367 1.100 95.1 95.1 1.9 70.4 91.1
8 0.5 0.4 0.4 0.937 0.338 0.323 0.353 0.330 94.7 94.8 6.5 52.9 84.5
8 0.8 0.4 0.4 0.937 0.681 0.583 0.771 0.618 93.4 94.4 1.0 57.6 85.9
8 0.5 0.6 0.4 1.405 0.512 0.476 0.519 0.480 95.3 95.1 6.0 65.0 88.2
8 0.8 0.6 0.4 1.405 0.767 0.710 0.781 0.718 95.7 95.9 1.0 72.6 92.4
$n=1000$
4 0.5 0.4 0.0 0.000 0.068 0.064 0.068 0.064 95.4 95.3 95.1 98.5 94.6
4 0.8 0.4 0.0 0.000 0.081 0.076 0.082 0.076 95.5 95.4 94.1 97.1 95.0
4 0.5 0.6 0.0 0.000 0.094 0.089 0.095 0.089 95.1 95.1 94.8 99.0 94.9
4 0.8 0.6 0.0 0.000 0.103 0.099 0.103 0.098 95.5 95.5 94.2 98.7 94.6
8 0.5 0.4 0.0 0.000 0.057 0.058 0.057 0.058 96.5 96.6 96.2 99.1 95.7
8 0.8 0.4 0.0 0.000 0.069 0.066 0.069 0.066 95.9 95.9 94.9 98.4 94.7
8 0.5 0.6 0.0 0.000 0.086 0.083 0.087 0.083 94.8 94.6 94.7 99.4 94.4
8 0.8 0.6 0.0 0.000 0.094 0.089 0.094 0.089 96.3 96.1 95.5 99.2 95.6
4 0.5 0.4 0.2 0.662 0.121 0.127 0.121 0.127 96.3 96.3 0.0 63.6 91.0
4 0.8 0.4 0.2 0.662 0.210 0.203 0.211 0.204 94.0 94.1 0.0 71.5 89.5
4 0.5 0.6 0.2 0.993 0.187 0.179 0.188 0.179 94.2 93.5 0.0 72.9 88.5
4 0.8 0.6 0.2 0.993 0.282 0.261 0.283 0.261 94.8 94.9 0.0 80.4 91.2
8 0.5 0.4 0.2 0.662 0.104 0.104 0.105 0.104 95.8 95.7 0.0 66.4 90.0
8 0.8 0.4 0.2 0.662 0.164 0.159 0.165 0.160 95.1 94.7 0.0 73.1 90.4
8 0.5 0.6 0.2 0.993 0.152 0.154 0.152 0.154 95.0 94.8 0.0 75.2 91.1
8 0.8 0.6 0.2 0.993 0.219 0.213 0.220 0.213 95.6 95.4 0.0 81.3 92.5
4 0.5 0.4 0.4 0.937 0.169 0.167 0.170 0.166 94.7 95.1 0.0 50.8 83.7
4 0.8 0.4 0.4 0.937 0.286 0.284 0.291 0.284 94.3 94.9 0.0 60.2 87.4
4 0.5 0.6 0.4 1.405 0.225 0.227 0.228 0.228 95.8 96.4 0.0 63.3 89.4
4 0.8 0.6 0.4 1.405 0.357 0.337 0.360 0.338 94.6 95.2 0.0 73.9 91.1
8 0.5 0.4 0.4 0.937 0.132 0.131 0.135 0.133 94.5 94.4 0.0 53.3 84.8
8 0.8 0.4 0.4 0.937 0.218 0.212 0.221 0.214 94.9 95.2 0.0 61.8 88.3
8 0.5 0.6 0.4 1.405 0.196 0.189 0.197 0.191 95.0 95.0 0.0 65.4 86.7
8 0.8 0.6 0.4 1.405 0.261 0.272 0.262 0.272 96.9 96.8 0.0 74.6 94.3
Note: $p$ denotes number of measurement items $Y_{j}$, $\pi_{Y}$ marginal proportion of $Y_{j}=1$,
         $R_{Y}^{2}$ and $R^{2}_{\eta}$ the $R^{2}$ statistics in models for $Y_{j}$ and $\eta_{2}$, and $\beta_{1}$ true value of $\beta_{1}$.
$\dagger$ Average % of the variance of 2-step estimator accounted for by step-2 variance alone,
       and coverage of 95% confidence interval if only this variance is included.
Results for 1-step and 2-step estimates exclude simulations where the estimation did not converge.
This occurred in 1–11 of the 1000 simulations in four settings.
Table A.11: Simulation results for point estimates of structural regression coefficient $\beta_{1}$ of a latent covariate $\eta_{1}$ for a conditionally normally distributed response $Z$ (Case C1).
Bias RMSE MAE
$p$ $\pi_{Y}$ $R^{2}_{Y}$ $R^{2}_{Z}$ $\beta_{1}$ 2-step 1-step 3-step 2-step 1-step 3-step 2-step 1-step 3-step
$n=200$
4 0.5 0.4 0.0 0.000 -0.003 -0.003 -0.003 0.067 0.070 0.067 0.041 0.042 0.040
4 0.8 0.4 0.0 0.000 -0.003 -0.003 -0.003 0.089 0.104 0.088 0.050 0.057 0.050
4 0.5 0.6 0.0 0.000 -0.000 -0.000 -0.000 0.039 0.040 0.039 0.024 0.025 0.024
4 0.8 0.6 0.0 0.000 -0.000 -0.000 -0.000 0.044 0.046 0.044 0.027 0.029 0.027
8 0.5 0.4 0.0 0.000 0.001 0.001 0.001 0.059 0.060 0.059 0.038 0.038 0.038
8 0.8 0.4 0.0 0.000 -0.002 -0.002 -0.002 0.068 0.071 0.068 0.042 0.043 0.042
8 0.5 0.6 0.0 0.000 0.002 0.002 0.002 0.038 0.038 0.037 0.025 0.025 0.025
8 0.8 0.6 0.0 0.000 0.001 0.001 0.001 0.040 0.040 0.040 0.025 0.025 0.025
4 0.5 0.4 0.2 0.302 0.003 0.012 0.003 0.111 0.113 0.111 0.072 0.071 0.071
4 0.8 0.4 0.2 0.302 0.008 0.027 0.008 0.159 0.159 0.156 0.088 0.087 0.090
4 0.5 0.6 0.2 0.201 -0.000 0.003 0.000 0.060 0.061 0.061 0.041 0.040 0.041
4 0.8 0.6 0.2 0.201 -0.005 -$0$0.004 0.065 0.063 0.066 0.044 0.042 0.044
8 0.5 0.4 0.2 0.302 -0.004 0.001 -0.004 0.083 0.085 0.083 0.056 0.057 0.056
8 0.8 0.4 0.2 0.302 0.005 0.018 0.006 0.114 0.119 0.114 0.070 0.068 0.069
8 0.5 0.6 0.2 0.201 0.000 0.001 0.000 0.051 0.051 0.051 0.035 0.034 0.035
8 0.8 0.6 0.2 0.201 -0.002 0.001 -0.002 0.055 0.056 0.056 0.037 0.038 0.037
4 0.5 0.4 0.4 0.427 0.002 0.012 0.003 0.134 0.129 0.134 0.083 0.082 0.084
4 0.8 0.4 0.4 0.427 -0.001 0.028 0.006 0.181 0.191 0.188 0.101 0.091 0.104
4 0.5 0.6 0.4 0.285 0.001 0.002 0.002 0.073 0.070 0.074 0.049 0.045 0.050
4 0.8 0.6 0.4 0.285 -0.005 0.003 -0.003 0.086 0.082 0.087 0.057 0.056 0.058
8 0.5 0.4 0.4 0.427 0.005 0.013 0.005 0.107 0.107 0.107 0.067 0.070 0.070
8 0.8 0.4 0.4 0.427 0.008 0.024 0.010 0.137 0.146 0.139 0.083 0.083 0.084
8 0.5 0.6 0.4 0.285 0.002 0.003 0.002 0.061 0.061 0.061 0.039 0.040 0.039
8 0.8 0.6 0.4 0.285 -0.003 0.002 -0.003 0.067 0.068 0.067 0.044 0.045 0.046
$n=1000$
4 0.5 0.4 0.0 0.000 -0.000 -0.000 -0.000 0.028 0.029 0.028 0.019 0.019 0.019
4 0.8 0.4 0.0 0.000 -0.001 -0.001 -0.001 0.034 0.034 0.034 0.022 0.022 0.022
4 0.5 0.6 0.0 0.000 0.000 0.000 0.000 0.017 0.018 0.017 0.012 0.012 0.012
4 0.8 0.6 0.0 0.000 -0.000 -0.000 -0.000 0.020 0.020 0.020 0.012 0.012 0.012
8 0.5 0.4 0.0 0.000 -0.001 -0.001 -0.001 0.026 0.026 0.025 0.016 0.016 0.016
8 0.8 0.4 0.0 0.000 0.002 0.002 0.002 0.030 0.030 0.030 0.019 0.020 0.020
8 0.5 0.6 0.0 0.000 0.001 0.001 0.001 0.016 0.016 0.016 0.011 0.011 0.011
8 0.8 0.6 0.0 0.000 0.000 0.000 0.000 0.017 0.017 0.017 0.011 0.011 0.011
4 0.5 0.4 0.2 0.302 0.003 0.003 0.002 0.046 0.044 0.046 0.030 0.029 0.030
4 0.8 0.4 0.2 0.302 0.006 0.008 0.007 0.060 0.057 0.061 0.041 0.039 0.040
4 0.5 0.6 0.2 0.201 -0.001 $0$0.001 0.026 0.025 0.026 0.017 0.017 0.017
4 0.8 0.6 0.2 0.201 -0.002 -0.001 -0.002 0.031 0.030 0.031 0.022 0.021 0.021
8 0.5 0.4 0.2 0.302 -0.000 0.001 -0.001 0.035 0.035 0.035 0.023 0.023 0.023
8 0.8 0.4 0.2 0.302 -0.000 0.003 -0.000 0.045 0.046 0.046 0.031 0.031 0.031
8 0.5 0.6 0.2 0.201 -0.002 -0.001 -0.002 0.022 0.022 0.022 0.015 0.015 0.015
8 0.8 0.6 0.2 0.201 0.000 0.001 -0.000 0.025 0.025 0.025 0.016 0.017 0.017
4 0.5 0.4 0.4 0.427 0.002 0.003 0.002 0.058 0.052 0.058 0.039 0.034 0.039
4 0.8 0.4 0.4 0.427 -0.001 0.002 -0.002 0.074 0.065 0.074 0.049 0.043 0.050
4 0.5 0.6 0.4 0.285 0.001 0.002 0.001 0.031 0.030 0.031 0.021 0.020 0.021
4 0.8 0.6 0.4 0.285 0.000 0.002 0.000 0.037 0.034 0.037 0.025 0.022 0.026
8 0.5 0.4 0.4 0.427 -0.001 0.001 -0.001 0.044 0.043 0.044 0.029 0.028 0.029
8 0.8 0.4 0.4 0.427 -0.001 0.003 -0.001 0.053 0.052 0.053 0.035 0.034 0.035
8 0.5 0.6 0.4 0.285 -$0$$0$0.000 0.026 0.026 0.026 0.017 0.017 0.018
8 0.8 0.6 0.4 0.285 0.000 0.001 0.000 0.028 0.029 0.029 0.019 0.019 0.020
Note: $p$ denotes number of measurement items $Y_{j}$, $\pi_{Y}$ marginal proportion of $Y_{j}=1$,
         $R_{Y}^{2}$ and $R^{2}_{Z}$ the $R^{2}$ statistics in models for $Y_{j}$ and $Z$, and $\beta_{1}$ true value of $\beta_{1}$.
Table A.12: Simulation results for standard error estimates of structural regression coefficient $\beta_{1}$ of a latent covariate $\eta_{1}$ for a conditionally normally distributed response $Z$ (Case C1).
Simulation standard deviation 2-step
vs. mean of est. std. error Coverage of without
2-step 1-step 95% conf. interval step-1 var.
$p$ $\pi_{Y}$ $R^{2}_{Y}$ $R^{2}_{Z}$ $\beta_{1}$ s.d. m(se) s.d. m(se) 2-st. 1-st. 3-st. %var. cover.
$n=200$
4 0.5 0.4 0.0 0.000 0.067 0.065 0.070 0.066 97.9 97.6 94.7 92.5 95.2
4 0.8 0.4 0.0 0.000 0.089 0.085 0.105 0.093 98.6 98.4 91.8 86.4 93.3
4 0.5 0.6 0.0 0.000 0.039 0.039 0.040 0.040 97.7 97.7 96.1 95.7 96.3
4 0.8 0.6 0.0 0.000 0.045 0.044 0.046 0.045 98.1 97.8 95.0 93.7 95.6
8 0.5 0.4 0.0 0.000 0.059 0.058 0.060 0.059 96.5 96.2 94.6 95.9 94.7
8 0.8 0.4 0.0 0.000 0.068 0.067 0.072 0.069 97.2 96.6 94.3 92.9 94.5
8 0.5 0.6 0.0 0.000 0.038 0.036 0.038 0.037 96.0 95.9 95.4 96.9 95.3
8 0.8 0.6 0.0 0.000 0.040 0.039 0.040 0.040 96.9 96.7 95.4 96.0 95.3
4 0.5 0.4 0.2 0.302 0.111 0.105 0.112 0.102 92.5 92.7 72.9 46.8 79.6
4 0.8 0.4 0.2 0.302 0.159 0.145 0.157 0.141 89.6 91.9 67.6 43.1 74.7
4 0.5 0.6 0.2 0.201 0.060 0.058 0.060 0.057 93.5 92.9 73.2 50.2 79.4
4 0.8 0.6 0.2 0.201 0.065 0.066 0.063 0.064 93.6 93.9 75.7 47.8 82.0
8 0.5 0.4 0.2 0.302 0.083 0.080 0.085 0.081 92.1 92.5 76.9 48.1 80.0
8 0.8 0.4 0.2 0.302 0.113 0.105 0.118 0.107 91.1 92.3 72.7 42.1 77.1
8 0.5 0.6 0.2 0.201 0.051 0.048 0.051 0.048 93.5 93.7 77.7 51.8 80.6
8 0.8 0.6 0.2 0.201 0.055 0.055 0.056 0.054 93.0 93.2 77.0 48.1 80.4
4 0.5 0.4 0.4 0.427 0.134 0.131 0.128 0.119 92.7 92.7 59.9 33.4 73.5
4 0.8 0.4 0.4 0.427 0.181 0.182 0.189 0.163 93.5 94.6 58.4 31.4 73.9
4 0.5 0.6 0.4 0.285 0.073 0.072 0.070 0.067 94.1 93.9 60.8 35.9 74.5
4 0.8 0.6 0.4 0.285 0.086 0.083 0.082 0.078 93.4 93.7 59.0 34.1 71.9
8 0.5 0.4 0.4 0.427 0.107 0.101 0.106 0.100 93.5 94.2 63.5 30.3 69.7
8 0.8 0.4 0.4 0.427 0.137 0.133 0.144 0.131 92.8 93.9 61.0 26.8 67.9
8 0.5 0.6 0.4 0.285 0.061 0.059 0.061 0.058 93.9 93.7 65.9 33.3 71.0
8 0.8 0.6 0.4 0.285 0.067 0.067 0.068 0.066 93.8 93.5 65.6 30.8 71.2
$n=1000$
4 0.5 0.4 0.0 0.000 0.028 0.028 0.029 0.028 95.8 95.5 94.9 98.4 95.1
4 0.8 0.4 0.0 0.000 0.034 0.033 0.034 0.033 96.0 95.9 95.1 97.3 95.1
4 0.5 0.6 0.0 0.000 0.017 0.017 0.018 0.017 95.2 95.2 94.9 99.1 95.0
4 0.8 0.6 0.0 0.000 0.020 0.019 0.020 0.019 95.4 95.4 94.6 98.7 94.7
8 0.5 0.4 0.0 0.000 0.026 0.025 0.026 0.025 95.0 95.0 94.8 99.1 94.8
8 0.8 0.4 0.0 0.000 0.030 0.028 0.030 0.028 93.9 93.8 93.3 98.4 93.3
8 0.5 0.6 0.0 0.000 0.016 0.016 0.016 0.016 94.5 94.5 94.3 99.4 94.1
8 0.8 0.6 0.0 0.000 0.017 0.017 0.017 0.017 95.8 95.8 95.1 99.2 95.3
4 0.5 0.4 0.2 0.302 0.046 0.045 0.043 0.043 94.7 93.4 74.7 46.9 82.1
4 0.8 0.4 0.2 0.302 0.060 0.058 0.057 0.055 93.8 94.2 70.5 43.9 78.2
4 0.5 0.6 0.2 0.201 0.026 0.025 0.025 0.025 93.1 92.8 76.4 50.5 82.2
4 0.8 0.6 0.2 0.201 0.031 0.029 0.030 0.028 92.5 92.6 73.2 48.4 78.4
8 0.5 0.4 0.2 0.302 0.035 0.035 0.035 0.035 95.2 95.6 80.5 47.5 83.8
8 0.8 0.4 0.2 0.302 0.045 0.044 0.046 0.044 94.4 94.4 73.7 42.8 78.0
8 0.5 0.6 0.2 0.201 0.022 0.021 0.022 0.021 94.4 94.1 79.8 52.2 82.6
8 0.8 0.6 0.2 0.201 0.025 0.024 0.025 0.024 95.3 95.2 77.9 48.7 81.5
4 0.5 0.4 0.4 0.427 0.058 0.057 0.052 0.051 94.0 94.9 60.9 33.7 74.6
4 0.8 0.4 0.4 0.427 0.074 0.074 0.065 0.064 93.4 94.0 55.1 32.2 74.0
4 0.5 0.6 0.4 0.285 0.031 0.032 0.030 0.030 95.9 95.3 64.1 35.7 77.9
4 0.8 0.6 0.4 0.285 0.037 0.036 0.034 0.034 94.2 94.8 61.9 34.7 75.7
8 0.5 0.4 0.4 0.427 0.044 0.044 0.043 0.043 93.9 93.9 65.5 30.4 71.3
8 0.8 0.4 0.4 0.427 0.053 0.055 0.052 0.053 95.4 95.8 62.7 27.8 70.3
8 0.5 0.6 0.4 0.285 0.026 0.026 0.026 0.025 95.3 94.9 66.8 33.7 72.0
8 0.8 0.6 0.4 0.285 0.028 0.030 0.029 0.029 95.6 94.9 68.7 31.5 74.7
Note: $p$ denotes number of measurement items $Y_{j}$, $\pi_{Y}$ marginal proportion of $Y_{j}=1$,
         $R_{Y}^{2}$ and $R^{2}_{Z}$ the $R^{2}$ statistics in models for $Y_{j}$ and $Z$, and $\beta_{1}$ true value of $\beta_{1}$.
$\dagger$ Average % of the variance of 2-step estimator accounted for by step-2 variance alone,
       and coverage of 95% confidence interval if only this variance is included.
Table A.13: Simulation results for point estimates of structural regression coefficient $\beta_{1}$ of a latent covariate $\eta_{1}$ for a conditionally skew-normally distributed response $Z$ (Case C2).
Bias RMSE MAE
$p$ $\pi_{Y}$ $R^{2}_{Y}$ $R^{2}_{Z}$ $\beta_{1}$ 2-step 1-step 3-step 2-step 1-step 3-step 2-step 1-step 3-step
$n=200$
4 0.5 0.4 0.0 0.000 0.003 0.003 0.003 0.068 0.070 0.067 0.040 0.042 0.040
4 0.8 0.4 0.0 0.000 0.005 0.008 0.003 0.085 0.113 0.084 0.043 0.048 0.043
4 0.5 0.6 0.0 0.000 0.000 0.000 0.000 0.039 0.040 0.039 0.025 0.026 0.025
4 0.8 0.6 0.0 0.000 0.001 0.001 -0.000 0.044 0.046 0.044 0.026 0.027 0.026
8 0.5 0.4 0.0 0.000 0.002 0.002 0.002 0.056 0.057 0.056 0.037 0.038 0.037
8 0.8 0.4 0.0 0.000 0.001 0.001 -0.000 0.064 0.067 0.064 0.041 0.043 0.041
8 0.5 0.6 0.0 0.000 -0.000 -0.000 -0.000 0.036 0.036 0.036 0.024 0.025 0.024
8 0.8 0.6 0.0 0.000 0.004 0.004 0.003 0.040 0.040 0.040 0.026 0.026 0.026
4 0.5 0.4 0.2 0.302 0.007 0.016 0.007 0.114 0.113 0.113 0.071 0.067 0.070
4 0.8 0.4 0.2 0.302 0.022 0.043 0.008 0.173 0.193 0.160 0.093 0.090 0.089
4 0.5 0.6 0.2 0.201 -0.002 0.001 -0.002 0.059 0.057 0.059 0.038 0.036 0.038
4 0.8 0.6 0.2 0.201 0.008 0.014 0.001 0.076 0.076 0.071 0.048 0.047 0.048
8 0.5 0.4 0.2 0.302 0.004 0.009 0.004 0.086 0.088 0.085 0.057 0.058 0.057
8 0.8 0.4 0.2 0.302 0.019 0.035 0.010 0.119 0.128 0.111 0.069 0.069 0.066
8 0.5 0.6 0.2 0.201 -0.001 0.001 -0.001 0.050 0.050 0.050 0.034 0.034 0.034
8 0.8 0.6 0.2 0.201 0.005 0.010 0.000 0.059 0.061 0.055 0.037 0.037 0.035
4 0.5 0.4 0.4 0.427 -0.000 0.014 -0.000 0.142 0.133 0.142 0.088 0.082 0.087
4 0.8 0.4 0.4 0.427 0.020 0.052 0.009 0.193 0.206 0.194 0.107 0.097 0.109
4 0.5 0.6 0.4 0.285 -0.003 0.002 -0.003 0.072 0.068 0.072 0.050 0.046 0.049
4 0.8 0.6 0.4 0.285 0.006 0.015 -0.001 0.089 0.088 0.087 0.056 0.053 0.055
8 0.5 0.4 0.4 0.427 0.003 0.012 0.002 0.112 0.112 0.111 0.071 0.074 0.071
8 0.8 0.4 0.4 0.427 0.020 0.034 0.010 0.145 0.150 0.140 0.087 0.083 0.085
8 0.5 0.6 0.4 0.285 0.001 0.003 0.001 0.062 0.063 0.062 0.040 0.040 0.040
8 0.8 0.6 0.4 0.285 0.003 0.009 -0.002 0.069 0.070 0.067 0.046 0.046 0.045
$n=1000$
4 0.5 0.4 0.0 0.000 0.001 0.001 0.001 0.029 0.029 0.029 0.019 0.019 0.019
4 0.8 0.4 0.0 0.000 0.000 0.000 -0.000 0.033 0.033 0.032 0.021 0.021 0.021
4 0.5 0.6 0.0 0.000 0.000 0.000 0.000 0.018 0.018 0.018 0.012 0.012 0.012
4 0.8 0.6 0.0 0.000 0.001 0.001 0.001 0.018 0.018 0.018 0.013 0.013 0.013
8 0.5 0.4 0.0 0.000 -0.000 -0.000 -0.000 0.024 0.025 0.024 0.016 0.016 0.016
8 0.8 0.4 0.0 0.000 0.001 0.001 0.001 0.028 0.029 0.028 0.018 0.019 0.019
8 0.5 0.6 0.0 0.000 0.000 0.000 0.000 0.017 0.017 0.017 0.011 0.011 0.011
8 0.8 0.6 0.0 0.000 0.001 0.001 0.001 0.017 0.017 0.017 0.011 0.011 0.011
4 0.5 0.4 0.2 0.302 0.001 0.002 0.000 0.045 0.043 0.045 0.030 0.029 0.030
4 0.8 0.4 0.2 0.302 0.017 0.021 0.003 0.065 0.064 0.058 0.040 0.037 0.038
4 0.5 0.6 0.2 0.201 0.000 0.002 0.000 0.026 0.025 0.026 0.017 0.016 0.017
4 0.8 0.6 0.2 0.201 0.007 0.007 0.000 0.033 0.032 0.030 0.021 0.021 0.020
8 0.5 0.4 0.2 0.302 0.001 0.003 0.001 0.035 0.036 0.035 0.023 0.023 0.023
8 0.8 0.4 0.2 0.302 0.008 0.013 -0.000 0.048 0.050 0.045 0.030 0.032 0.029
8 0.5 0.6 0.2 0.201 -0.001 -0.001 -0.001 0.022 0.022 0.022 0.015 0.015 0.014
8 0.8 0.6 0.2 0.201 0.004 0.005 -0.000 0.025 0.025 0.024 0.017 0.017 0.016
4 0.5 0.4 0.4 0.427 0.005 0.007 0.004 0.059 0.053 0.058 0.039 0.033 0.038
4 0.8 0.4 0.4 0.427 0.025 0.029 0.004 0.084 0.076 0.076 0.051 0.047 0.049
4 0.5 0.6 0.4 0.285 0.001 0.001 0.000 0.032 0.031 0.032 0.022 0.021 0.022
4 0.8 0.6 0.4 0.285 0.011 0.014 0.002 0.040 0.039 0.037 0.025 0.026 0.023
8 0.5 0.4 0.4 0.427 -0.000 0.002 -0.000 0.044 0.042 0.044 0.029 0.028 0.029
8 0.8 0.4 0.4 0.427 0.012 0.015 0.001 0.059 0.057 0.055 0.038 0.035 0.035
8 0.5 0.6 0.4 0.285 -0.001 -0.001 -0.001 0.025 0.025 0.025 0.017 0.017 0.017
8 0.8 0.6 0.4 0.285 0.006 0.007 -0.000 0.030 0.031 0.029 0.019 0.019 0.019
Note: $p$ denotes number of measurement items $Y_{j}$, $\pi_{Y}$ marginal proportion of $Y_{j}=1$,
         $R_{Y}^{2}$ and $R^{2}_{Z}$ the $R^{2}$ statistics in models for $Y_{j}$ and $Z$, and $\beta_{1}$ true value of $\beta_{1}$.
Table A.14: Simulation results for standard error estimates of structural regression coefficient $\beta_{1}$ of a latent covariate $\eta_{1}$ for a conditionally skew-normally distributed response $Z$ (Case C2).
Simulation standard deviation 2-step
vs. mean of est. std. error Coverage of without
2-step 1-step 95% conf. interval step-1 var.
$p$ $\pi_{Y}$ $R^{2}_{Y}$ $R^{2}_{Z}$ $\beta_{1}$ s.d. m(se) s.d. m(se) 2-st. 1-st. 3-st. %var. cover.
$n=200$
4 0.5 0.4 0.0 0.000 0.068 0.066 0.070 0.067 97.9 97.6 95.7 93.1 95.9
4 0.8 0.4 0.0 0.000 0.085 0.085 0.113 0.101 99.4 99.2 95.7 87.8 96.6
4 0.5 0.6 0.0 0.000 0.039 0.039 0.040 0.040 96.5 96.3 94.9 95.6 95.0
4 0.8 0.6 0.0 0.000 0.044 0.044 0.046 0.046 97.2 97.0 94.3 93.7 95.0
8 0.5 0.4 0.0 0.000 0.056 0.057 0.057 0.058 96.7 96.6 94.8 95.8 95.1
8 0.8 0.4 0.0 0.000 0.064 0.067 0.067 0.069 98.8 98.3 95.6 93.1 95.8
8 0.5 0.6 0.0 0.000 0.036 0.037 0.036 0.037 97.5 97.3 96.5 97.2 96.5
8 0.8 0.6 0.0 0.000 0.040 0.039 0.040 0.040 96.8 96.5 95.2 96.0 94.9
4 0.5 0.4 0.2 0.302 0.113 0.106 0.112 0.103 93.1 92.6 72.9 46.2 80.1
4 0.8 0.4 0.2 0.302 0.172 0.705 0.188 0.150 91.5 93.7 69.4 42.8 78.9
4 0.5 0.6 0.2 0.201 0.059 0.057 0.057 0.056 92.6 94.0 75.3 49.9 80.1
4 0.8 0.6 0.2 0.201 0.075 0.069 0.075 0.068 93.2 93.3 73.6 47.6 79.3
8 0.5 0.4 0.2 0.302 0.086 0.082 0.088 0.083 93.6 93.9 76.5 47.5 79.7
8 0.8 0.4 0.2 0.302 0.117 0.109 0.123 0.112 93.7 95.1 73.5 41.4 77.4
8 0.5 0.6 0.2 0.201 0.050 0.048 0.050 0.048 93.0 93.0 79.5 52.2 81.9
8 0.8 0.6 0.2 0.201 0.059 0.056 0.060 0.056 93.7 93.7 78.6 47.5 80.3
4 0.5 0.4 0.4 0.427 0.142 0.132 0.132 0.119 91.1 92.6 57.6 32.7 70.2
4 0.8 0.4 0.4 0.427 0.193 0.192 0.199 0.170 94.4 94.8 57.8 30.5 74.1
4 0.5 0.6 0.4 0.285 0.072 0.071 0.068 0.068 93.2 94.7 62.5 35.6 75.2
4 0.8 0.6 0.4 0.285 0.089 0.085 0.087 0.080 94.5 94.0 61.9 33.7 74.3
8 0.5 0.4 0.4 0.427 0.112 0.101 0.112 0.100 92.7 92.9 61.5 30.1 68.0
8 0.8 0.4 0.4 0.427 0.143 0.136 0.146 0.133 93.4 95.3 60.5 26.3 67.6
8 0.5 0.6 0.4 0.285 0.062 0.059 0.063 0.058 92.9 93.1 65.7 33.1 69.9
8 0.8 0.6 0.4 0.285 0.069 0.068 0.069 0.067 93.4 94.1 65.3 30.1 70.9
$n=1000$
4 0.5 0.4 0.0 0.000 0.029 0.028 0.029 0.028 95.1 95.0 94.0 98.4 94.1
4 0.8 0.4 0.0 0.000 0.033 0.033 0.033 0.033 95.8 95.7 94.3 97.4 94.5
4 0.5 0.6 0.0 0.000 0.018 0.017 0.018 0.017 95.0 95.0 94.7 99.0 94.7
4 0.8 0.6 0.0 0.000 0.018 0.019 0.018 0.019 96.5 96.5 96.3 98.8 96.4
8 0.5 0.4 0.0 0.000 0.024 0.025 0.025 0.025 96.3 96.1 95.7 99.2 95.7
8 0.8 0.4 0.0 0.000 0.028 0.028 0.029 0.029 95.3 95.3 95.0 98.5 94.9
8 0.5 0.6 0.0 0.000 0.017 0.016 0.017 0.016 94.3 94.2 94.1 99.4 94.1
8 0.8 0.6 0.0 0.000 0.017 0.017 0.017 0.017 96.6 96.6 96.5 99.2 96.3
4 0.5 0.4 0.2 0.302 0.045 0.045 0.043 0.043 94.8 95.8 74.9 46.8 83.3
4 0.8 0.4 0.2 0.302 0.063 0.060 0.060 0.057 94.5 94.5 70.9 43.2 78.4
4 0.5 0.6 0.2 0.201 0.026 0.025 0.025 0.025 95.1 94.5 76.2 50.5 82.0
4 0.8 0.6 0.2 0.201 0.032 0.030 0.031 0.029 94.6 94.9 75.1 48.2 78.9
8 0.5 0.4 0.2 0.302 0.035 0.036 0.036 0.036 94.6 94.6 79.5 47.6 82.1
8 0.8 0.4 0.2 0.302 0.047 0.045 0.049 0.045 94.8 94.5 74.1 42.3 77.9
8 0.5 0.6 0.2 0.201 0.022 0.021 0.022 0.021 94.1 93.5 79.2 52.0 81.8
8 0.8 0.6 0.2 0.201 0.025 0.024 0.025 0.024 94.9 94.6 79.8 48.3 81.4
4 0.5 0.4 0.4 0.427 0.059 0.057 0.053 0.051 95.5 94.8 59.1 33.7 74.5
4 0.8 0.4 0.4 0.427 0.081 0.078 0.071 0.067 95.7 95.3 56.7 31.8 72.1
4 0.5 0.6 0.4 0.285 0.032 0.032 0.031 0.030 95.0 94.3 61.0 35.8 74.5
4 0.8 0.6 0.4 0.285 0.038 0.038 0.037 0.035 94.3 93.3 64.3 34.0 75.7
8 0.5 0.4 0.4 0.427 0.044 0.044 0.042 0.043 94.1 94.7 67.6 30.4 72.4
8 0.8 0.4 0.4 0.427 0.057 0.057 0.055 0.054 95.0 95.5 61.8 27.0 69.1
8 0.5 0.6 0.4 0.285 0.025 0.026 0.025 0.025 94.8 94.6 69.7 33.7 76.5
8 0.8 0.6 0.4 0.285 0.030 0.030 0.030 0.029 95.5 94.1 69.5 30.8 74.5
Note: $p$ denotes number of measurement items $Y_{j}$, $\pi_{Y}$ marginal proportion of $Y_{j}=1$,
         $R_{Y}^{2}$ and $R^{2}_{Z}$ the $R^{2}$ statistics in models for $Y_{j}$ and $Z$, and $\beta_{1}$ true value of $\beta_{1}$.
$\dagger$ Average % of the variance of 2-step estimator accounted for by step-2 variance alone,
       and coverage of 95% confidence interval if only this variance is included.
Table A.15: Simulation results for point estimates of structural regression coefficient $\beta_{1}$ of latent covariate $\eta_{1}$ for a conditionally normally distributed latent response $\eta_{2}$, when each latent variable is measured by 4 measurement items with 4 ordered categories (Case D).
Bias RMSE MAE
$R^{2}_{Y}$ $R^{2}_{\eta}$ $\beta_{1}$ 2-step 1-step 3-step 2-step 1-step 3-step 2-step 1-step 3-step
$n=200$
0.4 0.0 0.000 0.001 0.001 0.001 0.110 0.113 0.079 0.068 0.070 0.048
0.6 0.0 0.000 -0.001 0.005 -0.001 0.096 0.204 0.079 0.063 0.063 0.051
0.4 0.2 0.447 0.003 0.014 -0.125 0.152 0.154 0.169 0.096 0.093 0.146
0.6 0.2 0.447 0.002 0.006 -0.078 0.129 0.130 0.134 0.087 0.085 0.106
0.4 0.4 0.632 0.003 0.018 -0.179 0.198 0.192 0.233 0.126 0.120 0.212
0.6 0.4 0.632 0.010 0.016 -0.103 0.157 0.154 0.169 0.102 0.098 0.134
$n=1000$
0.4 0.0 0.000 0.000 0.000 0.000 0.046 0.046 0.033 0.031 0.031 0.022
0.6 0.0 0.000 0.000 0.000 0.000 0.039 0.039 0.032 0.027 0.027 0.022
0.4 0.2 0.447 -0.001 0.002 -0.132 0.065 0.065 0.140 0.043 0.043 0.135
0.6 0.2 0.447 0.000 0.002 -0.081 0.055 0.055 0.094 0.036 0.036 0.084
0.4 0.4 0.632 -0.005 -0.001 -0.189 0.080 0.079 0.198 0.054 0.051 0.195
0.6 0.4 0.632 0.002 0.004 -0.112 0.065 0.062 0.125 0.045 0.044 0.113
Note: $R_{Y}^{2}$ and $R^{2}_{\eta}$ denote the $R^{2}$ statistics in models for $Y_{j}$ and $\eta_{2}$, and $\beta_{1}$ true value of $\beta_{1}$.
Table A.16: Simulation results for standard error estimates of the estimates considered in Table A.15.
Simulation standard deviation 2-step
vs. mean of est. std. error Coverage of without
2-step 1-step 95% conf. interval step-1 var.
$R^{2}_{Y}$ $R^{2}_{\eta}$ $\beta_{1}$ s.d. m(se) s.d. m(se) 2-st. 1-st. 3-st. %var. cover.
$n=200$
0.4 0.0 0.000 0.110 0.107 0.113 0.108 97.2 97.1 95.0 92.7 95.0
0.6 0.0 0.000 0.096 0.092 0.204 0.111 96.8 96.6 95.3 95.5 95.3
0.4 0.2 0.447 0.152 0.152 0.154 0.151 93.1 93.6 46.1 45.8 80.9
0.6 0.2 0.447 0.130 0.124 0.130 0.123 93.3 93.5 60.1 48.5 80.5
0.4 0.4 0.632 0.198 0.188 0.191 0.180 92.2 93.0 31.5 28.4 68.4
0.6 0.4 0.632 0.157 0.153 0.153 0.148 93.3 94.0 46.9 30.1 70.7
$n=1000$
0.4 0.0 0.000 0.046 0.046 0.047 0.046 95.3 95.3 95.2 98.4 95.2
0.6 0.0 0.000 0.039 0.039 0.039 0.039 95.1 95.1 95.0 99.1 95.0
0.4 0.2 0.447 0.065 0.066 0.065 0.065 95.0 95.4 9.5 45.6 81.3
0.6 0.2 0.447 0.055 0.054 0.055 0.054 94.3 94.5 30.1 48.9 82.4
0.4 0.4 0.632 0.080 0.081 0.079 0.077 95.6 95.3 3.2 28.9 70.9
0.6 0.4 0.632 0.065 0.066 0.062 0.064 95.2 96.3 15.5 30.5 72.2
$\dagger$ Average % of the variance of 2-step estimator accounted for by step-2 variance alone,
       and coverage of 95% confidence interval if only this variance is included.
Table A.17: Simulation results for point estimates of structural regression coefficient $\beta_{2}$ of latent covariate $\eta_{2}$ for a latent response $\eta_{3}$, in a model which also includes another latent covariate $\eta_{1}$ (which has coefficient $\beta_{1}$) (Case E).
Bias RMSE MAE
$R^{2}_{Y}$ $\beta_{1}$ $\beta_{2}$ 2-step 1-step 3-step 2-step 1-step 3-step 2-step 1-step 3-step
$n=100$
0.4 0.390 0.309 0.036 0.044 -0.037 0.451 0.756 0.235 0.198 0.214 0.150
0.6 0.390 0.309 0.034 0.036 -0.021 0.302 0.326 0.195 0.158 0.161 0.131
0.4 0.273 0.423 0.007 0.049 -0.120 0.524 0.623 0.291 0.230 0.249 0.223
0.6 0.273 0.423 0.053 0.074 -0.068 0.341 0.427 0.227 0.186 0.195 0.165
0.4 0.000 0.632 0.026 0.154 -0.257 0.514 0.954 0.367 0.284 0.302 0.336
0.6 0.000 0.632 0.057 0.108 -0.178 0.427 0.486 0.319 0.226 0.232 0.261
$n=200$
0.4 0.390 0.309 0.018 -0.019 -0.056 0.276 0.962 0.167 0.144 0.152 0.120
0.6 0.390 0.309 0.019 -0.017 -0.036 0.184 1.183 0.124 0.109 0.114 0.089
0.4 0.273 0.423 0.024 0.039 -0.127 0.260 0.272 0.193 0.161 0.151 0.164
0.6 0.273 0.423 0.010 0.005 -0.099 0.212 0.487 0.170 0.127 0.133 0.135
0.4 0.000 0.632 0.024 0.065 -0.270 0.346 0.350 0.323 0.211 0.195 0.311
0.6 0.000 0.632 0.011 0.039 -0.213 0.258 0.490 0.266 0.163 0.158 0.236
$n=1000$
0.4 0.390 0.309 0.003 0.004 -0.073 0.093 0.092 0.089 0.059 0.058 0.078
0.6 0.390 0.309 0.002 -0.245 -0.048 0.077 3.529 0.070 0.050 0.052 0.057
0.4 0.273 0.423 -0.001 -0.092 -0.148 0.106 2.134 0.159 0.070 0.064 0.154
0.6 0.273 0.423 0.002 -0.354 -0.107 0.085 5.133 0.120 0.058 0.059 0.110
0.4 0.000 0.632 0.007 0.012 -0.288 0.141 0.133 0.297 0.094 0.091 0.292
0.6 0.000 0.632 0.002 -0.567 -0.223 0.110 8.196 0.233 0.072 0.070 0.227
Note: $R_{Y}^{2}$ denotes the $R^{2}$ statistic in measurement models for items $Y_{j}$.
         Results for 1-step estimates exclude simulations where the estimation did not converge.
         This occurred in 0–9 of the 1000 simulations in different settings.
Table A.18: Simulation results for point estimates of structural regression coefficient $\beta_{2}$ of latent covariate $\eta_{2}$ for a latent response $\eta_{3}$, in a model which also includes another latent covariate $\eta_{1}$ (which has coefficient $\beta_{1}$) (Case E). Note: Here bias of the estimates is quantified by the difference between their median over the simulations and the true value of $\beta_{2}$, whereas in Table A.17 the simulation mean rather than the median was used for this. Other columns of this table identical to those of Table A.17.
Bias (median) RMSE MAE
$R^{2}_{Y}$ $\beta_{1}$ $\beta_{2}$ 2-step 1-step 3-step 2-step 1-step 3-step 2-step 1-step 3-step
$n=100$
0.4 0.390 0.309 -0.045 -0.033 -0.097 0.451 0.756 0.235 0.198 0.214 0.150
0.6 0.390 0.309 -0.020 -0.015 -0.066 0.302 0.326 0.195 0.158 0.161 0.131
0.4 0.273 0.423 -0.089 -0.059 -0.183 0.524 0.623 0.291 0.230 0.249 0.223
0.6 0.273 0.423 -0.006 -0.002 -0.109 0.341 0.427 0.227 0.186 0.195 0.165
0.4 0.000 0.632 -0.090 $0$0.317 0.514 0.954 0.367 0.284 0.302 0.336
0.6 0.000 0.632 -0.035 0.011 -0.230 0.427 0.486 0.319 0.226 0.232 0.261
$n=200$
0.4 0.390 0.309 -0.025 -0.013 -0.091 0.276 0.962 0.167 0.144 0.152 0.120
0.6 0.390 0.309 -0.003 -0.004 -0.057 0.184 1.183 0.124 0.109 0.114 0.089
0.4 0.273 0.423 -0.021 -0.009 -0.151 0.260 0.272 0.193 0.161 0.151 0.164
0.6 0.273 0.423 -0.022 -0.011 -0.120 0.212 0.487 0.170 0.127 0.133 0.135
0.4 0.000 0.632 -0.050 -0.008 -0.304 0.346 0.350 0.323 0.211 0.195 0.311
0.6 0.000 0.632 -0.026 -0.019 -0.233 0.258 0.490 0.266 0.163 0.158 0.236
$n=1000$
0.4 0.390 0.309 -0.002 -0.003 -0.077 0.093 0.092 0.089 0.059 0.058 0.078
0.6 0.390 0.309 -0.005 -0.005 -0.054 0.077 3.529 0.070 0.050 0.052 0.057
0.4 0.273 0.423 -0.012 -0.008 -0.154 0.106 2.134 0.159 0.070 0.064 0.154
0.6 0.273 0.423 -0.002 0.003 -0.110 0.085 5.133 0.120 0.058 0.059 0.110
0.4 0.000 0.632 -0.007 0.001 -0.292 0.141 0.133 0.297 0.094 0.091 0.292
0.6 0.000 0.632 -0.004 -0.002 -0.227 0.110 8.196 0.233 0.072 0.070 0.227
Note: $R_{Y}^{2}$ denotes the $R^{2}$ statistic in measurement models for items $Y_{j}$.
         Results for 1-step estimates exclude simulations where the estimation did not converge.
         This occurred in 0–9 of the 1000 simulations in different settings.
Table A.19: Simulation results for standard error estimates of structural regression coefficient $\beta_{2}$ of latent covariate $\eta_{2}$ for a latent response $\eta_{3}$, in a model which also includes another latent covariate $\eta_{1}$ (which has coefficient $\beta_{1}$) (Case E).
Simulation standard deviation 2-step
vs. mean of est. std. error Coverage of without
2-step 1-step 95% conf. interval step-1 var.
$R^{2}_{Y}$ $\beta_{1}$ $\beta_{2}$ s.d. m(se) s.d. m(se) 2-st. 1-st. 3-st. %var. cover.
$n=100$
0.4 0.390 0.309 0.450 1.937 0.755 2.350 93.2 94.2 63.8 58.9 87.8
0.6 0.390 0.309 0.300 0.299 0.324 0.303 93.6 93.1 68.7 64.1 89.2
0.4 0.273 0.423 0.524 0.581 0.621 0.513 88.6 91.1 49.2 53.4 80.3
0.6 0.273 0.423 0.337 0.332 0.421 0.343 92.7 92.7 60.1 56.3 85.9
0.4 0.000 0.632 0.514 1.893 0.942 0.657 88.4 90.3 34.9 42.9 76.8
0.6 0.000 0.632 0.423 0.427 0.474 0.429 89.4 91.1 44.2 43.8 79.4
$n=200$
0.4 0.390 0.309 0.276 0.255 0.962 0.486 95.2 95.2 56.6 64.2 89.3
0.6 0.390 0.309 0.183 0.181 1.184 0.304 94.0 94.3 70.3 69.7 89.9
0.4 0.273 0.423 0.259 0.277 0.269 0.268 93.3 94.5 45.8 56.4 86.0
0.6 0.273 0.423 0.212 0.203 0.487 0.225 91.6 91.9 52.0 59.7 84.3
0.4 0.000 0.632 0.345 0.354 0.344 0.330 91.0 93.6 22.8 44.0 79.8
0.6 0.000 0.632 0.258 0.256 0.488 0.368 92.6 93.1 29.1 45.4 80.8
$n=1000$
0.4 0.390 0.309 0.093 0.095 0.092 0.092 95.4 94.8 36.7 73.1 90.8
0.6 0.390 0.309 0.077 0.074 3.523 0.134 94.7 94.5 54.9 74.6 89.7
0.4 0.273 0.423 0.106 0.106 2.133 0.146 93.4 93.2 10.8 62.0 87.4
0.6 0.273 0.423 0.085 0.083 5.123 0.155 93.8 93.8 22.6 62.3 87.1
0.4 0.000 0.632 0.141 0.137 0.132 0.127 93.2 94.2 0.9 47.2 81.0
0.6 0.000 0.632 0.110 0.107 8.181 0.195 94.0 94.6 2.6 46.5 80.6
$\dagger$ Average % of the variance of 2-step estimator accounted for by step-2 variance alone,
       and coverage of 95% confidence interval if only this variance is included.
Results for 1-step estimates exclude simulations where the estimation did not converge.
This occurred in 0–9 of the 1000 simulations in different settings.
Table A.20: Simulation results for point estimates of structural regression coefficient $\beta_{2}$ of latent covariate $\eta_{2}$ for a latent response $\eta_{3}$, in a model which also includes another latent covariate $\eta_{1}$ (which has coefficient $\beta_{1}$) and where the measurement model includes cross-loadings from $\eta_{1}$ to measures of $\eta_{2}$ (Case F).
Bias RMSE MAE
$R^{2}_{Y}$ $\beta_{1}$ $\beta_{2}$ 2-step 1-step 3-step 2-step 1-step 3-step 2-step 1-step 3-step
$n=200$
0.4 0.390 0.309 0.019 0.073 -0.099 0.341 0.935 0.230 0.158 0.176 0.161
0.6 0.390 0.309 0.025 0.046 -0.067 0.216 0.346 0.177 0.123 0.127 0.127
0.4 0.273 0.423 0.026 0.068 -0.134 0.470 0.921 0.303 0.173 0.181 0.196
0.6 0.273 0.423 0.007 -0.202 -0.108 0.231 5.275 0.207 0.139 0.147 0.163
0.4 0.000 0.632 0.033 0.152 0.014 0.528 0.700 7.030 0.236 0.234 0.300
0.6 0.000 0.632 0.009 0.044 -0.159 0.291 0.313 0.272 0.176 0.177 0.226
$n=1000$
0.4 0.390 0.309 0.002 0.010 -0.116 0.192 0.112 0.135 0.072 0.070 0.125
0.6 0.390 0.309 0.002 -0.046 -0.087 0.082 1.613 0.106 0.055 0.055 0.093
0.4 0.273 0.423 -0.012 0.005 -0.163 0.230 0.124 0.181 0.081 0.083 0.173
0.6 0.273 0.423 0.001 0.006 -0.115 0.091 0.092 0.134 0.059 0.058 0.121
0.4 0.000 0.632 0.007 0.017 -0.239 0.160 0.155 0.260 0.099 0.095 0.251
0.6 0.000 0.632 0.005 -0.103 -0.169 0.122 3.527 0.192 0.080 0.083 0.176
Note: $R_{Y}^{2}$ denotes the $R^{2}$ statistic in measurement models for items $Y_{j}$.
Results for 1-step estimates exclude simulations where the estimation did not converge.
This occurred in 0–1 of the 1000 simulations in different settings.
Table A.21: Simulation results for point estimates of structural regression coefficient $\beta_{2}$ of latent covariate $\eta_{2}$ for a latent response $\eta_{3}$, in a model which also includes another latent covariate $\eta_{1}$ (which has coefficient $\beta_{1}$) and where the measurement model includes cross-loadings from $\eta_{1}$ to measures of $\eta_{2}$ (Case F). Note: Here bias of the estimates is quantified by the difference between their median over the simulations and the true value of $\beta_{2}$, whereas in Table A.20 the simulation mean rather than the median was used for this. Other columns of this table identical to those of Table A.20.
Bias (median) RMSE MAE
$R^{2}_{Y}$ $\beta_{1}$ $\beta_{2}$ 2-step 1-step 3-step 2-step 1-step 3-step 2-step 1-step 3-step
$n=200$
0.4 0.390 0.309 -0.023 -0.006 -0.132 0.341 0.935 0.230 0.158 0.176 0.161
0.6 0.390 0.309 -0.002 0.011 -0.088 0.216 0.346 0.177 0.123 0.127 0.127
0.4 0.273 0.423 -0.031 0.014 -0.178 0.470 0.921 0.303 0.173 0.181 0.196
0.6 0.273 0.423 -0.029 -0.007 -0.136 0.231 5.275 0.207 0.139 0.147 0.163
0.4 0.000 0.632 -0.076 -0.004 -0.279 0.528 0.700 7.030 0.236 0.234 0.300
0.6 0.000 0.632 -0.041 -0.013 -0.200 0.291 0.313 0.272 0.176 0.177 0.226
$n=1000$
0.4 0.390 0.309 -0.004 -0.001 -0.125 0.192 0.112 0.135 0.072 0.070 0.125
0.6 0.390 0.309 -0.004 $0$0.091 0.082 1.613 0.106 0.055 0.055 0.093
0.4 0.273 0.423 -0.015 -0.009 -0.173 0.230 0.124 0.181 0.081 0.083 0.173
0.6 0.273 0.423 -0.004 -0.001 -0.120 0.091 0.092 0.134 0.059 0.058 0.121
0.4 0.000 0.632 -0.013 0.005 -0.251 0.160 0.155 0.260 0.099 0.095 0.251
0.6 0.000 0.632 -0.005 -0.002 -0.176 0.122 3.527 0.192 0.080 0.083 0.176
Note: $R_{Y}^{2}$ denotes the $R^{2}$ statistic in measurement models for items $Y_{j}$.
Results for 1-step estimates exclude simulations where the estimation did not converge.
This occurred in 0–1 of the 1000 simulations in different settings.
Table A.22: Simulation results for standard error estimates of structural regression coefficient $\beta_{2}$ of latent covariate $\eta_{2}$ for a latent response $\eta_{3}$, in a model which also includes another latent covariate $\eta_{1}$ (which has coefficient $\beta_{1}$) and where the measurement model includes cross-loadings from $\eta_{1}$ to measures of $\eta_{2}$ (Case F).
Simulation standard deviation 2-step
vs. mean of est. std. error Coverage of without
2-step 1-step 95% conf. interval step-1 var.
$R^{2}_{Y}$ $\beta_{1}$ $\beta_{2}$ s.d. m(se) s.d. m(se) 2-st. 1-st. 3-st. %var. cover.
$n=200$
0.4 0.390 0.309 0.340 0.624 0.933 0.533 94.7 95.0 66.7 61.9 90.1
0.6 0.390 0.309 0.214 0.220 0.343 0.233 94.8 94.4 73.6 64.0 89.5
0.4 0.273 0.423 0.470 2.922 0.919 0.464 94.9 95.6 56.9 54.6 88.9
0.6 0.273 0.423 0.232 0.234 5.273 0.334 92.4 93.0 59.7 55.9 85.7
0.4 0.000 0.632 0.527 2.240 0.684 0.506 88.8 90.7 43.6 46.2 79.1
0.6 0.000 0.632 0.291 0.301 0.310 0.296 90.5 91.7 50.1 44.6 79.3
$n=1000$
0.4 0.390 0.309 0.192 0.164 0.111 0.109 95.1 94.9 38.9 70.8 90.2
0.6 0.390 0.309 0.082 0.081 1.613 0.094 94.8 94.4 48.4 70.5 90.5
0.4 0.273 0.423 0.230 2.564 0.124 0.121 94.1 94.5 24.1 61.8 85.9
0.6 0.273 0.423 0.091 0.092 0.092 0.093 95.1 95.2 34.2 60.1 87.2
0.4 0.000 0.632 0.160 0.159 0.154 0.151 93.3 94.3 12.1 50.5 83.9
0.6 0.000 0.632 0.122 0.120 3.527 0.162 94.2 92.7 23.0 47.2 80.5
$\dagger$ Average % of the variance of 2-step estimator accounted for by step-2 variance alone,
       and coverage of 95% confidence interval if only this variance is included.
Results for 1-step estimates exclude simulations where the estimation did not converge.
This occurred in 0–1 of the 1000 simulations in different settings.
Table A.23: Simulation results for point estimates of structural regression coefficient $\beta_{2}$ of latent covariate $\eta_{2}$ for a latent response $\eta_{3}$, in a model which also includes another latent covariate $\eta_{1}$ (which has coefficient $\beta_{1}$) and where the measurement model includes cross-loadings from $\eta_{1}$ to measures of $\eta_{2}$ (as in case G), but when the measurement model is estimated without the cross-loadings (as in case F) (Case G).
Bias RMSE MAE
$R^{2}_{Y}$ $\beta_{1}$ $\beta_{2}$ 2-step 1-step 3-step 2-step 1-step 3-step 2-step 1-step 3-step
$n=200$
0.4 0.390 0.309 0.246 0.314 0.026 0.634 1.064 0.174 0.263 0.371 0.095
0.6 0.390 0.309 0.252 0.320 0.074 0.461 1.022 0.178 0.247 0.341 0.095
0.4 0.273 0.423 0.369 0.542 -0.033 0.787 1.199 0.198 0.306 0.442 0.123
0.6 0.273 0.423 0.402 0.552 0.038 0.601 0.808 0.187 0.337 0.460 0.112
0.4 0.000 0.632 0.491 0.830 -0.217 0.864 1.356 0.285 0.345 0.543 0.262
0.6 0.000 0.632 0.550 0.796 -0.092 0.774 1.076 0.222 0.456 0.659 0.167
$n=1000$
0.4 0.390 0.309 0.215 0.278 0.007 0.294 0.443 0.065 0.196 0.251 0.044
0.6 0.390 0.309 0.234 0.230 0.061 0.278 2.043 0.089 0.225 0.305 0.061
0.4 0.273 0.423 0.288 0.418 -0.072 0.352 0.912 0.098 0.274 0.345 0.078
0.6 0.273 0.423 0.325 0.468 0.003 0.365 0.888 0.072 0.309 0.416 0.048
0.4 0.000 0.632 0.450 0.714 -0.238 0.527 1.496 0.249 0.408 0.538 0.244
0.6 0.000 0.632 0.482 0.572 -0.124 0.524 3.318 0.146 0.466 0.613 0.128
Note: $R_{Y}^{2}$ denotes the $R^{2}$ statistic in measurement models for items $Y_{j}$.
Results for 1-step estimates exclude simulations where the estimation did not converge.
This occurred in 1 of the 1000 simulations in one setting.
Table A.24: Simulation results for point estimates of structural regression coefficient $\beta_{2}$ of latent covariate $\eta_{2}$ for a latent response $\eta_{3}$, in a model which also includes another latent covariate $\eta_{1}$ (which has coefficient $\beta_{1}$) and where the measurement model includes cross-loadings from $\eta_{1}$ to measures of $\eta_{2}$ (as in case G), but when the measurement model is estimated without the cross-loadings (as in case F) (Case G). Note: Here bias of the estimates is quantified by the difference between their median over the simulations and the true value of $\beta_{2}$, whereas in Table A.23 the simulation mean rather than the median was used for this. Other columns of this table identical to those of Table A.23.
Bias (median) RMSE MAE
$R^{2}_{Y}$ $\beta_{1}$ $\beta_{2}$ 2-step 1-step 3-step 2-step 1-step 3-step 2-step 1-step 3-step
$n=200$
0.4 0.390 0.309 0.180 0.228 -0.008 0.634 1.064 0.174 0.263 0.371 0.095
0.6 0.390 0.309 0.201 0.273 0.049 0.461 1.022 0.178 0.247 0.341 0.095
0.4 0.273 0.423 0.263 0.380 -0.068 0.787 1.199 0.198 0.306 0.442 0.123
0.6 0.273 0.423 0.324 0.454 0.011 0.601 0.808 0.187 0.337 0.460 0.112
0.4 0.000 0.632 0.305 0.526 -0.251 0.864 1.356 0.285 0.345 0.543 0.262
0.6 0.000 0.632 0.456 0.656 -0.126 0.774 1.076 0.222 0.456 0.659 0.167
$n=1000$
0.4 0.390 0.309 0.192 0.249 0.003 0.294 0.443 0.065 0.196 0.251 0.044
0.6 0.390 0.309 0.225 0.304 0.058 0.278 2.043 0.089 0.225 0.305 0.061
0.4 0.273 0.423 0.274 0.345 -0.074 0.352 0.912 0.098 0.274 0.345 0.078
0.6 0.273 0.423 0.309 0.416 -0.003 0.365 0.888 0.072 0.309 0.416 0.048
0.4 0.000 0.632 0.408 0.537 -0.244 0.527 1.496 0.249 0.408 0.538 0.244
0.6 0.000 0.632 0.466 0.612 -0.128 0.524 3.318 0.146 0.466 0.613 0.128
Note: $R_{Y}^{2}$ denotes the $R^{2}$ statistic in measurement models for items $Y_{j}$.
Results for 1-step estimates exclude simulations where the estimation did not converge.
This occurred in 1 of the 1000 simulations in one setting.
Table A.25: Simulation results for standard error estimates of structural regression coefficient $\beta_{2}$ of latent covariate $\eta_{2}$ for a latent response $\eta_{3}$, in a model which also includes another latent covariate $\eta_{1}$ (which has coefficient $\beta_{1}$) and where the measurement model includes cross-loadings from $\eta_{1}$ to measures of $\eta_{2}$ (as in case G), but when the measurement model is estimated without the cross-loadings (as in case F) (Case G).
Simulation standard deviation 2-step
vs. mean of est. std. error Coverage of without
2-step 1-step 95% conf. interval step-1 var.
$R^{2}_{Y}$ $\beta_{1}$ $\beta_{2}$ s.d. m(se) s.d. m(se) 2-st. 1-st. 3-st. %var. cover.
$n=200$
0.4 0.390 0.309 0.585 1.555 1.017 1.639 99.5 99.5 73.9 64.9 95.0
0.6 0.390 0.309 0.386 1.111 0.971 0.519 98.0 98.4 78.5 70.4 92.1
0.4 0.273 0.423 0.695 1.499 1.071 4.246 98.7 98.7 63.9 60.9 92.3
0.6 0.273 0.423 0.446 0.460 0.590 0.545 96.7 97.0 73.7 68.1 86.2
0.4 0.000 0.632 0.712 1.156 1.073 1.398 99.2 97.9 32.5 55.4 93.9
0.6 0.000 0.632 0.544 0.557 0.723 0.671 97.2 97.7 57.4 62.9 83.1
$n=1000$
0.4 0.390 0.309 0.202 0.205 0.344 0.245 85.0 81.6 73.9 79.3 78.4
0.6 0.390 0.309 0.150 0.148 2.031 0.204 66.0 56.6 64.3 84.6 59.1
0.4 0.273 0.423 0.202 0.210 0.811 0.331 77.2 72.1 47.6 79.1 67.6
0.6 0.273 0.423 0.167 0.152 0.755 0.351 44.2 33.9 75.0 84.8 37.2
0.4 0.000 0.632 0.273 0.263 1.315 1.311 67.6 55.2 3.9 74.4 48.3
0.6 0.000 0.632 0.205 0.181 3.270 0.309 20.1 12.5 29.1 81.5 14.2
$\dagger$ Average % of the variance of 2-step estimator accounted for by step-2 variance alone,
       and coverage of 95% confidence interval if only this variance is included.
Results for 1-step estimates exclude simulations where the estimation did not converge.
This occurred in 1 of the 1000 simulations in one settings.
Table B.1: Estimated coefficients of structural models for extrinsic work values given different sets of covariates (with standard errors in parentheses), estimated for EVS2017 data for Netherlands. The table shows two-step (“2-st.”), one-step (“1-st.”) and naive three-step (“3-st.”) estimates. This example is considered in Section 5 of the paper, and the coefficients of gender of respondent (dummy variable for men) are also shown in Table 3 there.
Model (1) Model (2) Model (3)
2-st. 1-st. 3-st. 2-st. 1-st. 3-st. 2-st. 1-st. 3-st.
Intercept 1.403 1.398 1.699 1.385 1.404 1.694 1.155 1.160 1.518
(0.113) (0.116) (0.057) (0.231) (0.238) (0.172) (0.287) (0.294) (0.224)
Man -0.042 -0.033 -0.033 0.025 0.022 0.024 0.018 0.009 0.018
(0.103) (0.104) (0.082) (0.101) (0.104) (0.080) (0.102) (0.104) (0.081)
Age (vs. 30–49)
        15–29 0.472 0.489 0.380 0.448 0.462 0.364
(0.207) (0.214) (0.166) (0.208) (0.213) (0.167)
        50– -0.391 -0.408 -0.326 -0.191 -0.196 -0.162
(0.150) (0.155) (0.119) (0.152) (0.155) (0.123)
Has partner -0.019 -0.022 -0.017 -0.052 -0.056 -0.048
(0.163) (0.169) (0.133) (0.163) (0.167) (0.133)
Age of youngest person in household (vs. older)
        0–5 0.542 0.570 0.437 0.477 0.499 0.390
(0.194) (0.201) (0.151) (0.191) (0.195) (0.150)
        6–17 0.367 0.381 0.299 0.272 0.279 0.225
(0.150) (0.156) (0.120) (0.149) (0.153) (0.121)
Education level (vs. upper secondary)
         Lower -0.104 -0.105 -0.080
(0.144) (0.147) (0.117)
         Higher 0.034 0.031 0.039
(0.133) (0.135) (0.107)
Occupation-based social class (vs. Managerial/service contract)
     Mixed -0.074 -0.074 -0.069
(0.144) (0.147) (0.116)
     Self-empl. -0.350 -0.358 -0.273
(0.243) (0.248) (0.192)
     Labour 0.231 0.229 0.185
(0.144) (0.147) (0.116)
Currently in paid employment 0.449 0.468 0.358
(0.123) (0.125) (0.092)
Parents’ education level (highest, vs. upper secondary)
        Lower -0.122 -0.124 -0.108
(0.146) (0.149) (0.118)
        Higher -0.092 -0.099 -0.092
(0.159) (0.162) (0.128)
Table B.2: Estimated coefficients of structural models for intrinsic work values given different sets of covariates (with standard errors in parentheses), estimated for EVS2017 data for Netherlands. The table shows two-step (“2-st.”), one-step (“1-st.”) and naive three-step (“3-st.”) estimates. This example is considered in Section 5 of the paper, and the coefficients coefficients of gender of respondent (dummy variable for men) are also shown in Table 3 there.
Model (1) Model (2) Model (3)
2-st. 1-st. 3-st. 2-st. 1-st. 3-st. 2-st. 1-st. 3-st.
Intercept 0.624 0.611 0.955 0.474 0.449 0.847 1.311 1.087 1.480
(0.190) (0.165) (0.099) (0.415) (0.376) (0.305) (0.569) (0.430) (0.387)
Man 0.607 0.603 0.460 0.657 0.638 0.500 0.516 0.506 0.400
(0.193) (0.186) (0.142) (0.194) (0.176) (0.142) (0.188) (0.151) (0.139)
Age (vs. 30–49)
        15–29 0.783 0.749 0.573 0.568 0.507 0.417
(0.405) (0.365) (0.293) (0.390) (0.312) (0.288)
        50– -0.584 -0.545 -0.452 -0.183 -0.162 -0.159
(0.293) (0.261) (0.210) (0.283) (0.227) (0.213)
Has partner 0.412 0.371 0.311 0.168 0.108 0.132
(0.328) (0.293) (0.235) (0.311) (0.248) (0.229)
Age of youngest person in household (vs. older)
        0–5 -0.081 -0.072 -0.067 -0.276 -0.221 -0.228
(0.355) (0.324) (0.267) (0.349) (0.278) (0.259)
        6–17 0.028 0.010 0.029 -0.132 -0.130 -0.098
(0.282) (0.257) (0.213) (0.276) (0.221) (0.208)
Education level (vs. upper secondary)
         Lower -0.459 -0.374 -0.378
(0.279) (0.218) (0.202)
         Higher 0.446 0.365 0.342
(0.257) (0.201) (0.185)
Occupation-based social class (vs. Managerial/service contract)
     Mixed -0.588 -0.503 -0.454
(0.279) (0.216) (0.200)
     Self-empl. -0.452 -0.402 -0.337
(0.444) (0.355) (0.332)
     Labour -0.857 -0.748 -0.652
(0.289) (0.220) (0.200)
Currently in paid employment 0.304 0.252 0.237
(0.219) (0.172) (0.160)
Parents’ education level (highest, vs. upper secondary)
        Lower -1.009 -0.796 -0.758
(0.344) (0.239) (0.203)
        Higher -0.238 -0.180 -0.167
(0.306) (0.242) (0.221)
Table B.3: Estimated coefficients of gender (as dummy variable for men) in structural models for extrinsic work values, for each country in EVS2017 data. The table shows two-step, one-step and naive three-step estimates. The models also include as covariates the respondent’s age, education, occupation-based social class and current employment status (working vs. not). All the coefficients of the structrural model are estimated separately for each country, but the measurement model for work values is the same in all countries. The coefficients and confidence intervals from two-step estimation are also also shown in Figure 1 in Section 5 of the paper (with the further standardisation that they are expressed on a scale where the residual variance of extrinsic values for the Netherlands is 1).
country $n$        2-step (s.e.) 1-step (s.e.) 3-step (s.e.)
Albania 1054 -0.180 (0.276) -0.195 (0.320) -0.016 (0.069)
Armenia 1157 -0.133 (0.076) -0.149 (0.089) -0.124 (0.074)
Austria 1450 -0.066 (0.089) -0.045 (0.103) -0.056 (0.079)
Azerbaijan 1205 -0.300 (0.099) -0.346 (0.116) -0.240 (0.081)
Belarus 1452 -0.092 (0.109) -0.096 (0.127) -0.069 (0.078)
Bosnia & Herzegovina 1002 -0.022 (0.123) -0.023 (0.142) -0.019 (0.098)
Bulgaria 1327 -0.085 (0.137) -0.091 (0.156) -0.046 (0.081)
Croatia 1328 -0.313 (0.117) -0.352 (0.135) -0.214 (0.086)
Czechia 1546 -0.108 (0.114) -0.113 (0.131) -0.078 (0.076)
Denmark 3024 -0.143 (0.087) -0.118 (0.100) -0.106 (0.063)
Estonia 1272 -0.088 (0.088) -0.081 (0.104) -0.076 (0.084)
Finland 1035 -0.130 (0.127) -0.111 (0.146) -0.104 (0.099)
France 1729 -0.035 (0.095) -0.027 (0.106) -0.026 (0.078)
Georgia 1756 0.032 (0.077) 0.051 (0.090) 0.035 (0.072)
Germany 1890 -0.159 (0.095) -0.155 (0.109) -0.128 (0.076)
Great Britain 1673 -0.154 (0.110) -0.154 (0.127) -0.105 (0.082)
Hungary 1360 -0.041 (0.100) -0.041 (0.117) -0.023 (0.078)
Iceland 1547 -0.156 (0.075) -0.172 (0.089) -0.139 (0.072)
Italy 1782 -0.141 (0.068) -0.134 (0.077) -0.123 (0.064)
Latvia 1204 -0.389 (0.130) -0.441 (0.149) -0.292 (0.091)
Lithuania 1232 -0.274 (0.138) -0.313 (0.159) -0.142 (0.069)
Montenegro 716 -0.283 (0.135) -0.320 (0.153) -0.237 (0.112)
Netherlands 2068 -0.003 (0.108) 0.032 (0.123) -0.010 (0.077)
North Macedonia 748 0.017 (0.092) 0.022 (0.109) 0.014 (0.081)
Norway 1079 -0.322 (0.113) -0.334 (0.129) -0.266 (0.095)
Poland 1200 -0.213 (0.087) -0.244 (0.098) -0.202 (0.083)
Portugal 1092 0.169 (0.226) 0.215 (0.258) 0.077 (0.097)
Romania 1098 -0.205 (0.135) -0.232 (0.158) -0.139 (0.080)
Russia 1582 -0.397 (0.086) -0.435 (0.097) -0.363 (0.078)
Serbia 1168 0.117 (0.121) 0.133 (0.139) 0.082 (0.085)
Slovakia 1280 -0.493 (0.118) -0.560 (0.134) -0.391 (0.093)
Slovenia 954 -0.212 (0.117) -0.247 (0.140) -0.169 (0.088)
Spain 1002 0.007 (0.140) 0.020 (0.164) -0.003 (0.084)
Sweden 1053 -0.307 (0.146) -0.340 (0.169) -0.225 (0.106)
Switzerland 2841 0.023 (0.072) 0.066 (0.082) 0.029 (0.057)
Ukraine 1455 -0.423 (0.125) -0.476 (0.143) -0.253 (0.073)
Table B.4: Estimated coefficients of gender (as dummy variable for men) in structural models for intrinsic work values, for each country in EVS2017 data. The table shows two-step, one-step and naive three-step estimates. The models also include as covariates the respondent’s age, education, occupation-based social class and current employment status (working vs. not). All the coefficients of the structural model are estimated separately for each country, but the measurement model for work values is the same in all countries. The coefficients and confidence intervals from two-step estimation are also also shown in Figure 1 in Section 5 of the paper (with the further standardisation that they are expressed on a scale where the residual variance of intrinsic values for the Netherlands is 1).
country $n$        2-step (s.e.) 1-step (s.e.) 3-step (s.e.)
Albania 1054 0.152 (0.237) 0.141 (0.235) 0.092 (0.122)
Armenia 1157 -0.051 (0.201) -0.052 (0.196) -0.039 (0.142)
Austria 1450 0.010 (0.122) 0.014 (0.119) 0.006 (0.119)
Azerbaijan 1205 0.010 (0.229) 0.003 (0.222) 0.024 (0.160)
Belarus 1452 0.235 (0.191) 0.239 (0.184) 0.164 (0.141)
Bosnia & Herzegovina 1002 -0.269 (0.161) -0.268 (0.153) -0.229 (0.145)
Bulgaria 1327 0.277 (0.249) 0.269 (0.242) 0.185 (0.137)
Croatia 1328 -0.204 (0.184) -0.204 (0.178) -0.145 (0.137)
Czechia 1546 0.463 (0.196) 0.454 (0.190) 0.295 (0.130)
Denmark 3024 0.004 (0.095) -0.007 (0.092) 0.017 (0.079)
Estonia 1272 0.474 (0.175) 0.449 (0.169) 0.407 (0.146)
Finland 1035 0.105 (0.147) 0.096 (0.143) 0.112 (0.136)
France 1729 -0.024 (0.116) -0.035 (0.113) -0.026 (0.107)
Georgia 1756 0.854 (0.188) 0.813 (0.181) 0.558 (0.128)
Germany 1890 0.287 (0.116) 0.284 (0.114) 0.252 (0.101)
Great Britain 1673 0.087 (0.125) 0.077 (0.121) 0.076 (0.110)
Hungary 1360 0.268 (0.186) 0.253 (0.181) 0.203 (0.134)
Iceland 1547 -0.155 (0.121) -0.165 (0.118) -0.137 (0.109)
Italy 1782 0.214 (0.081) 0.192 (0.077) 0.248 (0.093)
Latvia 1204 0.049 (0.228) 0.058 (0.219) 0.033 (0.156)
Lithuania 1232 -0.033 (0.312) -0.036 (0.304) -0.047 (0.147)
Montenegro 716 -1.196 (0.475) -1.146 (0.454) -0.556 (0.213)
Netherlands 2068 0.366 (0.117) 0.355 (0.114) 0.334 (0.105)
North Macedonia 748 -0.063 (0.260) -0.068 (0.254) -0.007 (0.148)
Norway 1079 -0.190 (0.130) -0.196 (0.126) -0.174 (0.126)
Poland 1200 0.101 (0.131) 0.084 (0.125) 0.098 (0.131)
Portugal 1092 0.501 (0.311) 0.478 (0.302) 0.239 (0.147)
Romania 1098 -0.865 (0.277) -0.849 (0.271) -0.462 (0.134)
Russia 1582 0.054 (0.135) 0.048 (0.128) 0.039 (0.107)
Serbia 1168 -0.101 (0.238) -0.101 (0.230) -0.078 (0.156)
Slovakia 1280 -0.117 (0.180) -0.105 (0.173) -0.104 (0.145)
Slovenia 954 0.262 (0.320) 0.253 (0.312) 0.103 (0.106)
Spain 1002 0.147 (0.285) 0.134 (0.277) 0.085 (0.159)
Sweden 1053 -0.681 (0.156) -0.689 (0.153) -0.495 (0.116)
Switzerland 2841 0.084 (0.078) 0.077 (0.076) 0.086 (0.077)
Ukraine 1455 0.266 (0.220) 0.262 (0.213) 0.169 (0.138)